### v1

# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

Abstract
We introduce a stochastic model of a population with overlapping generations and arbitrary levels of self-fertilization versus outcrossing. We study how the global graph of reproductive relationships, or pedigree, influences the genealogical relationships of a sample of two gene copies at a genetic locus. Specifically, we consider a diploid Moran model with constant population size $N$ over time, in which a proportion of offspring are produced by selfing. We show that the conditional distribution of the pairwise coalescence time at a single locus given the random pedigree converges to a limit law as $N$ tends to infinity. This limit law generally differs from the corresponding traditional result obtained by averaging over all possible pedigrees. We describe three different behaviors in the limit depending on the relative strengths, from large to small, of selfing versus outcrossing: partial selfing, limited outcrossing, and negligible outcrossing. In the case of partial selfing, coalescence times are related to the Kingman coalescent, similar to what is found in traditional analyses. In the the case of limited outcrossing, the retained pedigree information forms a random graph coming from a fragmentation-coagulation process, with the limiting coalescence times given by the meeting times of random walks on this graph. In the case of negligible outcrossing, which represents complete or nearly complete selfing, coalescence times are determined entirely by the fixed times to common ancestry of diploid individuals in the pedigree.
1 Introduction
Modes of reproduction and genetic transmission are fundamental aspects of evolution (Charlesworth and Wright, 2001; Glémin et al., 2019). In this work, we consider autogamy or self-fertilization—hereafter just ‘selfing’—in which meiosis occurs then a zygote is formed by the union of two gametes from the same individual. This form of inbreeding has been especially important in the evolution of plants (Jain, 1976; Uyenoyama et al., 1993; Charlesworth, 2006; Olsen et al., 2021). The common alternative to selfing in plants and animals is outcrossing, in which zygotes are formed by the union of two gametes from different individuals. The selfing rate, defined here as the probability that an individual is produced by selfing, ranges from 00 to $1$. We describe three distinct kinds of coalescent models which hold for different magnitudes of the selfing rate when the process of coalescence is conditioned on the population pedigree, or the organismal genealogy of the population (Ball et al., 1990). This important latent variable has been mostly ignored in theoretical population genetics; see Section 1.3 of this paper and Diamantidis et al. (2024).
The conditional coalescent models we describe exist in the limit as the population size $N$ tends to infinity in the same way that the Wright-Fisher diffusion and the corresponding standard neutral coalescent model exist in this limit (Kingman, 1982; Möhle, 1999; Ewens, 2004). We use $\alpha_{N}$ to denote the selfing rate, with explicit dependence on $N$ so we can control the relative importance of selfing in the model. Our results for coalescence in large populations conditional on the pedigree depend strongly on how $\alpha_{N}$ changes as $N\to\infty$. Most previous analyses, in addition to not conditioning on the pedigree, have assumed a fixed selfing rate. This rate is often denoted $s$, though various notations have been adopted; e.g. $l$ in Haldane (1924), $h$ in Wright (1951), $\beta$ in Pollak (1987) and $\alpha$ in Karlin (1968a, b). As Uyenoyama (2024) recently emphasized, some authors use $s$ for the entire selfing rate while others use it to measure selfing in excess of a background rate $1/N$ due to random mating. Our model includes these as special cases, namely if we assume $\alpha_{N}=s$ or $\alpha_{N}=s+(1-s)/N$, respectively. These two models with fixed $s$ make the same predictions in the limit and both have been called ‘partial self-fertilization’ or ‘partial selfing’.
We use $\alpha\in[0,1]$ to denote the limit of $\alpha_{N}$ as $N\to\infty$. Conditioning coalescence on the pedigree reveals three markedly different behaviors near the boundary $\alpha=1$, depending on how quickly $\alpha_{N}\to 1$. The critical case, which we call ‘limited outcrossing’, is when $1-\alpha_{N}$ is of order $1/N$. Then the conditional coalescent process is characterized by a system of coalescing random walks on a random ancestral graph like the ones previously described for recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). If outcrossing is even less frequent, for example $1-\alpha_{N}$ of order $1/N^{2}$, the limiting conditional coalescent process is fully determined by the random pedigree of the population, which becomes just a Kingman coalescent tree connecting diploid individuals. We call this ‘negligible outcrossing’. If outcrossing is more frequent than in the critical case, for example $1-\alpha_{N}$ of order $1/\sqrt{N}$, the limiting conditional coalescent model resembles the previously described (unconditional) coalescent model of Nordborg and Donnelly (1997) and Möhle (1998). Here we adopt the previous term ‘partial selfing’. The latter model holds for all $\alpha\in[0,1]$ as long as the rate of approach to $\alpha=1$ is not too fast, whereas limited outcrossing and negligible outcrossing hold just for $\alpha=1$.
In the next two sections, we review previous works which generally use $s$ for the fixed selfing rate, take $s\in[0,1]$ and do not condition on the population pedigree. We will reserve our notation $\alpha_{N}$ for Section 1.3 and later. In Section 1.3, we explain how previous population-genetic models in fact average over pedigrees. In Section 1.1 and Section 1.2 we use $s$ to maintain continuity with what one finds in much of the evolutionary and population-genetic literature.
1.1 Prevalence, causes and consequences of selfing in evolution
Although the ability to self-fertilize is rare among animals (Jarne and Auld, 2006; Avise and Mank, 2009; Escobar et al., 2011) it is found in other taxa including eukaryotic microbes and marine invertebrates (Sasson and Ryan, 2017; Yadav et al., 2023) and is common in plants (Abbott and Gomes, 1989; Hartfield et al., 2017; Teterina et al., 2023). About $10$-$15$% of plant species are predominantly selfing (Wright et al., 2013) as are two of the best-studied genetic model species, Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). The empirical distribution of selfing rates among plant species is markedly bimodal, with relatively fewer species having $s\in(0.2,0.8)$ (Schemske and Lande, 1985; Vogler and Kalisz, 2001; Goodwillie et al., 2005). Genetical models of inbreeding depression can explain this bimodality (Lande and Schemske, 1985; Charlesworth and Willis, 2009). But, whether selfing is favored over outcrossing or vice versa in a given species depends on a number of ecological and evolutionary factors (Charlesworth, 2006; Wright et al., 2013).
Selfing is a form of inbreeding which has three important consequences for the distribution of genetic variation. First, selfing increases single-locus homozygosity. Haldane (1924) described this and the associated decrease in heterozygosity compared to Hardy-Weinberg expectations in an infinite population with selfing rate $s\in[0,1]$. His result may best be described using single-locus probabilities of identity by descent. See Rousset (2002) for a synthetic treatment of identity by descent and other measures of inbreeding and relatedness. In terms of Wright’s individual inbreeding coefficient $F$ (Wright, 1931, 1951), what Haldane (1924) showed was that the reduction in the proportion of heterozygous individuals in the population at equilibrium is given by $F=s/(2-s)$. A simple coalescent interpretation is that $F$ is the probability the two alleles or gene copies at a locus in an individual coalesce in the recent ancestry of the individual, rather than being separated into different individuals by an outcrossing event (Nordborg and Donnelly, 1997).
Second, selfing reduces the effective rate of recombination. Bennett and Binet (1956) discovered this effect by modeling genotype frequencies at two linked loci, and related it to the decrease in single-locus heterozygosity due to selfing. Weir and Cockerham (1973) confirmed this effect in their extensive study of two-locus identity coefficients. Recombination is similarly reduced in finite populations (Golding and Strobeck, 1980; Vitalis and Couvet, 2001). Nordborg (2000) provided an interpretation in terms of coalescence. Backward in time, a recombination event places ancestral genetic material onto two chromosomes. These are necessarily in the same parental individual which might itself have been produced by selfing. Then the same coefficient $F$ is the probability the recombination event gets healed or repaired by a coalescent event within the individual’s recent ancestry. As a result, only a fraction $1-F=2(1-s)/(2-s)$ of recombination events are observable.
Third, selfing generates non-random associations of diploid genotypes at different loci, even if the loci are unlinked. As Smith (2023) recently emphasized, both multi-locus homozygotes and multi-locus heterozygotes are overrepresented. Cockerham and Weir (1968) and Weir and Cockerham (1969, 1973) introduced the term ‘identity disequilibrium’ for this effect, treating it in further detail and with helpful overviews in Cockerham and Weir (1977, 1983, 1984). Haldane (1949) was the first to describe the effect, for inbreeding in general not just selfing, specifically as a consequence of variation in the level of inbreeding among individuals in the population. Bennett and Binet (1956) and Kimura (1958) found and quantified identity disequilibrium for a pair of loci subject to recombination and partial selfing. Weir and Cockerham (1973) partitioned their result into a sum of two terms, one due to linkage, which is equal to zero for unlinked loci, and one due to selfing, which is equal to the theoretical variance of $F$ among individuals in the population:
$\sigma^{2}_{F}=\frac{4s(1-s)}{(2-s)^{2}(4-s)}.$ (1)
Weir and Cockerham (1973) considered a number of different coefficients of relatedness, and use $\sigma^{2}_{F_{1}}$ in their version of (1). Narain (1966) described many of the same phenomena as Weir and Cockerham (1973) but in terms of two-locus probabilities of identity by descent directly, rather than in terms of identity disequilibria which are the corresponding (co)variances.
A coalescent interpretation of the identity disequilibrium for two unlinked loci is as follows. Let $Y_{i}$ be an indicator variable, equal to $1$ if the gene copies at locus $i$ in a particular individual coalesce in its recent ancestry, and equal to 00 otherwise. The recent ancestry of a randomly sampled individual is given by the number of generations of selfing in its ancestral line before the first outcrossing event. As Weir and Cockerham (1973) note, this formulation originates with Cockerham and Rawlings (1967). For an individual with $k$ generations of selfing in its immediate ancestry, the probability the two gene copies at a locus remain distinct (i.e. do not coalesce) in any of these generations is $2^{-k}$. If $U$ denotes the random number of selfing generations for an individual,
$\mathbb{P}(U=k)=s^{k}(1-s)\qquad k=0,1,2,\ldots$ (2)
and the coefficient $F$ can be seen as the average
$F\coloneq\mathbb{E}\left[Y_{i}\right]=\sum_{k=0}^{\infty}\left(1-2^{-k}\right)s^{k}(1-s)=\frac{s}{2-s}.$ (3)
The covariance of $Y_{i}$ and $Y_{j}$ for two unlinked loci, which are conditionally independent given $k$, is
$\mathrm{Cov}{\left[Y_{i},Y_{j}\right]}=\sum_{k=0}^{\infty}\left(1-2^{-k}\right)^{2}s^{k}(1-s)-\left(\frac{s}{2-s}\right)^{2}$ (4)
which is identical to (1), since $\sigma^{2}_{F}\coloneq\mathrm{Var}{\left[Y_{i}\right]}$ is also given by (4). Interpreting identity disequilibria as covariances is sensible because they occur between loci. In anticipation of later results conditional on the population pedigree, it is important that (2) here is intended as the theoretical distribution for a single individual sampled at random from a given population. In keeping with Haldane (1949), (4) is an average which quantifies variation among individuals.
Owing to these three effects—increased homozygosity, reduced recombination, and identity disequilibrium—partially selfing populations follow different evolutionary paths than randomly mating populations. Selfing alters the evolution of the recombination rate (Charlesworth et al., 1977; Roze and Lenormand, 2005; Stetsenko and Roze, 2022), leads to apparent or spurious multi-locus signals of selection (Cockerham and Rawlings, 1967; Ohta and Cockerham, 1974; Charlesworth, 1991), determines levels of inbreeding depression (Nei, 1965; Hansche et al., 1966; Campbell, 1986; Charlesworth et al., 1990; Lande et al., 1994; Schultz and Willis, 1995; Abu Awad and Roze, 2018), and in many ways alters the responses of populations to selection (Kondrashov, 1985; Wright and Cockerham, 1985; Charlesworth et al., 1993; Nordborg et al., 1996; Willis, 1996; Kelly, 1999a, b, 2007; Lande and Porcher, 2015; Roze, 2016; Sachdeva, 2019). Among these evolutionary studies, Cockerham and Rawlings (1967), Campbell (1986), Kelly (1999a, b, 2007) and Lande and Porcher (2015) explicitly model the variation in the number of selfing generations in the recent ancestry of individuals, given here under neutrality by (2).
The effects of selfing on the distribution of genetic variation are also important determinants of the evolution of the selfing rate (Charlesworth et al., 1990; Uyenoyama and Waller, 1991a, b, c; Abu Awad and Roze, 2020) but in themselves do not explain the prevalence of selfing. In the absence of other evolutionary forces, selfing is strongly favored over outcrossing (Fisher, 1941; Moran, 1962; Karlin and McGregor, 1974; Nagylaki, 1976; Lloyd, 1979; Wells, 1979). As Maynard Smith (1971) observed in discussing the evolution of sexual reproduction or meiosis, outcrossing individuals are at a disadvantage because they transmit only half of their genome to each offspring. The comparison in Maynard Smith (1971) was to parthenogenesis or asexual reproduction, but selfing is a way to alleviate this “cost of sex” in organisms which necessarily undergo meiosis. Thus, while Williams and Mitton (1973) attributed this cost to meiosis itself, Uyenoyama (1986) more precisely reckoned it a cost of outcrossing. The fact that outcrossing is so common can be explained in genetical terms by the avoidance of inbreeding depression (Charlesworth and Willis, 2009; Brown and Kelly, 2020) or the purging or deleterious mutations in the presence of recombination (Kondrashov, 1985; Keightley and Otto, 2006; Kamran-Disfani and Agrawal, 2014).
In addition, many ecological factors affect the evolutionary dynamics of selfing species (Jain, 1976; Solbrig, 1976). Pioneering work by Baker (1955) and Stebbins (1957) showed that selfing may be favored when the usual modes of cross-fertilization are blocked, in plants for example by excessive rainfall or lack of pollinators, or when the environment presents opportunities for colonization starting from one or a few individuals, for example if rare long-distance dispersal events can access open habitat or if there is rapid turnover of suitable habitat patches. Ingvarsson (2002) showed that the last of these, specifically extinction-recolonization dynamics in a metapopulation, could explain the common observation that levels of genetic variation in selfing species are often far below what is predicted by simple neutral population-genetic models (see Section 1.2). Selfing species with their characteristic suite of phenotypes have evolved many times from outcrossing ancestors, prompting questions about whether it might be an evolutionary “dead end” (Stebbins, 1957; Ornduff, 1969; Sicard and Lenhard, 2011; Wright et al., 2013; Cutter, 2019). Within a single species, subpopulations of selfers are commonly restricted to patches at the margins of the habitat (Baker, 1955; Stebbins, 1957), as for example in the case of Leavenworthia alabamica detailed in Busch (2005).
1.2 Population genetics of selfing under neutrality
Our goal in this work is to understand how selfing affects the sampling structure of gene genealogies under neutrality when the process of coalescence is conditioned on the pedigree of the population. We ask whether results conditional on the pedigree differ from those of previous theoretical treatments which have disregarded this feature of the population. For simplicity we use a diploid Moran model of reproduction (Moran, 1958, 1962; Linder, 2009; Coron and Le Jan, 2022), which we will describe in Section 2. Previous theoretical treatments, including those we review here, have assumed Wright-Fisher reproduction (Fisher, 1930; Wright, 1931).
Based on the considerations of Section 1.1, it may be that selfing species are even less likely than outcrossing ones to meet the assumptions of standard models of population genetics. Both the Wright-Fisher model and the Moran model apply to a single well mixed population of constant in size $N$ without selection. The simplest application relevant to selfing is to diploid, monoecious, randomly mating organisms, in which case selfing happens with probability $1/N$. When $N$ can be assumed to be large, both models converge to the Wright-Fisher diffusion with its corresponding coalescent process (Kingman, 1982; Hudson, 1983a; Tajima, 1983; Möhle, 1999; Ewens, 2004) which exist in the limit $N\to\infty$ with time measured in units proportional to $N$ generations. This provides a common framework for studying more complicated populations, as deviations from this simple null model. Selfing is one such deviation, which has been incorporated into the model by assuming that the selfing rate $s$ is a fixed constant when the population size $N$ tends to infinity.
Pollak (1987) took a forward-time approach to studying, among other things, the equilibrium probabilities of identity by descent of two gene copies in the same individual or in different individuals, when $s\gg N^{-1}$. Let us call these probabilities $\Phi_{1}$ and $\Phi_{2}$, respectively. For large $N$, Pollak (1987) found that the expression for $\Phi_{2}$ was identical to that for a randomly mating population, if $N_{e}=N/(1+F)=(2-s)N/2$ is used in place of $N$. As $s$ increases from 00 to $1$, $N_{e}$ decreases from $N$ to $N/2$, effectively taking the value of $\Phi_{2}$ from that for a diploid (randomly mating) population to that for a haploid population. Pollak (1987) identified this as a separation-of-time-scales result, stemming from the assumption $s\gg N^{-1}$, and suggested that the two-times-scales diffusion approximation of Ethier and Nagylaki (1980) could be applied to prove convergence to the Wright-Fisher diffusion with this value of $N_{e}$. The separation of time scales can be seen in the relationship Pollak (1987) found between $\Phi_{1}$ and $\Phi_{2}$, namely that $\Phi_{1}=F+(1-F)\Phi_{2}$, which may best be understood in terms of subsequent descriptions of the coalescent process.
[…]
It was in this context that Möhle (1998) proved an important general result (Möhle, 1998, Lemma 1, Theorem 1) for Markov processes with two time scales. Owing to the duality relation between the coalescent and diffusion models (Möhle, 1999) this result can be seen as the backward-time counterpart to the result of Ethier and Nagylaki (1980) cited by Pollak (1987). For the selfing model with fixed $s$ as $N\to\infty$, this gave rigorous justification to the model of Nordborg and Donnelly (1997): whenever a pair of lineages is in the same individual, they coalesce with probability $F=s/(2-s)$ instantaneously, otherwise ending up in different individuals with probability $1-F$, while samples from different individuals obey the standard neutral coalescent process with time in units of $(2-s)N$ generations (Möhle, 1998).
A somewhat different approach to the coalescent process with partial selfing was taken by Fu (1997), who approximated the single-generation backward-time transition probabilities for the number of ancestral lineages and the number of individuals containing them, up to order $1/N$, then used these in an algorithm to simulate coalescent ancestries. Time remains in generations, and the algorithm generates coalescence times of some few generations for pairs of gene copies within individuals. Something like this could be done by developing an algorithm for the pre-limiting ancestral process described in sections 3.2 and 3.3 in Möhle (1998). But it would not produce identical results, because Möhle (1998), also Nordborg and Donnelly (1997), used $s$ for the total proportion of selfing, while Fu (1997) used $s$ for the excess proportion of selfing, for a total of $s+(1-s)/N$.
1.3 Accounting for the pedigree of the population
The pedigree of a population is a graph representing the set of reproductive relationships of all individuals in the population for all time. Previous studies of the population-genetic effects of selfing, including those reviewed in Section 1.2, have followed standard practice in population genetics by averaging over all possible outcomes of reproduction, which is equivalent to averaging over the pedigree. See Diamantidis et al. (2024) for a recent detailed discussion of the issues surrounding this tradition in population genetics, which we review briefly here.
Averaging over outcomes of reproduction is often done implicitly, for example by stating that two ancestral lines coalesce with probability $1/(2N)$ each generation under the diploid monoecious Wright-Fisher model. Sometimes it is explicit, for example in equation (4.1) in Kingman (1982) for the coalescence probability under Cannings’ haploid exchangeable model (Cannings, 1974). For diploids in particular, the practice of averaging is highly questionable because all reproductive outcomes in the population for all time are encoded in a fixed structure, the genealogy or pedigree of the population through which all genetic transmission must also have occurred (Ball et al., 1990). In the most common application of coalescent models, which is to describe the distribution of gene genealogies or genetic variation among loci, the practice of averaging over reproductive outcomes is plainly wrong because the same pedigree holds for all loci.
A better-justified starting point for coalescent theory is to condition the ancestral genetic process on the pedigree. This gives the correct sampling structure for multiple unlinked loci: conditionally independent realizations of the process of genetic transmission within the same population/pedigree. The sampling structure implicit in traditionally formulated coalescent models is that each locus has its own pedigree, as if each locus was sampled from an independent population.
It turns out that the standard neutral coalescent model, at least, can still be applied because the coalescent model conditional on the pedigree is the same as the standard coalescent model for populations in the domain of the Kingman coalescent (Tyukin, 2015; Diamantidis et al., 2024). The reason is that pedigrees in large well mixed populations have a characteristic mixing time of $\log_{2}N$ generations (Chang, 1999; Derrida et al., 1999; Barton and Etheridge, 2011; Wakeley et al., 2012). But it is not true in general that extensions of the standard neutral coalescent model are similarly applicable to coalescent processes conditional on pedigrees generated under the corresponding population models (Wakeley et al., 2016; Wilton et al., 2017; Diamantidis et al., 2024).
The coalescent models of partial selfing in Fu (1997), Nordborg and Donnelly (1997) and Möhle (1998) average over the process of reproduction and thus implicitly over the pedigree. One consequence of instead conditioning on the pedigree is that each sampled individual has its own realized number of generations of selfing before the first outcrossing event in its ancestry. This, as we will show, is a single draw from the distribution (2). The end result of averaging is clear in previous approaches: $F$ in (3) is such an average. The averaging procedure is illustrated by the matrix $A^{m}$ in Möhle (1998, p. 496) which gives the $m$-generation transition probabilities of the fast process. For example,
$\left(A^{m}\right)_{32}=(1-s)\frac{1-(s/2)^{m}}{1-s/2}$
is the probability that the ancestors of two gene copies currently in the same individual remain distinct and are in two different individuals in past generation $m$. This can be expressed as a sum over the distribution (2), namely
$\left(A^{m}\right)_{32}=\sum_{k=0}^{m-1}\left(\frac{1}{2}\right)^{k}s^{k}(1-s).$ (5)
Lemma 1 and Theorem 1 of Möhle (1998) use the limit of the fast process, $P\coloneq\lim_{m\to\infty}A^{m}$, in which (5) converges to $2(1-s)/(2-s)=1-F$, or $\mathbb{E}\left[1-Y_{i}\right]$ for the indicator random variable $Y_{i}$ defined in Section 1.1.
Partial selfing is just one of three possibilities when the coalescent process is conditioned on the pedigree. Here we reintroduce our notation, in which $\alpha_{N}$ denotes the selfing rate in a population of size $N$, and $\alpha$ denotes its value in the limit $N\to\infty$ or in the special case that $\alpha_{N}=\alpha$ is constant. When selfing is the primary mode of reproduction, an entirely new sort of model emerges in the limit, one which is completely invisible in the unconditional pedigree-averaging process. This novel limit process occurs when the outcrossing rate $1-\alpha_{N}$ is of order $N^{-1}$. It is parameterized by $\lambda\coloneq\lim_{N\to\infty}N(1-\alpha_{N})$, which we note is the same way mutation rates are assumed to scale in the Wright-Fisher diffusion and corresponding standard neutral coalescent model (Kingman, 1982; Hudson, 1983a; Tajima, 1983; Ewens, 2004) and how migration rates are scaled in the structured coalescent model (Takahata, 1988; Notohara, 1990; Herbots, 1997). In the limit we consider, $\lambda$ is the instantaneous rate of outcrossing events per ancestral lineage.
Figure 1 shows simulation results of pedigrees and pairwise coalescence times for the model of selfing used by Möhle (1998), which corresponds to $\alpha_{N}=\alpha$ constant in our model. Specifically, for each of three selfing rates, $50$ independent pedigrees were simulated and for each of these two different individuals were sampled at random. Colored lines, one for each pedigree-plus-sample, depict the cumulative distribution function (CDF) of the coalescence time for two gene copies, one from each of the two sampled individuals. On the scale used in Figure 1, the corresponding expectation for the pedigree-averaging model (not depicted) would be the CDF of an exponential distribution with parameter $(2-\alpha)^{-1}$. Many of the CDFs in Figure 1(a) have something close to this shape, but those in Figure 1(b) and Figure 1(c) look nothing like it. Instead they are characterized by random jumps in probability at random times.
   Refer to caption   Refer to caption
Figure 1: Cumulative distribution functions of the coalescence times for a sample of size $n=2$ on each of $50$ pedigrees simulated under the Wright-Fisher model with partial selfing and $N=1000$ individuals. Panels (a), (b) and (c) show the results for three different values of the probability of selfing, respectively, for $\alpha\in\{0.9,0.99,0.999\}$. Coalescence probabilities in each generation given the pedigree and the sampled individuals were calculated numerically by the method described in Wakeley et al (2012).
The deviations from the predictions of the unconditional coalescent model displayed in Figure 1 are not due to the relatively small population size used in these simulations ($N=1000$). Instead they illustrate that the behavior we will prove for our Moran model occurs when $N(1-\alpha_{N})$ remains finite in the limit $N\to\infty$ is relevant even for such small populations. The value of $\alpha=0.99$ used in Figure 1(b) is similar to estimates of the selfing rates for the model species Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). Of course, $N=1000$ would be a very small population size for either of these species, though it may be realistic for some local populations. In any case, species or populations for which $N(1-\alpha)$ is not very large will have genetic ancestries very unlike what is expected under the unconditional or pedigree-averaging model. Instead, they are captured by the conditional coalescent model with ‘limited outcrossing’ which is characterized by a system of coalescing random walks on a random ancestral graph, described in Section 4.
1.4 Plan of the present work
Our analysis is based on a diploid Moran model of reproduction with constant population size $N$ and selfing rate $\alpha_{N}$, described in detail in Section 2. We focus on the coalescence time for a sample of size $n=2$, which already reveals the non-trivial effects of the pedigree on gene genealogies.
To establish the conditional limit for the limiting-outcrossing regime, we prove that the trajectories of potential ancestral individuals of the sample converge weakly to the ancestral graph, which is different from the proof for the partial-selfing regime. For the latter we adapt the method of Diamantidis et al. (2024), which itself was adapted from Birkner et al. (2013). We convert the problem of conditional convergence to $L^{2}$ convergence for the distribution of the minimum of two conditionally independent pairwise coalescence times for two independently assorting loci on the pedigree. Unlike the work in Diamantidis et al. (2024), here in both the conditional and unconditional case Möhle’s result (Möhle, 1998, Lemma 1, Theorem 1) does not apply, at least not in the obvious way. Some of our results could be obtained using the extension by Möhle and Notohara (2016), but here we take an alternate approach. We compute coalescence times by examining the random number of opportunities for coalescence between genetic lineages in different individuals, which we will refer to as overlap events, and the timings of these in relation to splitting events in the ancestral process, which result from outcrossing.
The organization of the paper is as follows. In Section 2, after detailing our Moran model, we make precise the pedigree and the gene genealogy. Our results for the unconditional and the conditional distributions of pairwise coalescence times are presented in Section 3 and Section 4, respectively. While we focus on $n=2$ in this paper, we briefly discuss the case $n>2$ and offer a conjecture in Section 4.2. In Section 5, we establish some fundamental properties of the scaling limit under limited outcrossing. This scaling limit is a coalescing random walk on a random ancestral graph. Our main convergence theorems will be proved in Section 6.
2 A discrete-time diploid Moran model with selfing
We consider a diploid, monoecious, panmictic (well mixed and randomly mating) population of constant $N$ over time. Individuals are labeled $I=\{1,...,N\}$. The population has overlapping generations and changes at discrete time-steps. At each time-step, a randomly chosen individual dies and is replaced by an offspring produced in one of two ways. With probability $1-\alpha_{N}$ the offspring has two distinct parents; with probability $\alpha_{N}$ the offspring has exactly one parent. The parent(s) are chosen uniformly at random from the previous time-step. In the case of two parents, these are chosen without replacement. The reproduction events at different time-steps are independent and identically distributed.
Explicitly, for each non-negative integer $k\in\mathbb{Z}_{+}=\{0,1,2,3,...\}$, the reproduction event between consecutive time-steps $k$ and $k+1$ in the past is as follows:
  1. 1.
    (distinct parents) With probability $1-\alpha_{N}$, a single offspring is produced by outcrossing. A triplet containing the two parents and one offspring $(\pi_{k,1},\,\pi_{k,2},\,\kappa_{k})\in I^{3}$ is chosen uniformly at random with $\pi_{k,1}\neq\pi_{k,2}$ (i.e. the parents are distinct). The offspring $\kappa_{k}$ is an individual in time-step $k$, while the offspring’s parents are the individuals in time-step $k+1$ with labels $\pi_{k,1}$ and $\pi_{k,2}$.
    The offspring has two genes copies, one inherited from each parent. Genetically, the offspring is produced according to Mendel’s laws, which means each of the two gene copies in a parent is equally likely to be the one transmitted to the offspring. An example outcrossing event in a population of size $N=7$ is depicted below.
    ($\kappa_{k}=3$)   time-step k($\pi_{k,1}=2,\pi_{k,2}=4$)  time-step k+1
  2. 2.
    (selfing) With probability $\alpha_{N}$, a parent-offspring pair with labels $(\pi_{k},\,\kappa_{k})\in I^{2}$ is chosen uniformly at random without replacement. The offspring $\kappa_{k}$ in time-step $k$ is the offspring of a single individual (the parent) with label $\pi_{k}$ at time-step $k+1$. This can be viewed as having $\pi_{k,2}=\pi_{k,1}$ in the previous reproduction event.
    As above, the offspring is produced according to Mendel’s laws which means each of the two gene copies in the parent is independently and equally likely to be inherited by each of those of the offspring. An example selfing event in a population of size $N=7$ is depicted below.
    ($\kappa_{k}=3$)    time-step k($\pi_{k}=4$)   time-step k+1
The reproduction events at different time-steps are independent. Note that at time-step $k$, the offspring $\kappa_{k}$ is a new individual introduced and it replaces the one with the same label in the previous step $k+1$ (death-birth for the label $\kappa_{k}$); all the remaining $N-1$ individuals in time-step $k$ are the same individuals as those in time-step $k+1$. Hence this model has overlapping generations.
Remark 2.1.
When $\alpha_{N}=0$, our model is exactly the one introduced in Coron and Le Jan (2022). When $\alpha_{N}$ is equal to a fixed $p$ in $[0,1)$, our model is exactly that in Linder (2009).
2.1 The pedigree as important partial information of the population
The above population dynamics generates a random graph, which we will call $\mathcal{G}_{N}$, with vertex set $I\times\mathbb{Z}_{+}$ (individuals) and directed edges joining each individual with its parent(s). This random graph $\mathcal{G}_{N}$ is called the pedigree of the population. In the following, we consider the pedigree without specifying outcomes of genetic transmission.
Between consecutive time-steps with two distinct parents, the pedigree has one edge from offspring $(\kappa_{k},k)$ to parent $(\pi_{k,1},k+1)$, one from offspring $(\kappa_{k},k)$ to parent $(\pi_{k,2},k+1)$, and $N-1$ single edges for the same individual from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\kappa_{k}\}$. The portion of the pedigree corresponding to the example above is
[…]
Between consecutive time-steps with selfing, the pedigree has a double edges from $(\kappa_{k},k)$ to $(\pi_{k},k+1)$, and again $N-1$ single edges from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\kappa_{k}\}$. The portion of the pedigree corresponding to the example above is
[…]
Hence, the offspring always has 2 edges coming out of it which trace upward (towards the past); and all the other $N-1$ individuals have a single edge.
Repetition of this process with these two possibilities for each time step backward into the indefinite past results in a single realized pedigree of the population. Patterns of genetic ancestry at each locus are outcomes of Mendelian transmission conditional on this one pedigree. Unlinked loci are transmitted independently through the pedigree. In Figure 2, we illustrate a realization of both the population dynamics with genetic transmission and the corresponding pedigree for the first 4 generations in the past for a population of size $6$.
Figure 2: A realization of our diploid Moran process with $N=6$ individuals (left image) and the corresponding pedigree (right image). On the left, two genes $(X_{0},Y_{0})=(2,9)$ are sampled in the present time-step 00, and their lineages are highlighted by the solid dots. Namely, $\{(X_{k},Y_{k})\}_{k=0}^{4}=\{(2,9),(2,9),(2,5),(6,5),(4,4)\}$. These two lineages coalesce in time-step $4$ because $X_{4}=Y_{4}$ but $X_{k}\neq Y_{k}$ for $k=0,1,2,3$.
2.2 Pairwise coalescence time: unconditional versus conditional on the pedigree
In our population model each of the $N$ individuals has two copies of each genetic locus. Suppose we number these $J=\{1,...,2N\}$ such that the individual with label $i\in I$ has gene copies $12i-$ 1 and $2i$. Broadly, our aim is to describe the statistical properties of random samples of gene copies from $J$ by following their ancestral genetic lineages backward in time.
Suppose we sample two gene copies $X_{0}$ and $Y_{0}$ without replacement from the $2N$ present at time $k=0$. Let $X_{k}$ and $Y_{k}$ be their ancestral lineages $k$ time-steps in the past. Then $(X_{k})_{k\in\mathbb{Z}_{+}}$ and $(Y_{k})_{k\in\mathbb{Z}_{+}}$ are two correlated Markov chains with state space $J$ that have the same transition probabilities. Our main object of study, for sample size $n=2$, is
$\tau^{(N,2)}:=\inf\{k\in\mathbb{Z}_{+}:\,X_{k}=Y_{k}\},$ (8)
the (pairwise) coalescence time of the sample of two genes. The left image in Figure 2 shows a realization of the population process in which $\tau^{(6,2)}=4$.
The value of $\tau^{(N,2)}$ is completely determined by the population process. However, knowing the pedigree is not enough to determine the exact value of $\tau^{(N,2)}$, due to Mendelian randomness. For instance, even if we know the pedigree is exactly the one the right in Figure 2 and we know that $X_{0}=2$, we do not know what $X_{3}$ is: $X_{3}$ can be in one of the genes $\{5,6,7,8\}$ with equal probability.
3 Unconditional distribution of pairwise coalescence time
[…]
${\text{diff}=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}&\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\bullet}&\end{pmatrix}\,\text{same}=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}&{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\bullet}\end{pmatrix},\ \text{coal}=\begin{pmatrix}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\bullet}&\end{pmatrix}}$
The one-step transition matrix $\mathbf{\Pi}_{N}$ of this Markov process, under our diploid Moran model, is
$\mathbf{\Pi}_{N}\,:=\,$ diff same coal diff $1-\frac{2}{N(N-1)}$ $\frac{1}{N(N-1)}$ $\frac{1}{N(N-1)}$ same $(1-\alpha_{N})\frac{1}{N}$ $\frac{N-1}{N}+\alpha_{N}\frac{1}{2N}$ $\alpha_{N}\frac{1}{2N}$ coal 00 00 $1$
For $q\in\{\text{diff},\,\text{same},\,\text{coal}\}$, we let $\mathbb{P}_{q}$ be the probability under which the initial state (i.e. the sampling configurations) is $q$. In what follows, we use $\,{\buildrel{d}\over{\longrightarrow}}\,$ and =$\,{\buildrel{d}\over{=}}\,$ to denote convergence in distribution and equal in distribution respectively.
Theorem 3.1 (Unconditional limiting distribution).
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ as $N\to\infty$. Then $N^{-2}\tau^{(N,2)}$ converges in distribution:
$N^{-2}\tau^{(N,2)}\,{\buildrel{d}\over{\longrightarrow}}\,\begin{cases}0\qquad&\text{under the law }\;\mathbb{P}_{\rm same}(\cdot\,|\mathcal{O}=0)\\ {\rm Exp}\left(\frac{2}{2-\alpha}\right)\qquad&\text{under the laws }\;\mathbb{P}_{\rm diff}(\cdot)\text{ and }\mathbb{P}_{\rm same}(\cdot\,|\mathcal{O}\geq 1)\end{cases}$
where
$\mathcal{O}:=|\{k\geq 1:\,\hat{X}_{k}=\hat{Y}_{k},\,\hat{X}_{k-1}\neq\hat{Y}_{k-1}\}|$
is the total number of time-steps when the two sample lineages transition from belonging to two distinct individuals to belonging to a single individual, and $\hat{X}_{k},\hat{Y}_{k}$ denote the labels of the individuals to which the genes $X_{k}$ and $Y_{k}$ belong.
The convergence to an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$ in Theorem 3.1, under $\mathbb{P}_{{\rm diff}}$, shows that our Moran model gives the same result as the Wright-Fisher model with partial selfing (Nordborg and Donnelly, 1997; Möhle, 1998) when one averages over the pedigree for a fixed rate of selfing $\alpha\in[0,1]$. Kogan et al. (2023, Lemma 2) recently obtained this result in the broader context of selfing plus recombination under Wright-Fisher reproduction (and averaging over the pedigree).
Two kinds of events are key to our analysis: overlap events where two sample lineages in distinct individuals transition to belonging to the same individual, and splitting events where two sample lineages in the same individual transition to belonging to two distinct individuals. We use $\mathcal{O}$ to denote the number of overlap events. Note that an overlap event results also in coalescence with probability $1/2$. In addition, under $\mathbb{P}_{\rm{same}}$ the two lineages may coalesce before they split, in which case $\mathcal{O}=0$. Under $\mathbb{P}_{{\rm diff}}$ it holds that $\mathcal{O}\geq 1$ almost surely. We prove in Lemma 8.1 that
$\mathbb{P}_{\rm diff}(\mathcal{O}=0)=0\quad\text{and}\quad\mathbb{P}_{\rm same}(\mathcal{O}=0)=\frac{\alpha_{N}}{2-\alpha_{N}}.$ (9)
Remark 3.2.
For any $k\in\mathbb{Z}_{+}$,
$\mathbb{P}_{\rm diff}(\tau^{(N,2)}>k)=$ $\,\left(1,0,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}$
$\mathbb{P}_{\rm same}(\tau^{(N,2)}>k)=$ $\,\left(0,1,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}$
Like its Wright-Fisher counterpart (Nordborg and Donnelly, 1997; Möhle, 1998), Theorem 3.1 is a separation-of-timescales result. Direct application of Möhle (1998, Lemma 1, Theorem 1) or its generalization by Möhle and Notohara (2016) is complicated by the extra factor of $N$ in our diploid Moran model, i.e. to go from time in discrete steps to time in units of $N$ generations. In Section 8.3, we instead use our notion of overlap events to prove convergence.
4 Main results: conditional coalescence times given the pedigree
Our main results are about the conditional distribution of $\tau^{(N,2)}$ given the pedigree and the sampled pair of individuals. We consider a sequence of our diploid Moran model indexed by $N$, under three different regimes for the selfing rate $\alpha_{N}$ in the limit. Our results are summarized in Table 1 and precised in Theorems 4.3 and 4.4. The first corresponds to the model of partial selfing previously studied by others, only excluding cases in which $\alpha_{N}\to 1$ more slowly than $1/N$. The second is what we call limited outcrossing, and happens when $\alpha_{N}\to 1$ at rate $\lambda/N$ for some $\lambda\in(0,\infty)$. The third is when outcrossing is negligible, because $\alpha_{N}\to 1$ faster than $1/N$. As we will show, limited outcrossing interpolates between partial selfing and negligible outcrossing. As $\lambda\to\infty$, it becomes indistinguishable from partial selfing with $\alpha=1$. As $\lambda\to 0$ it becomes indistinguishable from negligible outcrossing.
Following the notation in Diamantidis et al. (2024), we let $\mathcal{A}_{N}$ be the $\sigma$-algebra generated by both the pedigree $\mathcal{G}_{N}$ and the labels $\hat{X}_{0}$ and $\hat{Y}_{0}$ of the sampled individuals. That is, we let
$\mathcal{A}_{N}:=\sigma\left(\mathcal{G}_{N},\,\hat{X}_{0},\,\hat{Y}_{0}\right).$ (10)
Our main results say that the conditional distributions of $N^{-2}\tau^{(N,2)}$ given $\mathcal{A}_{N}$ converge to different distributions, depending on the sampling configuration (whether the two sampled genes are in the same individual or not) and on the three regimes of the selfing probability. Furthermore, the limiting distributions can retain information of the random pedigree as $N\to\infty$.
Before stating our results rigorously, we first give a quick summary in Table 1 and some illustrations in Figure 3. We are interested in the event ($E_{t}$ in Table 1) that the rescaled conditional coalescence time $N^{-2}\tau^{(N,2)}$ for a sample in state $q\in\{\textrm{diff},\textrm{same}\}$ is greater than a fixed time $t$, namely $\mathbb{P}_{\textrm{q}}(N^{-2}\tau^{(N,2)}>t\mid\mathcal{A}_{N})$ which we denote $\mathbb{P}_{{\rm diff}}(E_{t}\,|\,\mathcal{A}_{N})$ and $\mathbb{P}_{\rm{same}}(E_{t}\,|\,\mathcal{A}_{N})$ in Table 1.
regime selfing rate as $N\to\infty$ $\mathbb{P}_{{\rm diff}}(E_{t}\,|\,\mathcal{A}_{N})$ $\mathbb{P}_{\rm{same}}(E_{t}\,|\,\mathcal{A}_{N})$
partial selfing $N(1-\alpha_{N})\rightarrow\infty$ $e^{-\frac{2}{2-\alpha}t}$ $2^{1-\text{Geom}(1-\alpha)}e^{-\frac{2}{2-\alpha}t}$
limited outcrossing $N(1-\alpha_{N})\rightarrow\lambda\in(0,\infty)$ “random staircase” 0
negligible outcrossing $N(1-\alpha_{N})\rightarrow 0$ $\textbf{1}_{\{{\rm Exp}(2)>t\}}$ 0
Table 1: Summary of results for the conditional coalescence time of a sample of size $n=2$, as $N\to\infty$, where $E_{t}$ denotes the event $\{N^{-2}\tau^{(N,2)}>t\}$ for a fixed time $t$. In the last column, $\text{Geom}(1-\alpha)$ is a geometric random variable $K$ satisfying that $\mathbb{P}(K=k)=\alpha^{k-1}(1-\alpha)$ for $k\in\mathbb{Z}_{+}$. In the middle column, $\textbf{1}_{\{{\rm Exp}(2)>t\}}$ is the indicator of the event $\{{\rm Exp}(2)>t\}$, hence it is a Bernoulli random variable $B$ so that $\mathbb{P}(B=1)=e^{-2t}=1-\mathbb{P}(B=0)$.
From Table 1, we see that under the conditional probability $\mathbb{P}_{{\rm diff}}(\cdot\,|\,\mathcal{A}_{N})$, the limiting survival function of $N^{-2}\tau^{(N,2)}$ changes from deterministic to random, as the selfing rate increases from partial selfing to limited outcrossing. In the partial-selfing regime, the deterministic distribution ${\rm Exp}(\frac{2}{2-\alpha})$ is the same as that for the unconditional case (when we average over pedigrees as in Theorem 3.1). In the limited-outcrossing regime, the limiting survival function is a random random function which we call a “random staircase” for reasons which will become clear. In the negligible-outcrossing regime, the limiting survival function is the random function $\textbf{1}_{\{{\rm Exp}(2)>t\}}$. Note that the expectation of $\textbf{1}_{\{{\rm Exp}(2)>t\}}$, namely $e^{-2t}$, is the same as the unconditional probability $\lim_{N\to\infty}\mathbb{P}_{\text{diff}}(N^{-2}\tau^{(N,2)}>t)$ when $\alpha_{N}\to\alpha=1$. This makes sense because
$\mathbb{E}[\mathbb{P}_{\text{diff}}(N^{-2}\tau^{(N,2)}>t\mid\mathcal{A}_{N})]=\mathbb{P}_{\text{diff}}(N^{-2}\tau^{(N,2)}>t)$ (11)
for all $N\in\mathbb{N}$ and $t\in\mathbb{R}_{+}$. It is always true that the expectation of the conditional limit is the same as the corresponding unconditional limit.
Besides, there is a simple explanation for $\textbf{1}_{\{{\rm Exp}(2)>t\}}$ in the negligible-outcrossing regime in Table 1, by considering the extreme case in which $\alpha_{N}=1$ for all $N$. In this extreme case, there is no outcrossing for the two gene copies of any individual, so there is no splitting event for the lineages of the sampled individuals $\hat{X}_{0},\,\hat{Y}_{0}$. The lineages of the sampled individuals $\{\hat{X}_{0},\,\hat{Y}_{0}\}$ will overlap after a random number of time-steps distributed as a geometric random variable with mean $\frac{N(N-1)}{2}$, hence they overlap after approximately $N^{2}T_{0}$ many generations, where $T_{0}$ is an ${\rm Exp(2)}$ random variable. Upon the overlap, the two sampled genes will stay in the same individual(s) and coalesce after $O(N)$ many time-steps, i.e. coalesce immediately under the $N^{2}$ timescale.
Further, in the case of limited outcrossing, with $N(1-\alpha_{N})\rightarrow\lambda\in(0,\infty)$, we will see that our random walks on the pedigree will converge to a particle system where each ancestral lineage splits at rate $\lambda$ and each pair of ancestral lineages coalesce at rate $2$. In particular, no matter which ancestral lineage our two sample particles belong, they coalesce together at rate $2$, which is what we see in the unconditional case.
Under the conditional probability $\mathbb{P}_{\rm{same}}(\cdot\,|\,\mathcal{A}_{N})$, the limiting survival function of $N^{-2}\tau^{(N,2)}$ changes from random to deterministic (namely 0), as the selfing rate increases from partial selfing to limited outcrossing. Note that in the last column of Table 1, the expectation of the conditional probability
$\mathbb{E}\left[2^{1-\text{Geom}(1-\alpha)}e^{-\frac{2}{2-\alpha}t}\right]=\frac{2-2\alpha}{2-\alpha}e^{-\frac{2}{2-\alpha}t}$
is the same as the unconditional probability
$\lim_{N\to\infty}\mathbb{P}_{\text{same}}(N^{-2}\tau^{(N,2)}>t)=\lim_{N\to\infty}\mathbb{P}_{\text{same}}(\mathcal{O}\geq 1)\mathbb{P}_{\text{same}}(N^{-2}\tau^{(N,2)}>t\mid\mathcal{O}\geq 1)$
according to (9). This makes sense in view of (11).
In Figure 3, we illustrate the limiting behavior of the random, conditional cumulative distribution function (conditional CDF) $t\mapsto\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2)}\leq t|\mathcal{A}_{N})$ for our three regimes about $\alpha_{N}$.
Figure 3: The limiting conditional CDF $t\mapsto\lim_{N\to\infty}\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2)}\leq t|\mathcal{A}_{N})$ for our three assumptions about $\alpha_{N}$: partial selfing, limited outcrossing, and negligible outcrossing. (Left) Partial selfing, where $\lim_{N}N(1-\alpha_{N})=\infty$. The limiting CDF is deterministic and is exactly the CDF of an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$; the figure takes $\alpha=1$. (Center) Five realizations of the limiting CDF for $\lambda=5$ under limited outcrossing, i.e. with $\lim_{N}N(1-\alpha_{N})=\lambda=5$. The limiting CDF is random and is described precisely in Theorem 4.3, and the corresponding survival function is what we called a “random staircase” in Table 1; cf. also Figure 1. (Right) Five realizations of the limiting CDF for the case of negligible outcrossing, where $\lim_{N}N(1-\alpha_{N})=0$. The limiting CDF is random, and it is the CDF of a random constant that is exponentially distributed with rate 2, i.e. the CDF is a Heaviside function with exponentially distributed jump time.
To describe the “random staircase” for the case of limited outcrossing with $\lambda\in(0,\infty)$ in Table 1, we first consider the lineage of a single sampled gene from a single individual $\hat{X}_{0}$. We wish to describe a process which encapsulates (i) the set of potential ancestors from whom this present-day gene may have been inherited, and (ii) the corresponding probabilities that each potential ancestor contains the gene. A realization of this process, starting from the individual $\hat{X}_{0}$ and parameterized by time-step $k\in\mathbb{Z}_{+}$ in the past, is illustrated in Figure 4.
Figure 4: A realization of the set of individuals (potential ancestors) of the sampled gene within the individual $\hat{X}_{0}$, the oval at the bottom. The three thick edges depict selfing events. The fractional number in each individual is the probability that the sampled gene lies in that individual. After $3$ time-steps, the individual to which the sample lineage belongs undergoes a fragmentation due to an outcrossing event.
Consider $n=2$ sampled genes, one from each of the individuals $\hat{X}_{0}$ and $\hat{Y}_{0}$. There is a similar description to the set of potential ancestors, but now each potential ancestor has two probabilities, one for each gene. A realization of this process $\{Z^{N}_{k}\}_{k\in\mathbb{Z}_{+}}$ is illustrated in Figure 5.
Figure 5: A realization of the process $\{Z_{k}^{N}\}_{k\in\mathbb{Z}_{+}}$, where $Z_{k}^{N}$ represents the set of potential ancestors at time-step $k$ in the past of the two genes in the sampled individuals $\hat{X}_{0}$ and $\hat{Y}_{0}$, together with the corresponding probabilities (in red and blue respectively) that each potential ancestor contains each of the sample lineages. The individuals at the bottom are $\hat{X}_{0}$ on the left and $\hat{Y}_{0}$ on the right. The initial condition is $Z_{0}=\{((1,0),1),\,((0,1),2)\}$ where the first triple $((1,0),1)$ indicates that $\hat{X}_{0}$ has probabilities/masses $(1,0)$ and label 1, and the second triple $((0,1),2)$ indicates that $\hat{Y}_{0}$ has probabilities/masses $(0,1)$ and label 2.
The discrete-time process $Z^{N}=(Z_{k}^{N})_{k\in\mathbb{Z}_{+}}$ is a fragmentation-coagulation process (Bertoin, 2006) starting with two particles $\{((1,0),1),\,((0,1),2)\}$, in which each individual (a potential ancestor) is viewed as a particle and the pairs of probabilities in the first two components within each potential ancestor are the “masses” of the particle. The third component of the particle is its label.
For each $k\in\mathbb{Z}_{+}$, the value of $Z_{k}^{N}$ is a set of 3-tuples, where the sum of all the first two components of each 3-tuple is equal to $(1,1)$ because the total probability/mass for each of the two lineages must be 1. The third component is the label of the particle, which tracks the total structure of inheritance for the process. The state space $\mathcal{P}_{m,2}$ of $Z^{N}=(Z_{k}^{N})_{k\in\mathbb{Z}_{+}}$ can be represented as
$\mathcal{P}_{m,2}:=\bigcup_{j=1}^{\infty}\mathcal{P}_{m,2}^{(j)},$ (12)
where
$\mathcal{P}_{m,2}^{(j)}:=\{(p_{i},l_{i}):1\leq i\leq j,p_{i}\in[0,1]^{2},l_{i}\in\mathbb{Z}_{+}\}.$
Here $j$ represents the number of particles, and $p_{i}$ and $l_{i}$ represent, respectively, the masses and the label of the $i$-th particle.
Lemma 4.2 says that if one unit of time is $N^{2}$ time-steps and if $N\to\infty$, then under limited outcrossing $Z^{N}$ converges weakly to a continuous-time process described as follows.
Definition 4.1 (Fragmentation-coagulation process with rate $\lambda$).
Let $\lambda\in\mathbb{R}_{+}$ be a fixed deterministic number. A fragmentation-coagulation process with rate $\lambda$ is a continuous-time process taking values in $\mathcal{P}_{m,2}$ and has right-continuous sample paths almost surely, and such that
  • (i)
    Each particle fragments independently at rate $\lambda$ into two particles with equal masses which is half of that of the fragmenting particle, and each pair of particles coagulate independently at rate $2$ into a particle whose masses are the sum of those of the two coagulating particles.
  • (ii)
    If $i$ is the label of the fragmenting particle, then the labels of the two resultant particles are $i$ and $m$, where $m$ is the smallest positive integer such that all labels are distinct; if $i$ and $j$ are the labels of the two coagulating particles, then the resultant particle has label $\min\{i,j\}$.
Let $Z_{\lambda}=(Z_{\lambda}(t))_{t\in\mathbb{Z}_{+}}$ be a fragmentation-coagulation process with rate $\lambda$, starting with two immiscible unit masses $Z_{\lambda}(0)=\{((1,0),1),\,((0,1),2)\}$; i.e. it starts with two particles with masses $(1,0)$ and $(0,1)$ and with labels 1 and 2 respectively.
The terminology in Definition 4.1 is inspired by Bertoin (2006) who discusses fragmentation and coagulation processes of unit masses. The use of the word “immiscible" reflects the fact that the coagulation of two particles with masses $(p,q)$ and $(r,s)$ into a particle with masses $(p+r,q+s)$ results in no crossover between components; i.e., the two components of the masses do not mix.
Lemma 4.2.
Suppose $\lim_{N\to\infty}N(1-\alpha_{N})=\lambda\in\mathbb{R}_{+}$. Then $(Z_{\lfloor tN^{2}\rfloor}^{N})_{t\in\mathbb{R}_{+}}$ converges in distribution in the Skorokhod space $\mathcal{D}(\mathbb{R}_{+},\mathcal{P}_{m,2})$ under $\mathbb{P}_{{\rm diff}}$ to the process $Z_{\lambda}$ in Definition 4.1.
Lemma 4.2 is proven in Section 6.1. If we ignore the masses, e.g. ignoring the pairs of probabilities in the individuals in Figure 5, then we obtain a simpler process $G^{N}=(G_{k}^{N})_{k\in\mathbb{Z}_{+}}$ from $Z^{N}$ that keeps track of the set of potential ancestors and their ancestral relationships (i.e. which splits into, and which two particles coagulate). It is the particle process consisting solely of the labels of the particles of $Z^{N}$. This process $G^{N}$ has a graphical structure that represents all possible trajectories of $(\hat{X}_{k},\hat{Y}_{k})_{k\in\mathbb{Z}_{+}}$ (recall their definition in Theorem 3.1) and can be viewed as a subset of the population pedigree $\mathcal{G}_{N}$.
Similarly, ignoring the masses in $Z_{\lambda}$, we obtain a process which we denote by $G_{\lambda}=(G_{\lambda}(t))_{t\in\mathbb{R}_{+}}$. This continuous-time process is a particle system starting with two different particles, where each particle splits into two with rate $\lambda$ and each pair of particles coagulate independently at rate $2$. The labelling process of the particles is retained from the process $Z_{\lambda}$. It is equal in distribution to an ancestral recombination graph (Griffiths, 1991; Griffiths and Marjoram, 1997) or an ancestral selection graph (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). Throughout this paper, we call $G_{\lambda}$ an ancestral graph. A realization of an ancestral graph is presented on the left in Figure 6, and further examples are offered in Figure 7.
Figure 6: (Left) A realization $G_{\lambda}$ of the random ancestral graph $G_{\lambda}$ starting with two particles/nodes. For this particular realization, the overlap times of the sample lineages are $t_{1}<t_{2}<t_{3}$. (Right) The “random staircase” mentioned in Table 1 corresponding to this realization of $G_{\lambda}$. The conditional distribution of the limiting coalescence time $T_{\lambda}$, given $G_{\lambda}$, is obtained by tracing ancestral genetic lineages backwards in time along the graph $G_{\lambda}$. Given this $G_{\lambda}$, $T_{\lambda}$ must take values in $\{t_{1},t_{2},t_{3}\}$. One can read off the conditional CDF of $T_{\lambda}$ as follows: The node of $G_{\lambda}$ at time $t_{1}$ contains $\frac{3}{4}$ of the genetics of the right lineage and the whole of that of the left lineage, so $\mathbb{P}(T_{\lambda}=t_{1}|G_{\lambda})=\frac{3}{4}$. Between $t_{1}$ and $t_{2}$, half of the left lineage splits into the left handle. At $t_{2}$, $\frac{1}{4}$ of the right lineage meets half of the left lineage, so $\mathbb{P}(T=t_{2}|G_{\lambda})=\frac{1}{8}$. The remainder of the right lineage meets the remainder of the left lineage at $t_{3}$ so $\mathbb{P}(T_{\lambda}=t_{3})=\frac{1}{8}$.
Next, we consider two conditionally independent random walks on a given ancestral graph. Given $G_{\lambda}=(G_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and let $G_{\lambda}(0)=\{x(0),\,y(0)\}$ be the two different particles at $t=0$. We let $(x_{\lambda}(t),\,y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ be a continuous-time process starting at $(x(0),\,y(0))$, such that
  • (i)
    $\{x_{\lambda}(t),y_{\lambda}(t)\}\subset G_{\lambda}(t)$ for all $t\in\mathbb{R}_{+}$,
  • (ii)
    at any fragmentation event, each of $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ will follow each of the two paths available with equal (i.e. 1/2) probability.
In particular, $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ are conditionally independent random walks on the ancestral graph $G_{\lambda}$ starting at two different particles. The first meeting time of these random walks is
$T_{\lambda}:=\inf\{t\in\mathbb{R}_{+}:\,x_{\lambda}(t)=y_{\lambda}(t)\}.$ (13)
By Theorem 4.3, $T_{\lambda}$ under the law of $\mathbb{P}(\cdot|G_{\lambda})$ is the limit law of the pairwise coalescence time of two sample lineages given the pedigree under limited outcrossing. In Figure 7 below, we illustrate two different realizations of the process $(x_{\lambda},y_{\lambda})$ on the same ancestral graph $G_{\lambda}$.
Figure 7: Two realizations of the process $(x_{\lambda},y_{\lambda})$ on the same ancestral graph $G_{\lambda}$, one on the left and the other on right. In each realization, $x_{\lambda}$ is in blue and $y_{\lambda}$ in red. At each fragmentation of the graph, where a particle splits into two, each walk chooses which particle path to follow fairly i.e. with probability $\frac{1}{2}$ each. The thick purple dot represents the first meeting time $T_{\lambda}$ of the two paths $x_{\lambda}$ and $y_{\lambda}$. For these two realizations, $T_{\lambda}=t_{1}$ on the left and $T_{\lambda}=t_{2}$ on the right, where $\{t_{1},t_{2},t_{3}\}$ are shown in Figure 6,
4.1 Rigorous statements of convergence results
We are now ready to state our main results for coalescence times in Theorems 4.3 and 4.4. Recall the sigma field $\mathcal{A}_{N}$ in (10) and the pairwise coalescence time $\tau^{(N,2)}$ defined in (8).
Theorem 4.3.
For any $t\in\mathbb{R}_{+}$, as $N\to\infty$,
$\mathbb{P}_{{\rm diff}}\left(N^{-2}\tau^{(N,2)}>t\,|\,\mathcal{A}_{N}\right)\,{\buildrel{d}\over{\longrightarrow}}\,\begin{dcases}e^{-\frac{2}{2-\alpha}t}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\infty\;\text{ and }\,\alpha_{N}\rightarrow\alpha\in[0,1]\\ \mathbb{P}(T_{\lambda}>t\,|\,G_{\lambda})&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\\ \mathbbm{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow 0\end{dcases}.$
Theorem 4.3 offers a description of the conditional distribution of the coalescence time $\tau^{(N,2)}$ for a sample of two gene copies from different individuals in a population of size $N$ given the pedigree. Simulations of the random CDFs can be seen in Figure 3. It says that the law of $N^{-2}\tau^{(N,2)}$ for two lineages in different individuals is robust to the structure of the pedigree so long as the rate at which any individual undergoes a splitting $(1-\alpha_{N})N^{-1}$ is significantly faster than the rate at which any two sample lineages in distinct individuals coalesce, $N^{-2}+O(N^{-3})$, in other words that the ratio of the former over the latter tends to infinity. This robustness for the case of partial selfing is simply that the conditional limit agrees exactly with that in Theorem 3.1. If, on the other hand, the rate of coalescence is comparable to or dominates the rate of splitting of an individual, as in the cases of limited outcrossing and negligible outcrossing, the theorem demonstrates the non-robustness of the pairwise coalescence time to the pedigree. That is, the conditional pairwise coalescence time, more precisely its survival function or equivalently its CDF, does not converge to its mean.
From a technical point of view, Theorem 4.3 makes precise how the random probability measure $\xi_{N}:=\mathbb{P}_{{\rm diff}}\left(N^{-2}\tau^{(N,2)}\in\cdot\,|\,\mathcal{A}_{N}\right)$ converges, as $N\to\infty$, to a $\mathcal{P}(\mathbb{R}_{+})$-valued random variable $\xi$ which happens to be deterministic in the partial-selfing case and random in the other two cases. The random variable $\xi$ is the law of an ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$ random variable in the first case, the conditional law of $T_{\lambda}$ given $G_{\lambda}$ in the second case, and the (random) law $\textbf{1}_{\{{\rm Exp}(2)\in\cdot\}}$ in the third case. We anticipate that, with additional effort, one could justify the above convergence in the sense of weak convergence in the space $\mathcal{P}(\mathbb{R}_{+})$ of probability measures on $\mathbb{R}_{+}$.
Analogously to Theorem 4.3, we obtain the limiting conditional distribution of the coalescence time when our samples are taken from the same individual.
Theorem 4.4.
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ as $N\to\infty$. For all $t\in\mathbb{R}_{+}$ we have convergence in distribution
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})\,{\buildrel{d}\over{\longrightarrow}}\,\begin{dcases}e^{-2t}&\quad\text{if }\alpha=0\\ 2^{-U}e^{-2t}&\quad\text{if }\alpha\in(0,1)\\ 0&\quad\text{if }\alpha=1\end{dcases},$
where $U$ is a random variable satisfying $\mathbb{P}(U=k)=\alpha^{k}(1-\alpha)$ for $k\in\mathbb{Z}_{+}$.
Remark 4.5.
As in (2), the quantity $U$ represents the number of selfing events in the ancestry of the individual from whom our two sample lineages are taken, before the first splitting or outcrossing event. It is also one less than the geometric random variable $K$ introduced in Table 1, which was the number of generations back to the first outcrossing event in ancestry of the individual.
4.2 Conditional genealogy for general sample size $n>2$
We anticipate that our main results can readily be extended to general sample size $n>2$. We do not provide details of the proofs in this paper and leave it as a conjecture in this section and for future work.
Suppose we have a sample of $n$ genes from the population at time-step 00, where $n>2$ is an arbitrary fixed number. It is customary to summarize the ancestral relationships in the sample by an ancestral process $(\Pi_{k}^{N,n})_{k\in\mathbb{Z}_{+}}$ takes values in the space $\mathcal{E}_{n}$ of partitions of $\{1,\ldots,n\}$, where indices belong to the same block of $\Pi^{N,n}_{k}$ if and only if the samples with these indices have a common ancestor $k$ steps in the past.
As before, we let $\mathbb{P}_{{\rm diff}}$ be the probability under which our $n$ samples coming from $n$ distinct individuals. Under $\mathbb{P}_{{\rm diff}}$, the initial state $\Pi_{0}^{N,n}$ of the ancestral process has $n$ blocks of singleton. We also denote by $\mathcal{A}_{N}$ the $\sigma$-algebra generated by both the pedigree and the labels of the $n$ individuals from whom we sample the $n$ sample lineages. We make the following conjecture.
Conjecture 4.6.
Under the conditional law $\mathbb{P}_{{\rm diff}}(\cdot|\mathcal{A}_{N})$, the sequence of $\mathcal{E}_{n}$-valued processes $\Pi^{N,n}=(\Pi^{N,n}_{\lfloor tN^{2}\rfloor})_{t\in\mathbb{R}_{+}}$ converges in finite dimensional distributions, as $N\to\infty$, to a random variable
$\Pi=\begin{cases}\mathcal{K}^{(\alpha)}_{n}&\quad\text{if }N(1-\alpha_{N})\rightarrow\infty\quad\text{and }\alpha_{N}\rightarrow\alpha\\ P_{\lambda}\quad\text{under }\mathbb{P}(\cdot|G_{\lambda}^{n})&\quad\text{if }N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\end{cases}$ (14)
where $\mathcal{K}^{(\alpha)}_{n}$ is the standard Kingman $n$-coalescent with time rescaled by the factor $\frac{2}{2-\alpha}$. and $K^{(1)}_{n}$ is a single realization of $\mathcal{K}^{(1)}_{n}$.
Here we let $G_{\lambda}^{n}$ be an ancestral graph starting with $n$ initial nodes. Given $G_{\lambda}^{n}$. We let $G_{\lambda}^{n}(0)=\{x_{i}(0)\}_{1\leq i\leq n}$ be the set of the $n$ particles at $t=0$, and we let $((x_{i}(t))_{1\leq i\leq n})_{t\in\mathbb{R}_{+}}$ be the system of coalescing random walks on $G_{\lambda}^{n}$, starting at $(x_{i}(0))_{1\leq i\leq n}$ defined as before; before equation (13). Finally, $P_{\lambda}=(P_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ is a $\mathcal{E}_{n}$-valued process defined by the coalescence of the particles i.e. $i\sim_{P_{\lambda}(t)}j$ if and only if $x_{i}(t)=x_{j}(t)$.
Furthermore, as in Proposition 22, we expect that as $\lambda\rightarrow\infty$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to the Kingman $n$-coalescent $\mathcal{K}^{(\alpha)}_{n}$; and as $\lambda\rightarrow 0$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to a single realization of $\mathcal{K}^{(1)}_{n}$, which is the limit in the negligible-outcrossing regime.
5 Properties of the scaling limit under limited outcrossing
Here we provide some additional analyses of the novel scaling limits in Theorems 4.3 and 4.4, especially the coalescing random walk on the ancestral graph, motivated by their possible utility for a larger class of models than just our Moran model. For example, the simulations for the Wright-Fisher model in Figure 1 shows similar behavior as that for our Moran model in Figure 3.
Observe first that the limited-outcrossing regime in Theorem 4.3 includes the case $\lambda=0$, which means that the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ (when $\lambda=0$) is equal to that of $H$ in the negligible-outcrossing regime, where $H$ is a constant random variable whose value is exponentially distributed with rate 2. Corollary 22 below covers the other extreme: as $\lambda\to\infty$, the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ converges to ${\rm Exp}(2)$, which corresponds to the partial-selfing regime with $\alpha=1$. Hence, for lineages in different individuals, the limited-outcrossing regime covers all possible cases for which $\alpha_{N}\to 1$, as we tune the magnitude of $\lambda$. This highlights the significance of the limited-outcrossing regime.
We focus our analysis on the variance of the conditional survival probability $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ (for all $t\in\mathbb{R}_{>0}$) and that of the conditional expectation $\mathbb{E}\left[T_{\lambda}|G_{\lambda}\right]$. We obtain explicit formulas and their asymptotics (as $\lambda\to 0$ and $\lambda\to\infty$). The starting point of our analysis is a connection to a pair of two conditionally independent gene genealogies of the sample, in the spirit of Diamantidis et al. (2024) and Birkner et al. (2013).
Lemma 5.1.
For any $\lambda\in\mathbb{R}_{>0}$ and $t\in\mathbb{R}_{>0}$,
$\mathrm{Var}{\left(\mathbb{P}(T_{\lambda}>t|G_{\lambda})\right)}=\mathbb{P}(T_{\lambda}\wedge T_{\lambda}^{\prime}>t)-e^{-4t}\quad\text{ and }\quad\mathrm{Var}{\left(\mathbb{E}\left[T_{\lambda}|G_{\lambda}\right]\right)}=\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)},$ (15)
where $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are two conditionally independent copies of the pairwise hitting time given the same $G_{\lambda}$; they are defined in (13) with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$ and $x_{\lambda}(0)\neq y_{\lambda}(0)$.
Proof.
We proof only the first equality in (15) which is equivalent to $\mathbb{E}[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}]=\mathbb{P}(T_{\lambda}\wedge T_{\lambda}^{\prime}>t)$. The second follows from the same argument.
$\mathbb{E}\left[\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t|G_{\lambda})^{2}\right]=$ $\,\mathbb{E}\left[\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t|G_{\lambda})\;\mathbb{P}_{x_{\lambda}^{\prime}(0)\neq y_{\lambda}^{\prime}(0)}(T^{\prime}_{\lambda}>t|G_{\lambda})\right]$
$=$ $\,\mathbb{E}\left[\mathbb{P}_{x_{\lambda}^{\prime}(0)=x_{\lambda}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)}(T_{\lambda}>t,\;T^{\prime}_{\lambda}>t|G_{\lambda})\right]$
$=$ $\,\mathbb{P}_{x_{\lambda}^{\prime}(0)=x_{\lambda}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)}(T_{\lambda}\wedge T^{\prime}_{\lambda}>t),$
where the second equality above follows from the conditional independence of $T_{\lambda}$ and $T_{\lambda}^{\prime}$. ∎
Recall from Lemma 5.1 that the random variables $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are the pairwise meeting times for two conditionally independent processes $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ with the same starting point, i.e. with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$. Each of these two processes represents the dynamics of a pair of particles on the ancestral graph $G_{\lambda}$. Note also that in Lemma 5.1 we required $x_{\lambda}\neq x_{\lambda}^{\prime}$, meaning that $T_{\lambda}$ and $T_{\lambda}^{\prime}$ apply to the case in which two individuals are sampled and two copies of the pairwise coalescence process are followed starting at these same two individuals.
In order to model the full two-pairs process, we need to consider all possible samples and all possible ancestral states of the samples. For simplicity, we combine states where possible by symmetry and restrict ourselves to a process with the following five states.
$s_{0}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{1}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{2}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{\Delta,1}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{\Delta,2}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
Here, two dots of the same color (red or blue) correspond to two particles of the same pair; say, the two red dots correspond to $(x_{\lambda},y_{\lambda})$ and the two blue dots correspond to $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$. Two dots within a parenthesis correspond to taking the same value in $G_{\lambda}$, i.e. being in the same individual. Hence, for $i\in\{0,1,2\}$, the state $s_{i}$ corresponds to when $i$ individuals or particles in $G_{\lambda}$ contain two conditionally independent lineages.
[…]
Lemma 5.2.
Let $S=(S_{t})_{t\in\mathbb{R}_{+}}$ be the process, with state space $\{s_{0},\,s_{1},\,s_{2},\,s_{\Delta,1},\,s_{\Delta,2}\}$, which tracks the two pairs $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ at time $t$ (backward) on the coalescence time-scale, then its rate matrix $Q_{\lambda}$ is
$Q_{\lambda}=\begin{pmatrix}-12&8&0&4&0\\ \frac{\lambda}{2}&-6-\frac{\lambda}{2}&2&4&0\\ 0&\lambda&-2-\lambda&0&2\\ 0&0&0&-2&2\\ 0&0&0&0&0\end{pmatrix}$ (16)
The proof of Lemma 5.2 is detailed in the Appendix. Observed that $Q_{\lambda}$ is equivalent to the corresponding matrices for the Kingman coalescent with recombination between two loci (Kaplan and Hudson, 1985; Pluzhnikov and Donnelly, 1996; Simonsen and Churchill, 1997); see also Wakeley (2009, Chapter 7) and Durrett (2008, Chapter 3). More precisely, our model here becomes that in Wakeley (2009, Figure 7.7) if we divide all our rates by two, and then make the substitution $\lambda/2\to\rho$. This $1/2$ on $\lambda$ comes from the fact that each splitting event is only observed with probability $1/2$.
Then, the covariances of $T_{\lambda}$ and $T_{\lambda}^{\prime}$ for the different sample configurations are exactly the same as the classical formulas in two-locus models with recombination, given for example by equations (7.30), (7.29) and (7.28) in Wakeley (2009). Thus we have
$\mathrm{Cov}_{s_{0}}{\left(T_{1},T_{2}\right)}$ $=\frac{4}{\lambda^{2}+26\lambda+72}$ (17)
$\mathrm{Cov}_{s_{1}}{\left(T_{1},T_{2}\right)}$ $=\frac{6}{\lambda^{2}+26\lambda+72}$ (18)
$\mathrm{Cov}_{s_{2}}{\left(T_{1},T_{2}\right)}$ $=\frac{1}{2}\frac{\lambda+36}{\lambda^{2}+26\lambda+72}.$ (19)
Proposition 5.3.
For any $\lambda\in\mathbb{R}_{+}$, the variance of $\mathbb{E}{\left[T_{\lambda}|G_{\lambda}\right]}$ is equal to
$\mathrm{Var}{\left(\mathbb{E}{\left[T_{\lambda}|G_{\lambda}\right]}\right)}=\mathrm{Cov}{\left(T_{\lambda},T^{\prime}_{\lambda}\right)}=\frac{1}{2}\frac{\lambda+36}{\lambda^{2}+26\lambda+72}.$ (20)
In particular,
$\mathrm{Var}{\left(\mathbb{E}{\left[T_{\lambda}|G_{\lambda}\right]}\right)}=\begin{dcases}\frac{1}{4}-\frac{1}{12}\lambda+O(\lambda^{2})&\text{ as }\lambda\to 0\\[4.0pt] \frac{1}{2\lambda}+\frac{5}{\lambda^{2}}+O(\lambda^{-3})&\text{ as }\lambda\to\infty\end{dcases}.$
Proof.
The first equality follows directly from (15). Since $(T_{\lambda},T_{\lambda}^{\prime})=(T_{1},T_{2})$ given state $s_{2}$, (19) is precisely the second equality in (20). ∎
In the Discussion, we will consider the significance of (20) for an understanding how coalescence times depend on the pedigree, represented in this case by $G_{\lambda}$.
Next, we consider the variance of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$. We shall compute the second moment of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ in Lemma 5.7 and use it to show the following.
Proposition 5.4.
For any $t\in\mathbb{R}_{>0}$,
$\mathrm{Var}\left(\mathbb{P}(T_{\lambda}>t|G_{\lambda})\right)=\begin{dcases}e^{-2t}-e^{-4t}+\lambda\frac{e^{-2t}}{8}\left(1-4t-e^{-4t}\right)+O(\lambda^{2})&\quad\text{ as }\lambda\to 0\\ \frac{2e^{-4t}}{\lambda}\left(1+\frac{2(3+16t)}{\lambda}\right)+O(\lambda^{-3})&\quad\text{ as }\lambda\to\infty\end{dcases}.$ (21)
In particular, for any $t\in\mathbb{R}_{>0}$,
$\mathbb{P}(T_{\lambda}>t|G_{\lambda})\,{\buildrel{d}\over{\longrightarrow}}\,\begin{dcases}\textbf{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{ as }\lambda\to 0\\ e^{-2t}&\quad\text{ as }\lambda\to\infty\end{dcases}$ (22)
Lemma 5.7 and Proposition 5.4 will be proved in Section 5.1.
We observe that the covariance $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is connected to the error between the conditional survival function $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ and the unconditional survival function $\mathbb{P}(T_{\lambda}>t)=e^{-2t}$.
Proposition 5.5.
For any $\lambda\in\mathbb{R}_{>0}$,
[…]
Proof.
By Lemma 5.1, the LHS is equal to
[…]
where by $T_{\lambda}\wedge T_{\lambda}^{\prime}$ we mean the smaller of $T_{\lambda}$ and $T_{\lambda}^{\prime}$. ∎
Remark 5.6.
In agreement with previous work—see equations (3) and (52) in Simonsen and Churchill (1997)—from (19) and the fact $\mathrm{Var}{\left(T_{\lambda}\right)}=1/4$, we see that the correlation coefficient $\mathrm{Corr}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is exactly the same as $\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$. Equivalently, $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}=\frac{1}{4}\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$, and it is also equal to the left hand side of (23).
Depending on $\lambda\in(0,\infty)$, limited outcrossing displays a range of behaviors between the two qualitatively different extremes of negligible outcrossing and partial selfing. Figure 8 illustrates this using the law of total variance for the coalescence time of two gene copies sampled from two different individuals. Let $T$ and $G$ represent this “diff” coalescence time and its associated ancestral graph, for simplicity dropping the subscript $\lambda$ we attached to these previously in Section 5. Then we can write
$\textrm{Var}{\left(T\right)}=\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}+\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ (24)
and note that $\textrm{Var}{\left(T\right)}=1/4$, because $T$ by itself is exponentially distributed with rate parameter $2/(2-1)=2$. The second term on the right-hand side of (24), $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, is given by (20). The first term on the right-hand side of (24) is just $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}=1/4-\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, but both terms are displayed in Figure 8 for the sake of illustration.
Figure 8: The two components, $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$ and $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, of $\textrm{Var}{\left(T\right)}$ as functions of $\lambda$ under limited outcrossing. By the law of total variance, $\textrm{Var}{\left(T\right)}=\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}+\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$. Also, $\textrm{Var}{\left(T\right)}=1/4$ in this case because it is the average over ancestral graphs ($G$), that is to say over pedigrees, and because $\alpha_{N}\to 1$.
When $\lambda$ is very small, on the far left in Figure 8, variation in the mean coalescence time among graphs, $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, accounts for the bulk of variation in $T$. Correspondingly, there is little variation in coalescence times given the graph. Figure 8 displays the expectation of the latter, $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$, but $\textrm{Var}{\left(T|G\right)}$ itself will be small on any graph in this case. With probability approaching one as $\lambda\to 0$, the individuals containing the two ancestral lineages will experience only selfing events in their ancestries, up to and including the time they descend from a common ancestral individual. Pairwise coalescence times at every locus in the genome starting from these two individuals will converge on the time of this ancestor and $\textrm{Var}{\left(T|G\right)}$ should be minimal. Variation in the waiting time to this first common ancestral individual among graphs will be the primary source of variation in $T$. In the limit $\lambda\to 0$, this waiting time is exponentially distributed with rate parameter $2$, so $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ approaches $\textrm{Var}{\left(T\right)}=1/4$. It is given by a Kingman coalescent process made effectively haploid by selfing, just as under negligible outcrossing.
Across the middle of Figure 8, ancestral graphs will include increasing numbers of splitting or outcrossing events as $\lambda$ increases. Such graphs contain multiple pathways to coalescence because genetic lineages may trace back to either parent at each outcrossing event. They include multiple possible coalescence times, as already evident in the CDFs of $T|G$ in Figure 6 for a hypothetical graph with three splitting events and three possible coalescence times, in Figure 3 for each of five simulated ancestral graphs with $\lambda=2$, and in Figure 1 for each of fifty pedigrees in three versions of a finite-$N$ Wright-Fisher model corresponding to $\lambda\in\{1,10,100\}$. As a result, $\textrm{Var}{\left(T|G\right)}$ is positive and $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$ becomes the main source of variation in $T$ as $\lambda$ increases, eventually swamping variation of the mean coalescence times among graphs.
When $\lambda$ is very large, on the far right in Figure 8, single ancestral graphs do little to constrain $T|G$. These graphs are large. In their construction backward in time, the number of ancestral lineages grows quickly to reach a quasi-stable distribution, specifically a zero-truncated Poisson distribution with mean $\lambda$, as Mano (2009) showed for the ancestral recombination graph. The waiting time to the event that completes the graph, when there is just one lineage left, will be long, $\sim 2e^{\lambda}/\lambda^{2}$ in expectation using equation (1.1) in Griffiths and Marjoram (1997). Given the graph, each of the pair of genetic lineages starting from the two sampled individuals will trace back through a large number of splitting events before the two meet in a single individual and coalesce. As $\lambda\to\infty$, all variation in $T$ will be due to this random process of coalescence given the graph. Corollary 22 shows that this conditional coalescent process converges on that of the partial-selfing model with $\alpha_{N}\to 1$. Proposition 5.5 shows further that $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ in (20) is a key measure (in $L^{2}$) of the distance of the CDF under limited outcrossing to that under the corresponding model of partial selfing.
To summarize what is shown in Figure 8, limited outcrossing becomes like negligible outcrossing as $\lambda\to 0$ and like partial selfing as $\lambda\to\infty$. With respect to the pre-limiting Moran model, the results and intuitions about Figure 8 concern the behavior at the boundary of complete or nearly complete selfing ($\alpha_{N}\to 1$). All at this boundary, we have: first negligible outcrossing, then limited outcrossing, then an extreme case of partial selfing with $\alpha=1$. The entire rest of the range of selfing rates $\alpha\in[0,1]$ is covered by the partial-selfing model.
5.1 Proof for Proposition 5.4
Observe that the two pairs of particles $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ absorbed at $T_{\lambda}\wedge T_{\lambda}^{\prime}$ is a continuous-time Markov chain, with state space $\{s_{0},s_{1},s_{2},s_{\Delta}\}$, where we collapsed $s_{\Delta,2}$ and $s_{\Delta,2}$ into a single state called $s_{\Delta}$. This absorbed process, called $\widetilde{S}=(\widetilde{S}_{t})_{t\in\mathbb{R}_{+}}$, has 4 states and has transition rate matrix $R_{\lambda}$ obtained from (16) by collapsing $s_{\Delta,2}$ and $s_{\Delta,2}$ into a single state $s_{\Delta}$. By Lemma 5.2, the process $\widetilde{S}$ has rate matrix
$R_{\lambda}=\begin{pmatrix}A_{\lambda}&{\bf v}^{\,\rm T}\\ {\bf 0}&0\end{pmatrix},$ (25)
where $A_{\lambda}$ is the matrix (27) and ${\bf v}=\begin{pmatrix}4&4&2\end{pmatrix}$. Note that
$T_{\lambda}\wedge T_{\lambda}^{\prime}=\inf\{t\in\mathbb{R}_{+}:\,S_{t}\in\{s_{\Delta,1},s_{\Delta,2}\}\}=\inf\{t\in\mathbb{R}_{+}:\,\widetilde{S}_{t}=s_{\Delta}\}$ (26)
is the first time the process $\widetilde{S}$ reaches $s_{\Delta}$. The joint distribution of $T_{\lambda}$ and $T_{\lambda}^{\prime}$ can be computed as in (Diamantidis et al., 2024, Lemma A1) but we omit this here since it is not needed.
We now compute the second moment (and hence also the variance) of the conditional survival probability $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$.
Lemma 5.7.
For any $\lambda\in\mathbb{R}_{>0}$ and $t\in\mathbb{R}_{>0}$, $\mathbb{E}[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}]$ is equal to the sum of the entries of the last row of the 3 by 3 matrix $e^{tA_{\lambda}}$, where
$A_{\lambda}=\begin{pmatrix}-12&8&0\\ \frac{\lambda}{2}&-6-\frac{\lambda}{2}&2\\ 0&\lambda&-2-\lambda\end{pmatrix}.$ (27)
Proof of Lemma 5.7. We first show that
$\mathbb{E}[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}]=\begin{pmatrix}0&0&1&0\end{pmatrix}e^{tR_{\lambda}}\begin{pmatrix}1&1&1&0\end{pmatrix}^{T}.$ (28)
Recall the absorbed chain $\widetilde{S}_{\lambda}=(\widetilde{S}_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ mentioned earlier in this section. Since $x_{\lambda}(0)=x_{\lambda}^{\prime}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)$, we have that the initial condition $\widetilde{S}_{\lambda}(0)=s_{2}$. Then for any $t\in\mathbb{R}_{+}$, by (15),
$\mathbb{E}\left[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}\right]=$ $\,\mathbb{P}(T_{\lambda}\wedge T^{\prime}_{\lambda}>t)$
$=$ $\,\mathbb{P}_{s_{2}}(\widetilde{S}_{\lambda}(t)\in\{s_{0},s_{1},s_{2}\})$
$=$ $\,(0,0,1,0)\,e^{tR_{\lambda}}\,(1,1,1,0)^{T},$
where the second last equality follows from (26). Hence (28) is established. The right hand side of (28) is the sum of the entries of the last row of the 3 by 3 matrix $e^{tA_{\lambda}}$. Hence Lemma 5.7 follows. ∎
Proof of Proposition 5.4. We will employ equation (28) in the proof of Lemma 5.7. We begin by establishing the rate of convergence as $\lambda\to 0$. We write $R_{\lambda}=C+D\lambda$, where
$C:=\begin{pmatrix}-12&8&0&4\\ 0&-6&2&4\\ 0&0&-2&2\\ 0&0&0&0\end{pmatrix}\quad\text{and}\quad D:=\begin{pmatrix}0&0&0&0\\ \frac{1}{2}&-\frac{1}{2}&0&0\\ 0&1&-1&0\\ 0&0&0&0\end{pmatrix}.$ (29)
Then, using for instance Higham (2008, eqn. 10.40),
[…]
By applying the derivative in $\lambda$ at $\lambda=0$ to the Suzuki-Trotter identity, e.g. Higham (2008, eqn. 10.9), for the matrix exponential we simultaneously obtain that
[…]
Combining (31) and (30) gives
$\frac{de^{tR_{\lambda}}}{d\lambda}\Bigg{|}_{\lambda=0}=\begin{pmatrix}I,0\end{pmatrix}\ \exp{\left(t\begin{pmatrix}C&D\\ 0&C\end{pmatrix}\right)}\begin{pmatrix}0\\ I\end{pmatrix}.$
By a Wolfram Research, Inc. (2024, Mathematica) calculation, detailed in Newman (2024),
$(0,0,1,0)\begin{pmatrix}I&0\end{pmatrix}\exp\left(t\begin{pmatrix}C&D\\ 0&C\end{pmatrix}\right)\begin{pmatrix}0\\ I\end{pmatrix}(1,1,1,0)^{T}=\frac{e^{-2t}}{8}\left(1-4t-e^{-4t}\right)$ (32)
The result for $\lambda\to 0$ thus follows by Taylor’s theorem for matrix functions and (28).
The proof for $\lambda\to\infty$ is done via a Wolfram Research, Inc. (2024) calculation in Newman (2024). The proof of Proposition 5.4 is complete. ∎
6 Proofs for Theorem 4.3 and Theorem 4.4
In subsections 6.1-6.3 below, we consider convergences for each of the three regimes. The proofs for Theorems 4.3 and 4.4 will then be summarized in subsection 6.4.
6.1 Proofs for the limited-outcrossing regime
In this section, we prove our main results for the limited-outcrossing regime, in which $N(1-\alpha_{N})\to\lambda\in\mathbb{R}_{+}$.
We first describe a discrete-time Markov process $\tilde{Z}^{N}$, for each integer $N>1$, which not only keeps track of the labels of the potential ancestors of our sample, but also yields the fragmentation-coagulation process $Z^{N}$ described before Definition 4.1 via projection. This process takes values in finite subsets of $[0,1]^{2}\times\mathbb{Z}_{+}\times\mathbb{Z}_{+}$ and has initial state
$\tilde{Z}_{0}^{N}=\{((1,0),1,\hat{X}_{0}),((0,1),2,\hat{Y}_{0})\}.$
The third component in each set tracks a new label that we will create and the fourth component corresponds to the label given by the population. For simplicity we denote by $P(k)$ the collection of possible progenitors in the past, i.e. all those elements $i$ in $I_{N}$ such that $i$ is the fourth component of some 4-tuple in $\tilde{Z}_{k}^{N}$. A new label will be created whenever the fourth component of one of the four-tuples undergoes a split where neither parent belongs to $P$.
To understand the transition of this process in a single time-step, we first demonstrate its transition from time-step $k=0$ to time-step $k=1$. Precisely, $\tilde{Z}_{1}^{N}$ is equal to
$\tilde{Z}_{1}^{N}=\begin{cases}\{((1,0),1,\hat{X}_{0}),((0,1),2,\hat{Y}_{0})\},&\text{if neither $\hat{X}_{0},\hat{Y}_{0}$ is $\kappa_{0}$}\\ \{((1,0),1,\pi_{0,1}),((0,1),2,\hat{Y}_{0})\},&\text{if $\kappa_{0}=\hat{X}_{0}$ self-fertilizes and $\pi_{0,1}\neq\hat{Y}_{0}$}\\ \{((1,0),1,\hat{X}_{0}),((0,1),2,\pi_{0,1})\},&\text{if $\kappa_{0}=\hat{Y}_{0}$ self-fertilizes and $\pi_{0,1}\neq\hat{X}_{0}$}\\ \{((\frac{1}{2},0),1,\pi_{0,1}),((\frac{1}{2},0),3,\pi_{0,2}),((0,1),2,\hat{Y}_{0})\},&\text{if $\kappa_{k}=\hat{X}_{0}$, $\hat{Y}_{0}\notin\{\pi_{0,1},\pi_{0,2}\}$}\\ \{((1,0),1,\hat{X}_{0}),((0,\frac{1}{2}),2,\pi_{0,1}),((0,\frac{1}{2}),3,\pi_{0,2})\},&\text{if $\kappa_{k}=\hat{X}_{0}$, $\hat{Y}_{0}\notin\{\pi_{0,1},\pi_{0,2}\}$}\\ \{(\frac{1}{2},0,1,j),((\frac{1}{2},1),2,\hat{Y}_{0})\},&\text{if $\kappa_{k}=\hat{X}_{0}$, $\hat{Y}_{0}\in\{\pi_{0,1},\pi_{0,2}\},j\in\{\pi_{0,1},\pi_{0,2}\}\setminus\{\hat{Y}_{0}\}$}\\ \{((1,\frac{1}{2}),1,\hat{X}_{0}),((0,\frac{1}{2}),2,j)\},&\text{if $\kappa_{k}=\hat{Y}_{0}$, $\hat{X}_{0}\in\{\pi_{0,1},\pi_{0,2}\},j\in\{\pi_{0,1},\pi_{0,2}\}\setminus\{\hat{X}_{0}\}$}\\ \{((1,\frac{1}{2}),1,\hat{X}_{0}),((0,\frac{1}{2}),2,j)\},&\text{if $\kappa_{k}=\hat{Y}_{0}$, $\hat{X}_{0}\in\{\pi_{0,1},\pi_{0,2}\},j\in\{\pi_{0,1},\pi_{0,2}\}\setminus\{\hat{X}_{0}\}$}\\ \{((1,1),1,\pi_{0,1}))\},&\text{if $\kappa_{0}\in\{\hat{X}_{0},\hat{Y}_{0}\}$ and $\pi_{0,1}=\pi_{0,2}\in\{\hat{X}_{0},\hat{Y}_{0}\}$}\\ \end{cases}$
Suppose now that $\tilde{Z}_{j}^{N}$ is defined for all $j\leq k$. Write $\tilde{Z}_{k}^{N}=z=\{z_{i}\}_{i=1}^{m}$ and $\hat{z}_{r,\ldots,i_{r}}=z\setminus\{z_{i_{n}}\}_{n=1}^{r}$, where each $z_{i}=(p_{i},q_{i},l_{i},j_{i})$. Let $m_{k}$ denote the smallest positive integer not contained in $l_{i}$, and $P(k)=\{j_{i}\}_{i=1}^{m}$. Then the transition from time-step $k$ to $k+1$ is defined in such a way that $\tilde{Z}_{k+1}^{N}$ is equal to
$\tilde{Z}_{k+1}^{N}=\begin{cases}z,&\text{if $\kappa_{k}\notin P(k)$}\\ \hat{z}_{r}\cup\{((p_{r},q_{r}),l_{r},\pi_{k,1})\},&\text{if $\kappa_{0}=j_{r}$, and $\pi_{k,1}=\pi_{k,2}\notin P(k)$}\\ \hat{z}_{r}\cup\{((\frac{p_{r}}{2},\frac{q_{r}}{2}),l_{r},\pi_{k,1}),(\frac{p_{r}}{2},\frac{q_{r}}{2}),m_{k},\pi_{k,2})\},&\text{if }\kappa_{0}=j_{r},\pi_{k,1}\neq\pi_{k,2}\\ &\text{ and }\pi_{k,1},\pi_{k,2}\notin P(k)\\ \hat{z}_{rs}\cup\{((p_{r}+\frac{p_{s}}{2},q_{r}+\frac{q_{s}}{2}),l_{r},j_{r}),((\frac{p_{s}}{2},\frac{q_{s}}{2}),l_{s},j)\},&\text{if $\kappa_{0}=j_{s}$,$\{\pi_{k,1},\pi_{k,2}\}\cap P(k)=\{j_{r}\}$}\\ &\text{ and }j\in\{\pi_{k,1},\pi_{k,2}\}\setminus\{j_{r}\}\\ \hat{z}_{rst}\cup\{((p_{r}+\frac{p_{t}}{2},q_{r}+\frac{q_{t}}{2}),l_{r},j_{r}),((p_{s}+\frac{p_{t}}{2},q_{s}+\frac{q_{t}}{2}),l_{s},j_{s})\},&\text{if $\kappa_{0}=j_{t},$ $\{\pi_{k,1},\pi_{k,2}\}\cap P(k)=\{j_{r},j_{s}\}$}\\ \hat{z}_{rs}\cup\{((p_{r}+p_{s},q_{r}+q_{s}),l_{r},\pi_{k,1})\},&\text{if $\kappa_{k}=j_{s}$ and $\pi_{k,1}=\pi_{k,2}=j_{r}\in P(k)$}\end{cases}$
where we assume $r<s<t$. The transition probabilities are given by
$\tilde{Z}_{k+1}^{N}=\begin{cases}z,&\text{with probability }\frac{N-m}{N}\\ \hat{z}_{r}\cup\{((p_{r},q_{r}),l_{r},\pi_{k,1})\},&\text{w/prob. }\alpha_{N}\frac{1}{N}\frac{N-m-1}{N-1}\\ \hat{z}_{r}\cup\{((\frac{p_{r}}{2},\frac{q_{r}}{2}),l_{r},\pi_{k,1}),(\frac{p_{r}}{2},\frac{q_{r}}{2}),m_{k}+1,\pi_{k,2})\},&\text{w/prob. }(1-\alpha_{N})\frac{1}{N}\frac{N-m-1}{N-1}\frac{N-m-2}{N-2}\\ \hat{z}_{rs}\cup\{((p_{r}+\frac{p_{s}}{2},q_{r}+\frac{q_{s}}{2}),l_{r},j_{r}),((\frac{p_{s}}{2},\frac{q_{s}}{2}),l_{s},j)\},&\text{w/prob. }(1-\alpha_{N})\frac{2}{N}\frac{1}{N-1}\frac{N-m-2}{N-2}\\ \hat{z}_{rst}\cup\{((p_{r}+\frac{p_{t}}{2},q_{r}+\frac{q_{t}}{2}),l_{r},j_{r}),((p_{s}+\frac{p_{t}}{2},q_{s}+\frac{q_{t}}{2}),l_{s},j_{s})\},&\text{w/prob. }(1-\alpha_{N})\frac{1}{N}\frac{2}{(N-1)(N-2)}\\ \hat{z}_{rs}\cup\{((p_{r}+p_{s},q_{r}+q_{s}),l_{r},\pi_{k,1})\},&\text{w/prob.}\alpha_{N}\frac{1}{N}\frac{1}{N-1}\end{cases}$
We can now define the process $Z^{N}$ from $\tilde{Z}^{N}$, by forgetting the labels at the population level. Denote by $\pi:\,[0,1]^{2}\times\mathbb{Z}_{+}\times\mathbb{Z}_{+}\to[0,1]^{2}\times\mathbb{Z}_{+}$ the projection to the first three components (i.e. ignoring the last component). Then we define $Z^{N}=(\pi(\tilde{Z}_{k}^{N}))_{k\in\mathbb{Z}_{+}}$. Note that, as $\tilde{Z}^{N}$ is, conditional on $\mathcal{A}_{N}$, deterministic, so is $Z^{N}$.
Also note that $Z^{N}$ is in distribution the same as the discrete-time Markov process with state space $\mathcal{P}_{m,2}$ defined in (12). We endow $\mathcal{P}_{m,2}$ with a metric $d$ such that $(\mathcal{P}_{m,2},d)$ is a Polish space (a complete and separable metric space); see Section 8.2.
We now describe the transition probabilities of $Z^{N}$, and then prove weak convergence using standard theory of Markov processes.
$Z_{1}^{N}=\begin{cases}\{((1,0),1),((0,1),2)\}&\text{neither $\hat{X}_{0}$ nor$\hat{Y}_{0}$ is a offspring}\\ \{((\frac{1}{2},0),1),((\frac{1}{2},0),3)((0,1),2)\}&\text{$\hat{X}_{0}$ splits without overlap with $\hat{Y}_{0}$}\\ \{((1,0),1)((0,\frac{1}{2}),2),((0,\frac{1}{2}),3)\}&\text{$\hat{Y}_{0}$ splits without overlap with $\hat{X}_{0}$}\\ \{((\frac{1}{2},0),1),((\frac{1}{2},1),2)\}&\text{$\hat{X}_{0}$ splits and oneparent is $\hat{Y}_{0}$}\\ \{((1,\frac{1}{2}),1),((0,\frac{1}{2}),2)\}&\text{$\hat{Y}_{0}$ splits and oneparent is $\hat{X}_{0}$}\\ \{((1,1),1)\},&\text{selfing event between $\hat{X}_{0}$ and $\hat{Y}_{0}$}\end{cases}$
Suppose $Z_{k}^{N}=z=\{(p_{i},l_{i})\}_{i=1}^{m}$ for $p_{i}\in[0,1]^{2}$ and $l_{i}\in\mathbb{Z}_{+}$. Let $m_{k}$ denote the smallest positive integer not contained in $\{l_{i}\}$. Recall $P(k)=\{j_{i}\}_{1\leq i\leq m}$ denotes the corresponding set of population-level labels such that $l_{i}$ is the label of the individual to whom the the label $l_{i}$ (i.e. such that $(p_{i},l_{i},j_{i})\in\tilde{Z}_{k}^{N}$). By $\hat{z}_{rs...i_{j}}$ we denote the set $z\setminus[z_{i_{l}}:1\leq l\leq j]$. The one-step dynamics are described in such a way that $Z_{k+1}^{N}$ is equal to
$\begin{cases}z,&\text{if }\kappa_{k}\notin l,\text{ or }\{\kappa_{k},\pi_{k,1},\pi_{k,2}\}\cap P(k)=\{\kappa_{k}\},\pi_{k,1}=\pi_{k,2}\\ \{(p_{r}+p_{s},l_{r})\}\cup\hat{z}_{rs},&\text{if }\kappa_{k}=j_{s},\pi_{k,1}=\pi_{k,2}=j_{r}\\ \{(p_{r}+\frac{1}{2}p_{t},l_{r}),((p_{s}+\frac{1}{2}p_{t}),l_{s})\}\cup\hat{z}_{rst},&\text{if }l_{t}=\kappa_{k},P(k)\cap\{\pi_{k,2},\pi_{k,1}\}=\{l_{s},l_{t}\}\\ \{(p_{r}+\frac{1}{2}p_{s},l_{r}),((\frac{1}{2}p_{s}),l_{s})\}\cup\hat{z}_{rs},&\text{if }l_{s}=\kappa_{k},P(k)\cap\{\pi_{k,1},\pi_{k,2}\}=\{l_{r}\}\\ \{(\frac{1}{2}p_{r},l_{r}),(\frac{1}{2}p_{r},m_{k})\}\cup\hat{z}_{r},&\text{if}l_{r}=\kappa_{k},P(k)\cap\{\pi_{k,1},\pi_{k,2}\}=\emptyset.\par\end{cases}$
for any $r<s<t$.
The one-step transition probabilities $\{\mathbb{P}(Z_{k+1}^{N}=\cdot|Z_{k}^{N}=q)\}$ can thus be calculated as follows:
$z\rightarrow\begin{cases}z,&\text{w/prob. }\frac{N-m}{N}+\alpha_{N}\frac{m}{N}\frac{N-m-1}{N-1}\\ \{(p_{r}+p_{s},l_{r})\}\cup\hat{z}_{rs},&\text{w/prob. }\alpha_{N}\frac{2}{N}\frac{1}{N-1}\\ \{(p_{r}+\frac{1}{2}p_{t},l_{r}),((p_{s}+\frac{1}{2}p_{t}),l_{s})\}\cup\hat{z}_{rst},&\text{w/prob. }(1-\alpha_{N})\frac{2}{N(N-1)(N-2)}\\ \{(p_{r}+\frac{1}{2}p_{s},l_{r}),((\frac{1}{2}p_{s}),l_{s})\}\cup\hat{z}_{rs},&\text{w/prob. }(1-\alpha_{N})\frac{2}{N(N-1)}\frac{N-m-2}{N-2}\\ \{(\frac{1}{2}p_{r},l_{r}),(\frac{1}{2}p_{r},m_{k})\}\cup\hat{z}_{r},&\text{w/prob. }(1-\alpha_{N})\frac{1}{N}\frac{N-m-1}{N-1}\frac{n-M-2}{n-2}\par\end{cases}$ (33)
for each fixed $r$ and $s\neq r,s\neq t,t\neq r$.
In the critical regime, these transitions are
$z\rightarrow\begin{cases}z,&\text{w/prob. }1-O(N^{-2})\\ \{(p_{r}+p_{s},l_{r})\}\cup\hat{z}_{rs},&\text{w/prob. }2N^{-2}+O(N^{-3})\\ \{(p_{r}+\frac{1}{2}p_{t},l_{r}),((p_{s}+\frac{1}{2}p_{t}),l_{s})\}\cup\hat{z}_{rst},&\text{w/prob. }o(N^{-3})\\ \{(p_{r}+\frac{1}{2}p_{s},l_{r}),((\frac{1}{2}p_{s}),l_{s})\}\cup\hat{z}_{rs},&\text{w/prob. }o(N^{-3})\\ \{(\frac{1}{2}p_{r},l_{r}),(\frac{1}{2}p_{r},m_{k})\}\cup\hat{z}_{r},&\text{w/prob. }\lambda N^{-2}+o(N^{-2})\par\end{cases},$
where elements of $O(f(N))$ are sequences $a_{N}$ such that $a_{N}f(N)^{-1}$ is a bounded sequence; elements of $o(f(N))$ are sequences $a_{N}$ such that $a_{N}f(N)^{-1}$ converges to zero as $N\to\infty$.
Remark 6.1.
The labels of $Z_{k}^{N}$, for any $k\in\mathbb{Z}_{+}$, are contained in $\{1,\ldots,\sup_{i\leq k}|Z_{i}^{N}|\}$.
We let $G_{k}^{N}$ be the projection of the set $Z_{k}^{N}$ to its labels in $\mathbb{Z}_{+}$. From the above calculations, $(G_{k}^{N})_{k\in\mathbb{Z}_{+}}$ is itself a Markov process with one-step transition probabilities
$l\rightarrow\begin{cases}\hat{l}_{r}\cup\{m_{k}\}&\text{w/prob. }\quad\lambda|G_{k}^{N}|N^{-2}+o(N^{-2})\\ l&\text{w/prob. }\quad 1-O(N^{-2})\\ \hat{l}_{r}&\text{w/prob. }\quad|G_{k}^{N}|(|G_{k}^{N}|-1)N^{-2}+O(N^{-2})\end{cases}$ (34)
for some $r$. $G_{k}^{N}$ takes values in finite subsets of $\mathbb{Z}_{+}$ $\Lambda=\bigsqcup_{m=1}^{\infty}\{l_{i}\}_{i=1}^{m}$ for $l_{i}\in\mathbb{Z}_{+}$, which is a Polish space under the metric $d(l,m)=|l\cap m|$. Importantly, this metric engenders the same topology as by the projection map from $\mathcal{P}_{m,2}$.
With the above precise description of the process $G^{N}$, we proceed with a proof of Lemma 4.2.
Proof of Lemma 4.2. We first prove the weak convergence of the processes $\{G^{N}\}_{N\in\mathbb{N}}$ (i.e. ignoring the masses). Precisely, we shall show that if $N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}$, then $(G_{\lfloor tN^{2}\rfloor}^{N})_{t\in\mathbb{R}_{+}}$ converges in distribution under $\mathbb{P}_{{\rm diff}}(\cdot|\mathcal{A}_{N})$ in $\mathcal{D}(\mathbb{R}_{+},\Lambda)$ to a continuous-time Markov process $G_{\lambda}=(G_{\lambda}(t))_{t\geq 0}$ with initial distribution $G_{\lambda}(0)=\{1,2\}$ and transition rates
$l\rightarrow\begin{cases}l\cup\{m(l)\}&\text{ with rate}\quad\lambda|l|\\ \hat{l}_{r}&\text{ with rate}\quad|l|(|l|-1)\\ \end{cases},$ (35)
where $m(l)$ denotes the smallest positive integer not contained in $l$.
Let $T^{(N)}$ be the linear operator on the space $\mathcal{C}_{b}(\Lambda)$ of bounded continuous functions on $\Lambda$ defined by $T^{(N)}f(n)=\mathbb{E}\left[f(G^{N}_{1})\,|\,G^{N}_{0}=l\right]$. The generator $\mathcal{L}^{N}$ of the discrete-time process $G^{N}$ is given by
$\mathcal{L}^{N}f(n):=\mathbb{E}\left[f(G^{N}_{1})-f(G^{N}_{0})\,|\,G^{N}_{0}=l]=(T^{(N)}-I)f(l)\right]$ (36)
which can be explicitly computed using (34). Let $\mathcal{L}$ be the infinitesimal generator of the continuous-time process $G_{\lambda}=(G_{\lambda}(t))_{t\geq 0}$. That is, by (35),
$\mathcal{L}f(l)=\lambda|l|[f(l\cup\{m(l)\})-f(l)]+\sum_{r=1}^{|l|}f(\hat{l}_{r})r(r-1)].$ (37)
It follows from (34) that for all $f\in\mathcal{C}_{b}(\Lambda)$ with finite support,
$\sup_{l\in\Lambda}|N^{2}\mathcal{L}^{N}f(l)-\mathcal{L}f(l)|\to 0\quad\text{as}N\to\infty.$ (38)
Let $\{T(t)\}_{t\in\mathbb{R}_{+}}$ be the semigroup on $\mathcal{C}_{b}(\Lambda)$ of the continuous-time process $G_{\lambda}$. By Theorem 6.5 of Ethier and Kurtz (2009, chapter 1) and (38), it holds that
$\lim_{N}\sup_{0\leq t\leq t_{0}}\sup_{l\in\Lambda}|\left(T^{(N)}\right)^{\lfloor tN^{2}\rfloor}f(l)-T(t)f(l)|=0$ (39)
for all $f$ in the domain of $\mathcal{L}$.
Next, since the fragmentation rate is linear while the coagulation rate is quadratic, we can show as in Griffiths (1991, Theorem 3) that the compact containment condition (Ethier and Kurtz, 2009, (7.9), page 129) holds. That is, for any $\varepsilon\in(0,1)$ and $T\in(0,\infty)$, there is a constant $K_{\varepsilon,T}\in(0,\infty)$ such that
$\limsup_{N\to\infty}\mathbb{P}\left(\sup_{t\in[0,T]}|G^{N}_{[tN^{2}]}|\geq K_{\varepsilon,T}\right)\leq\varepsilon.$ (40)
We now demonstrate how to prove (40) by a martingale argument by examining the continuous-time Markovian process $L:=|G_{\lambda}|$ taking values in $\mathbb{Z}_{+}$ capturing the size of the cross-section of $G_{\lambda}$ at any time-point. Let $\mathcal{Q}$ denote the generator of $L$. We can see that, for any $f$ in $\mathcal{C}_{b}(\mathbb{Z}_{+})$ that
$\mathcal{Q}f(n)=\lambda n(f(n+1)-f(n))+n(n-1)(f(n-1)-f(n))$ (41)
First, we invoke a general relation between Markov processes and martingales. For any $f\in\mathcal{C}_{b}(\mathbb{Z}_{+})$,
[…]
is a martingale with quadratic variation $\langle M\rangle_{t}=\int_{0}^{t}\mathcal{Q}(f^{2})(L_{s})-2f(L_{s})\mathcal{Q}f(L_{s})\,ds$; see Kipnis and Landim (1998, Lemma 5.1) or Ethier and Kurtz (2009, Proposition 4.1.7). A truncation argument will enable us to take $f$ to be the identity function (i.e. when $f(n)=n$ for all $n\in\mathbb{Z}_{+}$) to obtain from (37) that
$L_{t}=L_{0}+\int_{0}^{t}-L_{s}^{2}+(\lambda+1)L_{s}\,ds+M^{\rm Id}(t)\quad\text{for }t\in\mathbb{R}_{+},$ (43)
where $M^{\rm Id}$ is a martingale with quadratic variation $\langle M^{\rm Id}\rangle_{t}=\int_{0}^{t}L_{s}^{2}+(\lambda-1)L_{s}\,ds$. In the above, we used the fact that, from (37), when $f$ is the identity function,
$\mathcal{Q}f(n)=$ $\lambda n-n(n-1)=-n^{2}+(\lambda+1)n\leq\frac{(\lambda+1)^{2}}{4}$ (44)
$\mathcal{Q}f^{2}(n)=$ $\lambda n[(n+1)^{2}-n^{2}]+n(n-1)[(n-1)^{2}-n^{2}]$
$\,=-2n^{3}+(3+2\lambda)n^{2}+(\lambda-1)n$ (45)
$\mathcal{Q}f^{2}(n)-2f(n)\mathcal{Q}f(n)=$ $n^{2}+(\lambda-1)n$
Therefore, from (43) and (44),
$\sup_{t\in[0,T]}L_{t}\leq L_{0}+\frac{(\lambda+1)^{2}}{4}\,T+\sup_{t\in[0,T]}M^{\mathrm{Id}}(t).$
By Doob’s maximal inequality, the above formula for $\langle M^{\mathrm{Id}}\rangle_{t}$, and the fact that $\sup_{t\in[0,T]}\mathbb{E}[L^{2}_{t}]<\infty$, which follows from (45), there exists a constant $C_{T}\in(0,\infty)$ such that
$\mathbb{P}\left(\sup_{t\in[0,T]}M^{\mathrm{Id}}(t)\geq K\right)\leq\frac{C_{T}}{K^{2}}\quad\text{for all }K\in(0,\infty).$
From these last two inequalities, for any $\varepsilon\in(0,1)$ there exists a constant $K_{\varepsilon,T}$ such that
$\mathbb{P}\left(\sup_{t\in[0,T]}L_{t}\geq K_{\varepsilon,T}\right)\leq\varepsilon.$ (46)
[…]
The proof for the process $\{Z^{N}\}$ follows from the same argument, using (33) instead of (34). ∎
Next, we couple random walks on $(Z_{k}^{N})_{k\in\mathbb{Z}_{+}}$ with those on $\mathcal{G}_{N}$ to ensure that their limits conditional on the pedigree are identical. Using an argument like that of Birkner et al. (2013) and Diamantidis et al. (2024), we show that, as $\lambda\to\infty$, the limits conditional on the pedigree converge to the unconditional limit. This proof technique is further elaborated in Section 6.3 to establish the limit conditional on the pedigree for the partial-selfing regime.
6.1.1 Potential ancestors of two sample lineages
To model the ancestry of our two sample lineages, we use a branching-coalescing-overlapping graph which differs from the ancestral recombination graph (Griffiths, 1991; Griffiths and Marjoram, 1997; Hudson, 1983b) and the ancestral selection graph (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997) only in how ancestral genetic material is traced back in time through the graph.
Recall from (34) that $G_{k}^{N}$ is the projection of $Z_{k}^{N}$ to its labels in $\mathbb{Z}_{+}$. The projected process $G^{N}:=(G_{k}^{N})_{k\in\mathbb{Z}_{+}}$ can be viewed graphically as having two initial nodes $(1,0)$ and $(0,1)$ under the sampling scheme $\mathbb{P}_{{\rm diff}}$. These two nodes correspond to the masses of the two lineages in the two individuals from whom we sampled the two lineages at time-step 00, namely $\hat{X}_{0}$ and $\hat{Y}_{0}$.
We now consider two conditionally independent simple random walks on $G^{N}$. Let $x_{0}^{N}=(1,0)$ and $y_{0}^{N}=(0,1)$. Suppose $x_{j}^{N}$ is defined for all $j\leq k$, for induction. If $\hat{X}_{k}$ undergoes an outcrossing at time-step $k$, then we define $x_{k}^{N}$ to follow the path of the outcrossing at $Z_{k}^{N}$ that corresponds to the edge along which $\hat{X}_{k}^{N}$ travels. Similarly, $y_{k}^{N}$ can be defined in the same way. This gives two discrete-time random walks on $G^{N}$ starting at $((1,0),\,(0,1))$ and satisfying
  • (i)
    $\{x_{k}^{N},y_{k}^{N}\}\subset G_{k}^{N}$ for all $k\in\mathbb{Z}_{+}$,
  • (ii)
    at any fragmentation event, each of $(x_{k}^{N})_{k\in\mathbb{Z}_{+}}$ and $(y_{k}^{N})_{k\in\mathbb{Z}_{+}}$ will follow each of the two paths available with equal (i.e. 1/2) probability.
Lemma 4.2 can readily be extended to the joint convergence in Lemma 6.2 which we omit the proof.
Lemma 6.2.
Suppose $\lim_{N\to\infty}N(1-\alpha_{N})=\lambda\in\mathbb{R}_{+}$. Then $(Z_{\lfloor tN^{2}\rfloor}^{N},x_{\lfloor tN^{2}\rfloor}^{N},y_{\lfloor tN^{2}\rfloor}^{N})_{t\in\mathbb{R}_{+}}$ converges in distribution in $\mathcal{D}(\mathbb{R}_{+},\mathcal{P}_{m,2}\times[0,1]^{2}\times[0,1]^{2})$ under $\mathbb{P}_{{\rm diff}}$ to a process $(Z_{\lambda},x_{\lambda},y_{\lambda})$, where $Z_{\lambda}=(Z_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ is a fragmentation-coagulation process of two immiscible unit masses with rate $\lambda$. The joint process $(x_{\lambda},y_{\lambda})=(x_{\lambda}(t),y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ is a continuous-time $G_{\lambda}\times G_{\lambda}$-valued process describe before (13).
6.1.2 Conditional limit for distinct individuals
Here, in the continuous-time limit, coalescence between overlapping particles is instantaneous, while this is in general not true in the discrete-time model. We first show that this behavior is true in the limit for our random walks on the pedigree by the following lemma.
Lemma 6.3.
Suppose $\lim_{N\to\infty}N(1-\alpha_{N})=\lambda\in\mathbb{R}_{+}$. Let $\tau_{O}^{N}$ be the time of the first overlap of our two sample lineages, that is,
$\tau_{O}^{N}:=\inf\{k\in\mathbb{Z}_{+}:\hat{X}_{k}=\hat{Y}_{k}\}.$
Then $\tau^{(N,2)}-\tau_{O}^{N}$ converges to 00 in distribution as $N\to\infty$ under $\mathbb{P}_{{\rm diff}}(\cdot|\mathcal{A}_{N})$.
Proof. Fix $\varepsilon>0$. Observe that
$\mathbb{P}_{{\rm diff}}\big{(}N^{-2}(\tau^{(N,2)}-\tau_{O}^{N})>\varepsilon|\mathcal{A}_{N}\big{)}\leq\frac{1}{\varepsilon}\mathbb{P}_{{\rm diff}}\big{(}N^{-2}(\tau^{(N,2)}-\tau_{O}^{N})>\varepsilon\big{)}$
by the conditional Markov inequality. The upper bound established here goes to zero as the probability that there is only a single overlap before coalescence is $\frac{1}{2-\alpha_{N}}$, which converges to one, by Lemma 8.1 and the time it takes between coalescence and that last overlap has mean going to zero by Lemma 8.2. ∎
We are now prepared to prove Theorem 4.3 in the limited-outcrossing regime.
Theorem 6.4.
Let $x_{\lambda}$ and $y_{\lambda}$ be independent random walks on a given ancestral graph $G_{\lambda}$, and $T_{\lambda}:=\inf\{t\in\mathbb{R}_{+}:\,x_{\lambda}(t)=y_{\lambda}(t)\}$ be their first meeting time defined in (13). Suppose $N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{>0}$ as $N\to\infty$. Then, for any fixed $t>0$,
$\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})\,{\buildrel{d}\over{\longrightarrow}}\,\mathbb{P}_{x_{0}\neq y_{0}}(T_{\lambda}>t|G_{\lambda}).$
Proof. Let $x_{k}^{N}$ and $y_{k}^{N}$ denote the projection of $\hat{X}_{k}$ and $\hat{Y}_{k}$ to $G_{k}^{N}$, described before Lemma 6.2. These are paths on $G^{N}$ which follow each path of each outcrossing with equal likelihood. Let $T_{\lambda}^{N}$ denote the pairwise coalescence time of $x_{k}^{N}$ and $y_{k}^{N}$ on $G_{k}^{N}$ i.e.
$T_{\lambda}^{N}:=\inf\{k\in\mathbb{Z}_{+}:x_{k}^{N}=y_{k}^{N}\}.$
Further, let
$\tau_{O}^{N}:=\inf\{k\in\mathbb{Z}_{+}:\hat{X}_{k}=\hat{Y}_{k}\}$
be the first time to overlap for the random walks on the pedigree.
By construction $\tau_{O}^{N}=T_{\lambda}^{N}$ pointwise and knowing the pedigree tells exactly as much information about $\tau_{O}^{N}$ as knowing $G_{\lambda}$. In particular,
$\mathbb{P}_{{\rm diff}}(N^{-2}\tau_{O}^{N}>t|\mathcal{A}_{N})=\mathbb{P}_{x_{0}^{N}\neq y_{0}^{N}}(N^{-2}T_{\lambda}^{N}>t|G^{N}).$ (47)
It follows from Lemma 6.2 that the right hand side of (47) converges as $N\to\infty$ to $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$.
Indeed, for any graph $\tilde{G}$ in $\mathcal{D}(\mathbb{R}_{+},\Lambda)$ with initial particles $\tilde{x},\tilde{y}$, we can define $\tilde{T}$, as in (13),
$\tilde{T}:=\inf\{t\geq 0:\tilde{x}(t)=\tilde{y}(t)\}.$
The functional $\Psi$, defined by
$\Psi(\tilde{G}):=\mathbb{P}(\tilde{T}>t\mid\tilde{G}),$
is a continuous functional of the graph $\tilde{G}$. $\Psi$ depends only on the discrete structure of the graph $\tilde{G}$ and its edge lengths up to time $t$. By lemma 6.2 the convergence of $G^{N}$ to $G_{\lambda}$ in $\mathcal{D}(\mathbb{R}_{+},\Lambda)$ and Proposition 5.3 of Ethier and Kurtz (2009, Chapter 3), the discrete structure of $G^{N}$ is eventually fixed for large enough $N$ on any compact interval $I$ of which $[0,t]$ is a proper subspace. Further, on $I$ the edge lengths converge uniformly in $N$.
From the definition of $\Psi$, we have $\Psi(G^{N})=\mathbb{P}(\tilde{T}_{\lambda}^{N}>t\mid G^{N})$, where $\tilde{T}_{\lambda}^{N}$ is the continuous-time analogue of $T_{\lambda}^{N}$; i.e.
$\tilde{T}_{\lambda}^{N}:=\inf\{t\geq 0:x_{\lfloor tN^{2}\rfloor}^{N}=y_{\lfloor tN^{2}\rfloor}^{N}\}.$
Note that $\tilde{T}_{\lambda}^{N}\geq N^{-2}T_{\lambda}^{N}$ for all $N$ and their difference $\tilde{T}_{\lambda}^{N}-N^{-2}T_{\lambda}^{N}$ is pointwise bounded by $\frac{1}{N^{2}}$. Thus, their limits conditional on the pedigree are the same. Therefore $\mathbb{P}(N^{-2}\tau_{O}^{N}>t|\mathcal{A}_{N})$ converges as $N\to\infty$ to $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$.
The result follows from the fact that the limits conditional on the pedigree of $\tau_{O}^{N}$ and $\tau^{(N,2)}$ under $\mathbb{P}_{\text{diff}}(\cdot\mid\mathcal{A}_{N})$ must be the same by Lemma 6.3. ∎
6.1.3 Conditional limit under $\mathbb{P}_{\rm{same}}$
Selfing is nearly ubiquitous in the limited-outcrossing regime. In particular, any sample of two lineages in the same individual will undergo very many selfing reproduction events before they have the chance to undergo a split. This is reflected in Lemma 6.5.
Lemma 6.5.
If $\alpha_{N}\rightarrow 1$, then $N^{-2}\tau^{(N,2)}$ converges to 00 in distribution under $\mathbb{P}_{\rm{same}}(\cdot|\mathcal{A}_{N})$.
Proof. Fix $\varepsilon>0$. By the conditional Markov inequality, for $t>0$,
$\mathbb{P}(\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})>\varepsilon)\leq\frac{1}{\varepsilon}\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t).$ (48)
The limit conditional on the pedigree therefore follows from the unconditional limit established in Theorem 8.3 as $\mathbb{P}_{\rm{same}}(\mathcal{O}=0)=\frac{\alpha_{n}}{2-\alpha_{N}}$, which converges to $1$, and under the law of $\mathbb{P}_{\rm{same}}(\cdot|\mathcal{O}=1)$ $\tau^{(N,2)}$ converges in distribution to 00 by Theorem 3.1. ∎
6.2 Proofs for the negligible-outcrossing regime
The proofs for the case of negligible outcrossing are straightforward. They follow directly from the results in the limited-outcrossing regime.
Corollary 6.6.
If $N(1-\alpha_{N})\rightarrow 0$, then for any $t\in\mathbb{R}_{+}$, $\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2})>t|\mathcal{A}_{N})$ converges in distribution as $N\to\infty$ to a Bernoulli random variable with success probability $e^{-2t}$.
Proof. By Theorem 6.4, $\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})$ converges as $N\to\infty$ to $\mathbb{P}_{x_{0}\neq y_{0}}(T_{0}>t|Z_{0})$. $Z_{0}$ branches at rate 00 and coagulates the two existing nodes at rate $2$. In particular, conditional on $Z_{0}$, $T_{0}$ is known exactly, and it distributed unconditionally as an exponential random variable with rate $2$. Therefore
$\mathbb{P}(\mathbb{P}_{x_{0}\neq y_{0}}(T_{0}>t|Z_{0})=1)=e^{-2t}=1-\mathbb{P}(\mathbb{P}_{x_{0}\neq y_{0}}(T_{0}>t|Z_{0})=0)$
This is exactly the claim. ∎
6.3 Proofs for the partial-selfing regime
In this section, we assume that $N(1-\alpha_{N})\to\infty$ and $\alpha_{N}\to\alpha\in[0,1]$ as $N\to\infty$.
We shall see that, for samples in different individuals, the distribution of $\tau^{(N,2)}$ conditioned on the pedigree is related to the minimum of two pairwise coalescence times conditionally independent with respect to the pedigree. These two conditionally independent pairwise coalescence times are the pairwise coalescence times of two conditionally independent pairs of random walks, $(X,Y)$ and $(X^{\prime},Y^{\prime})$, on $J\times\mathbb{Z}_{+}$. This is a slightly more complicated argument than was presented in the proof of Corollary 22 in Section 5.1.
6.3.1 Two conditionally independent pairs of genes
In the partial-selfing, direct analysis regime of $\tau^{(N,2)}$ conditional on $\mathcal{A}_{N}$ is more difficult. We can proceed by connecting the behavior of $\tau^{(N,2)}$ to that of two copies of $\tau^{(N,2)}$ conditionally independent with respect to the pedigree with the following lemma.
Lemma 6.7.
For any fixed $t\in\mathbb{R}_{+}$ and $N\geq 2$,
$\mathbb{E}_{\rm{diff}}\left[\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{N,2}>t|\mathcal{A}_{N})^{2}\right]=\mathbb{P}_{{\rm diff}}\left(\tau\wedge\tau^{\prime}>N^{2}t\right).$
The proof follows that of Lemma 5.1 and is omitted.
Let $\tau$ and $\tau^{\prime}$ be conditionally independent pairwise coalescence times with respect to $\mathcal{A}_{N}$. $\tau$ and $\tau^{\prime}$ are the pairwise coalescence times of two conditionally independent pairs of random walks on the same pedigree, denoted by $(X,Y)$ and $(X^{\prime},Y^{\prime})$. The individual occupied by $X_{k},Y_{k},X^{\prime}_{k},Y^{\prime}_{k}$ are denoted respectively by $\hat{X}_{k},\hat{Y}_{k},\hat{X}^{\prime}_{k},\hat{Y}^{\prime}_{k}$ respectively. We denote by
$\hat{X}_{0}=\hat{X}^{\prime}_{0}\quad\text{and}\quad\hat{Y}_{0}=\hat{Y}^{\prime}_{0}$ (49)
the sampled individuals. The set of individuals occupied by (at least) one of these random walks $k$ time-steps in the past is denoted by
$\widehat{I}_{k}:=\{\hat{X}_{k},\hat{Y}_{k},\hat{X}^{\prime}_{k},\hat{Y}^{\prime}_{k}\}.$ (50)
In other words, $\widehat{I}_{k}$ is the set of ancestral individuals for two conditionally independent genealogies $k$ time-steps in the past.
Just as in the unconditional case, the key objects of study are overlaps and splits. In this case, though, we need to track 4 sample lineages, not simply two. Luckily, at least in the partial-selfing regime, our analysis will work in much the same way.
We now define the overlap and splitting times. We let $\tau_{0}^{O}=0$ by convention, and, for $i\geq 1$ we let
$\tau_{i}^{O}=\inf\{k\in\mathbb{Z}_{+}:k>\tau_{i-1}^{O},|\widehat{I}_{k}|<|\widehat{I}_{k-1}|\}.$
These can be understood as the overlap times, where (at least) two of our four sample lineages transition to belonging in fewer individuals than before, with each other.
With respect to splits we define
$\tau_{0}^{D}:=\inf\{k\in\mathbb{Z}_{+}:|\widehat{I}_{k}|=4\},$
the first time that all four sample lineages are in four distinct individuals, if it exists. We then iteratively define, for $i\geq 1$,
$\tau_{i}^{D}:=\inf\{k\in\mathbb{Z}_{+}:|\widehat{I}_{k}|>|\widehat{I}_{k-1}|\}.$
We let $\mathcal{O}$ denote the total number of finite overlap times before $\tau\wedge\tau^{\prime}$ i.e.,
$\mathcal{O}:=|\{i\in\mathbb{Z}_{>}:\tau_{i}^{O}\leq\tau\wedge\tau^{\prime}\}|.$
Lemma 6.8.
Let $\Gamma_{N}$ denote the event
$\left\{0=\tau_{0}^{O}<\tau_{0}^{D}<\tau_{1}^{O}<\tau_{1}^{D}<\ldots<\tau_{\mathcal{O}-1}^{D}<\tau_{\mathcal{O}}^{O}\leq\tau\wedge\tau^{\prime}\right\}.$ (51)
Then $\mathbb{P}_{{\rm diff}}(\Gamma_{N})$ converges to $1$ as $N$ tends to infinity if $N(1-\alpha_{N})\rightarrow\infty$. Further, $\mathcal{O}$ under $\mathbb{P}_{{\rm diff}}(\cdot|\Gamma_{N})$ is geometric with parameter $\frac{1}{6}\frac{\alpha_{N}}{2-\alpha_{N}}$.
The primary use of this lemma is to get a telescoping sum representation of $\tau\wedge\tau^{\prime}$:
$\tau\wedge\tau^{\prime}=\tau\wedge\tau^{\prime}-\tau_{\mathcal{O}}^{O}+\sum_{i=0}^{\mathcal{O}}\left(\tau_{i}^{O}-\tau_{i-1}^{D}\right)+\sum_{i=1}^{\mathcal{O}-1}\left(\tau_{i}^{D}-\tau_{i}^{O}\right).$ (52)
In order to prove Lemma 6.8, we first prove two other lemmas. We begin by calculating the probability that $\tau_{0}^{D}<\tau_{1}^{O}$. We then calculate the probability that any overlap which occurs before $\tau\wedge\tau^{\prime}$ is followed either by a split or coalescence. We also quantify the times of the splits.
Lemma 6.9.
As $N\to\infty$,
$\mathbb{P}_{{\rm diff}}(\tau_{0}^{D}<\tau_{1}^{O})\to 1.$
Further, $N^{-2}\tau_{0}^{D}$ converges to 00 in distribution under the law $\mathbb{P}_{{\rm diff}}(\cdot\,\mid\tau_{0}^{D}<\tau_{1}^{O})$.
Proof. Recall that the two sampled individuals for the two pairs of conditionally independent gene copies are given by (49). The probability that either individual splits in a generation is $(1-\alpha_{N})\frac{2}{N}$. Let $\lambda_{N}$ be the rate at which either individual splits without an overlap event for either pair of the gene copies. Then
$\lambda_{N}=(1-\alpha_{N})\frac{2}{N}\left[\frac{N-2}{N-1}\frac{N-3}{N-2}\frac{1}{2}+\frac{2}{N-1}\frac{1}{4}\right]=(1-\alpha_{N})\frac{N-2}{N(N-1)}.$
The term $\frac{N-2}{N-1}\frac{N-3}{N-2}$ is the probability that immediately after splitting, the parents of the splitting individual are both different from the other individual. The term $\frac{2}{N-1}$ is the probability that immediately after splitting, one of the parents of the splitting individual is the other individual.
Let $\mu_{N}$ be the rate at which there is an overlap event for either pair of the gene copies. Then
$\mu_{N}=(1-\alpha_{N})\frac{2}{N}\frac{2}{N-1}\frac{3}{4}+\alpha_{N}\frac{2}{N(N-1)}=\frac{3-\alpha_{N}}{N(N-1)}.$
Therefore, the probability, that either individual splits without an overlap event is $\frac{\lambda_{N}}{\mu_{N}+\lambda_{N}}=\frac{(1-\alpha_{N})(N-2)}{3-\alpha_{N}+(1-\alpha_{N})(N-2)}$. The time it takes for such a desired split to occur, $b_{0}$, can therefore be observed to be geometric with parameter
$\frac{\lambda_{N}}{1-\frac{\mu_{N}}{\lambda_{N}}}=\frac{\lambda_{N}^{2}}{\lambda_{N}-\mu_{N}}.$
Once either individual splits without an overlap event, we have $4$ sample lineages in $3$ individuals. The rate $\lambda_{N}^{\prime}$ at which the occupied individual containing two sample lineages splits without an overlap event is
$(1-\alpha_{N})\frac{1}{N}\left(\frac{2}{N-1}\frac{N-3}{N-2}+\frac{N-3}{N-1}\frac{N-4}{N-2}\right)=\frac{(1-\alpha_{N})(N-3)(2N-7)}{2N(N-1)(N-2)}.$
The rate $\mu_{N}^{\prime}$ depends more precisely on which individual is chosen as the offspring. If the offspring contains two conditionally independent sample lineages and one parent is occupied, then there is a $\frac{3}{4}$ chance of having an overlap event. Regardless of which occupied offspring is chosen, if both parents are occupied then an overlap event is guaranteed. If an overlap is due to selfing then an overlap occurs if and only if both parent and offspring are occupied. Thus, $\mu_{N}^{\prime}$ can be calculated as
$(1-\alpha_{N})\left(\frac{1}{N}\frac{4}{N-1}\frac{N-3}{N-2}\frac{3}{4}+\frac{2}{N}\frac{4}{N-1}\frac{N-3}{N-2}\frac{1}{2}+\frac{3}{N}\frac{2}{N-1}\frac{1}{N-2}\right)+\alpha_{N}\frac{6}{N(N-1)}.$
By some simplification this becomes
$\frac{N(7-\alpha_{N})-15+3\alpha_{N}}{N(N-1)(N-2)}.$
Therefore the probability that we have four lineages in four individuals before any overlap event is
$1-\frac{\mu_{N}^{\prime}}{\lambda_{N}^{\prime}+\mu_{N}^{\prime}}=1-\frac{2(N(7-\alpha_{N})-15+3\alpha_{N})}{(1-\alpha_{N})(2N^{2}-15N+18)+12N-12},$
and the time, $b_{1}$ it takes for this to occur is geometric with parameter $\frac{(\lambda_{N}^{\prime})^{2}}{\lambda_{N}^{\prime}-\mu_{N}^{\prime}}$. This gives the claim using the fact that $N(1-\alpha_{N})\rightarrow\infty$. ∎
Recall $\widehat{I}_{k}$ defined in (50). We consider the case where there are two sample lineages from the same pair in the same individual, and where the other two sample lineages are in distinct individuals. Given this state, what is the probability $r_{N}$ that we return to four sample lineages in four individuals before coalescence, or that there is a coalescence before another overlap event? This is answered in the limit as $N\to\infty$ by the following lemma.
Lemma 6.10.
As $N\to\infty$, we have $r_{N}\to 1$ where
$r_{N}:=\mathbb{P}_{{\rm diff}}\left(\tau\wedge\tau^{\prime}<\tau_{i+1}^{O}\text{ or }\tau_{i+1}^{D}<\tau\wedge\tau^{\prime}\Big{|}\;|\widehat{I}_{\tau_{i}^{O}}|=3,\hat{X}_{\tau_{i}^{O}}=\hat{Y}_{\tau_{i}^{O}}\text{ or }\hat{X^{\prime}}_{\tau_{i}^{O}}=\hat{Y^{\prime}}_{\tau_{i}^{O}}\right).$
Proof. The probability $\lambda$ that an occupied individual containing two sample lineages splits or else has these two lineages coalesce via splitting in a single time-step, both without an overlap event, is
$(1-\alpha_{N})\frac{1}{N}\frac{N-3}{N-1}\frac{N-4}{N-2}+(1-\alpha_{N})\frac{1}{N}\frac{4}{N-1}\frac{N-3}{N-2}\frac{1}{2}+\alpha_{N}\frac{1}{N}\frac{N-3}{N-1}\frac{1}{2}$ (53)
$=\,\frac{(N-3)((2-\alpha_{N})N+2(1-\alpha_{N})-2)}{2N(N-1)(N-2)}$ (54)
The probability the coalescent event occurs before the splitting event can be calculated as the ratio of selfing coalescence probability and $\lambda$. The probability $\mu$ that there is an overlap event involving the three occupied individuals is
$\frac{N(7-\alpha_{N})-3(5-\alpha_{N})}{N(N-1)(N-2)}.$
Therefore the probability of the two sample lineages in the same individual coalescing or splitting before another overlap event is
$\frac{\lambda}{\lambda+\mu}=1-\frac{\mu}{\lambda+\mu}=1-\frac{(14-2\alpha_{N})N-15+3\alpha_{N}}{(2-\alpha_{N})N^{2}+(8-\alpha_{N})N-30+12\alpha_{N}}.$
This gives the claim. ∎
Proof of Lemma 6.8. The behavior of $\mathcal{O}$ is determined by the structure of the representation (51). Indeed, as $\tau_{0}^{D}$ is finite on $\Gamma_{N}$ we reach a state where all four lineages are in four distinct individuals. From this state, it is only possible for $|\widehat{I}_{k}|$ to transition from four to three. The presence of splits between each overlap indicates that $|\widehat{I}_{k}|$ oscillates between 3 and 4 after $\tau_{0}^{D}$ and before $\tau\wedge\tau^{\prime}$.
Therefore, the relevant overlaps of pairs occur are those that occur from the state in which all four lineages are in distinct individuals. There are $\binom{4}{2}$ equally likely possible pairs of lineages to overlap, but only two, those between $x$ and $y$ or $x^{\prime}$ and $y^{\prime}$, that could result in coalescence. Therefore there is a $\frac{1}{3}$ chance of either of these two potentially coalescent pairs overlapping at an overlap time. These lineages coalesce instantaneously with a further probability of $\frac{1}{2}$, but they may coalesce before the next split with some positive probability.
Let $U$ denote the number of selfing events, in the single individual containing two sample lineages, before the next split, conditional on $\Gamma_{N}$. The probability, conditional on $\Gamma_{N}$, that we have a selfing event before a split is $\alpha_{N}$. Therefore, $\mathbb{P}(U=j)=\alpha_{N}^{j}(1-\alpha_{N})$ as in (2). When $U=k$, there is a $2^{-1}+\ldots+2^{-k}=1-2^{-k}$ chance of coalescing. Therefore the probability of coalescing due to any given overlap, conditional on $\Gamma_{N}$, is
$\frac{1}{3}\left(\frac{1}{2}+\frac{1}{2}\mathbb{E}\left[1-2^{-U})\right]\right)=\frac{1}{6}\frac{\alpha_{N}}{2-\alpha_{N}}.$
It suffices now to show that the probability of $\Gamma_{N}$ converges to one as $N\to\infty$. This is precisely that probability that $\tau_{0}^{D}<\tau_{1}^{O}$ multiplied by the probability that each of the next $\mathcal{O}-1$ splits each occur before the next overlap, or else coalesce. The probability that $\tau_{0}^{D}<\tau_{1}^{O}$ converges to one by Lemma 6.9.
The probability that we have the structure in $\Gamma_{N}$ after $\tau_{0}^{D}$ is exactly $\mathbb{E}\left[r_{N}^{\mathcal{O}}\right]$, where $r_{N}$ is the probability given by Lemma 6.10.
As $r_{N}$ converges to one as $N\to\infty$ and $\mathcal{O}$ is simply a geometric random variable with parameter strictly bounded away from 00, $\mathbb{E}\left[r_{N}^{\mathcal{O}}\right]$ converges to one as $N\to\infty$. This gives the claim. ∎
It remains simply to quantify the gap between overlaps and splits given by the following lemma, and the gap between the final overlap and coalescence.
Lemma 6.11.
As $N\to\infty$, the conditional expectations $\left\{\,\mathbb{E}\left[N^{-2}(\tau_{i}^{D}-\tau_{i-1}^{O})|\tau_{i}^{D}<\tau_{i}^{O}\right]\,\right\}_{i\in\mathbb{Z}_{>0}}$ and $\mathbb{E}\left[N^{-2}(\tau^{(N,2)}-\tau_{\mathcal{O}}^{O}))\;\Big{|}\;|\widehat{I}_{\tau_{\mathcal{O}}^{O}}|=3\right]$ all converge to 00 in distribution.
Proof. We demonstrate the first limit. The probability of splitting of the multiply-occupied individual in a single step without any overlap event is
$\frac{(1-\alpha_{N})(N-3)(N-4)}{N(N-1)(N-2)}.$
Therefore, conditioning on knowing $\tau_{i}^{D}<\tau_{i}^{O}$ implies $(\tau_{i}^{D}-\tau_{i-1}^{O})$ is geometric with this parameter. This gives the claim. We now prove the second limit. The difference $\tau^{(N,2)}-\tau_{\mathcal{O}}^{O}$ is stochastically dominated by the time it takes for the relevant occupied individual to split, which is geometric with parameter $(1-\alpha_{N})N^{-1}$. Therefore,
$\mathbb{E}\left[N^{-2}(\tau^{(N,2)}-\tau_{\mathcal{O}}^{O})\right]\leq\frac{N}{1-\alpha_{N}}N^{-2}=\frac{1}{N(1-\alpha_{N})}$
which converges to 00 as $N(1-\alpha_{N})\rightarrow\infty$. ∎
We are now prepared to prove the limiting behavior of $\tau\wedge\tau^{\prime}$ in the partial-selfing regime.
Theorem 6.12.
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ and $N(1-\alpha_{N})\rightarrow\infty$. Then as $N\rightarrow\infty$, $N^{-2}(\tau\wedge\tau^{\prime})$ converges in distribution to an exponential random variable with rate $\frac{4}{2-\alpha}$.
Proof. By Lemma 6.8, for any $t\in\mathbb{R}$,
$\mathbb{E}_{\rm{diff}}\left[e^{itN^{-2}(\tau\wedge\tau^{\prime})}\right]=\mathbb{E}_{\rm{diff}}\left[e^{itN^{-2}(\tau\wedge\tau^{\prime})}|\Gamma_{N}\right]+o(1).$ (55)
We therefore can decompose $\tau\wedge\tau^{\prime}$ as in (52).
Let $\sigma_{N}$ denote the characteristic function of $\tau_{i}^{D}-\tau_{i}^{O}$ conditional on $\Gamma_{N}$ for $i\geq 1$, $\sigma_{N}^{\prime}$ denote the characteristic function of $\tau_{0}^{D}$, and $\varphi_{N}$ denote the characteristic function of each of the $\tau_{i}^{O}-\tau_{i-1}^{D}$, conditional on $\Gamma_{N}$. The $\{\tau_{i}^{O}-\tau_{i-1}^{D}\}_{i=1}^{\mathcal{O}}$ all have the same characteristic function as they are all independent and identically distributed. Finally, let $\psi_{N}$ denote the characteristic function of $\tau\wedge\tau^{\prime}-\tau_{\mathcal{O}}^{O}$, again conditional on $\Gamma_{N}$.
By Lemmas 6.11 and 6.9 we have that $\psi_{N}(tN^{-2}),\,\sigma_{N}(tN^{-2}),\,\sigma_{N}^{\prime}(tN^{-2})$ all converge to 1 as $N\to\infty$. Note that $\varphi_{N}$ and $\psi_{N}$ are the characteristic functions of geometric random variables with rates $2\binom{4}{2}N^{-2}+O(N^{-3})$ and $\frac{1}{2}\alpha_{N}N^{-1}$, respectively. Therefore,
$\varphi_{N}(t)=\frac{2\binom{4}{2}N^{-2}+O(N^{-3})}{1-e^{it}(1-2\binom{4}{2}N^{-2}-O(N^{-3}))}\quad\text{and}\quad\psi_{N}(t)=\frac{\frac{1}{2}\alpha_{N}N^{-1}}{1-e^{it}(1-\frac{1}{2}\alpha_{N}N^{-1})},$
respectively. By Lemma 6.9,
$\mathbb{E}_{\rm{diff}}\left[e^{itN^{-2}(\tau\wedge\tau^{\prime})}|\Gamma_{N}\right]\,=\,\sigma_{N}^{\prime}(tN^{-2})\psi_{N}(tN^{-2})\sum_{j=1}^{\infty}\mathbb{P}_{{\rm diff}}(\mathcal{O}=j|\Gamma_{N})\varphi_{N}(tN^{-2})^{j}\sigma_{N}^{j-1}.$ (56)
As $\mathbb{P}_{{\rm diff}}(\mathcal{O}=j|\Gamma_{N})=\left(\beta_{N}\right)^{j}(1-\beta_{N})$ for $\beta_{N}=\frac{1}{6}\frac{\alpha_{N}}{2-\alpha_{N}}$ by Lemma 6.9, (56) becomes
$\sigma_{N}^{\prime}(tN^{-2})\psi_{N}(tN^{-2})\frac{\beta_{N}\varphi_{N}(tN^{-2})}{1-(1-\beta_{N})\varphi_{N}(tN^{-2})\psi_{N}(tN^{-2})}.$ (57)
Note, $\varphi_{N}$ is the characteristic function of a geometric random variable with parameter $2\binom{4}{2}N^{-2}+O(N^{-3})$. Therefore $\varphi_{N}(tN^{-2})$ converges to the characteristic function $\varphi$ of an exponential random variable with rate $122\binom{4}{2}=$ 12. Further, $\beta_{N}$ converges to $\beta=\frac{1}{6}\frac{\alpha}{2-\alpha}$. Therefore (57) converges as $N\to\infty$ to
$\frac{\beta\varphi}{1-(1-\beta)\varphi}$
which is the characteristic function of an exponential random variable with rate $12\beta=\frac{4}{2-\alpha}$. This is the claim. ∎
We can now arrive at the main proof of this section.
Theorem 6.13.
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ and $N(1-\alpha_{N})\rightarrow\infty$. Then for any fixed $t>0$, as $N\rightarrow\infty$, $\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{N,2}>t\mid\mathcal{A}_{N})$ converges to $e^{-\frac{2}{2-\alpha}t}$ in $L^{2}(\mathbb{P}_{{\rm diff}})$.
Proof.
Expanding out the squared $L^{2}(\mathbb{P}_{{\rm diff}})$ distance between $\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{N,2}>t\mid\mathcal{A}_{N})$ and $e^{-2t}$, we obtain
$\mathbb{P}_{{\rm diff}}(N^{-2}(\tau\wedge\tau^{\prime})>t)-2\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{N,2}>t)e^{-\frac{2}{2-\alpha}}+e^{-\frac{4}{2-\alpha}t}$ (58)
from Lemma 6.7.
By Theorem 3.1, the second summand in (58) converges, as $N\to\infty$, to $e^{-\frac{4}{w-\alpha}t}$. It follows that the squared $L^{2}(\mathbb{P}_{{\rm diff}})$ distance in equation (58) is
$\mathbb{P}_{{\rm diff}}(N^{-2}(\tau\wedge\tau^{\prime})>t)-e^{-\frac{4}{2-\alpha}t}+o(1)$ (59)
which converges to 0 by Lemma 6.12. ∎
In particular, in the partial-selfing regime, we have that the classical approximation by averaging over the pedigrees when we begin in distinct individuals is robust to the pedigree.
6.3.2 Pairwise conditional convergence starting from a single individual
We turn now to the case where both sample lineages are taken from the same individual. The results we have already established for $\mathbb{P}_{{\rm diff}}$ greatly facilitate the analysis of $\mathbb{P}_{\rm{same}}$.
Lemma 6.14.
Let $\tilde{b}$ denote the infimum over the time-steps $k$ such that $\hat{X}_{k}\neq\hat{Y}_{k}$. Then
$\mathbb{P}_{\rm{same}}(\tau\wedge\tau^{\prime}>\tilde{b})=\frac{4(1-\alpha_{N})}{4-\alpha_{N}}.$
Furthermore, $\mathbb{E}_{\rm{same}}\left[N^{-2}\tilde{b}\,\Big{|}\,\tilde{b}<\infty\right]$ converges to 00 in distribution.
Proof. Let $U$ denote the number of selfing events of the single occupied individual before splitting. The rate at which selfing events in this individual occur is $\alpha_{N}N^{-1}$. The rate of splitting is $(1-\alpha_{N})N^{-1}$. Therefore the probability of a splitting event before coalescing is $1-\alpha_{N}$. Therefore $\mathbb{P}_{\rm{same}}(U=k)=\alpha_{N}^{k}(1-\alpha_{N})$. That is, $U$ is 00 with probability $1-\alpha_{N}$, else it is a geometric random variable with parameter $1-\alpha_{N}$. In particular, $\mathbb{E}_{\rm{same}}\left[4^{-U}\right]=\frac{4(1-\alpha_{N})}{4-\alpha_{N}}$.
The probability that neither pair coalesces during a selfing event is $\frac{1}{4}$. Therefore
$\mathbb{P}_{\rm{same}}(\tau\wedge\tau^{\prime}>\tilde{b}|U=k)=4^{-k}.$
That $N^{-2}\tilde{b}$ converges to 00 follows from it being geometric with parameter $(1-\alpha_{N})N^{-1}$. ∎
Corollary 6.15.
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$. As $N\to\infty$,
$\mathbb{E}_{\rm{same}}\left[e^{itN^{-2}(\tau\wedge\tau^{\prime})}\right]\to\frac{3\alpha}{4-\alpha}+\left(1-\frac{3\alpha}{4-\alpha}\right)\left(1+it(\frac{4}{2-\alpha})^{-1}\right)^{-1}.$
Proof. We can decompose $\mathbb{E}_{\rm{same}}\left[e^{itN^{-2}(\tau\wedge\tau^{\prime})}\right]$ into where coalescence occurs before the first splitting time $\tilde{b}$. The probability that it occurs before the first split is $\frac{3\alpha_{N}}{4-\alpha_{N}}$ and the time for it to occur, given that we have coalesce converges in distribution to 00 with the time rescaling by Lemma 6.14. Therefore, conditional on $\{\tau\wedge\tau^{\prime}<\tilde{b}\}$, with the $\mathbb{P}_{\rm{same}}$ sampling scheme $\tau\wedge\tau^{\prime}$ converges in distribution to 00.
Similarly we can condition on $\{\tau\wedge\tau^{\prime}>\tilde{b}\}$, where we transition into the initial conditions for the two pairs under $\mathbb{P}_{{\rm diff}}$. Therefore
$\mathbb{P}_{\rm{same}}\Big{(}N^{-2}(\tau\wedge\tau^{\prime})>t\big{|}\tau\wedge\tau^{\prime}>\tilde{b}\Big{)}=\mathbb{P}_{{\rm diff}}(N^{-2}(\tau\wedge\tau^{\prime}+\tilde{b})>t).$ (60)
Again by Lemma 6.14 we know $N^{-2}\tilde{b}$, conditional on being finite, converges to 00 in distribution as $N\to\infty$ so (60) converges, by Theorem 6.13 to $e^{-\frac{4}{2-\alpha}t}$ as $N\to\infty$. In particular, then, $\tau\wedge\tau^{\prime}$ under the law $\mathbb{P}_{\rm{same}}(\cdot|\tau\wedge\tau^{\prime}>\tilde{b})$ converges to an exponential random variable with rate $\frac{4}{2-\alpha}$, who has a characteristic function $(1+it(\frac{4}{2-\alpha})^{-1})^{-1}$.
Combining the two conditional limits gives the result. ∎
Theorem 6.16.
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ and $N(1-\alpha_{N})\rightarrow\infty$. Then for any fixed $t>0$, as $N\to\infty$,
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})\to 2^{-U}e^{-\frac{2}{2-\alpha}t}$
where $\mathbb{P}(U=k)=\alpha^{k}(1-\alpha)$ for $k$ in $\mathbb{Z}_{+}$.
Proof. Let $U_{N}$ denote the number of overlap events that occur in the single occupied individual before this individual undergoes a splitting event and let $\tilde{b}$ denote the time-step of this split. The probability of $k$ overlaps before the first split in the single occupied individual is $\alpha_{N}^{k}(1-\alpha_{N})$. Therefore $U_{N}$ converges in distribution to a random variable $U$ satisfying $\mathbb{P}(U=k)=\alpha^{k}(1-\alpha)$, the same $U$ as in the statement of the theorem.
The key observation is that $\tau^{(N,2)}$ coalesces “instantaneously” with probability $1-2^{-U_{N}}$, else it reaches a state where the two lineages are in distinct individuals and that these two distinct individuals are directly readable from the pedigree, reducing to the case where we started in distinct individuals. Equivalent to the statement is showing that $\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})$ converges to $2^{-U}e^{-\frac{2}{2-\alpha}t}$ in $L^{2}(\mathbb{P}_{\rm{same}})$.
Let $B_{N}$ denote the set where $N^{-2}\tilde{b}<t$. $B_{N}$ converges to one in probability by Lemma 6.14. In particular,
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})=\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N},B_{N})+o(1).$
Further,
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N},B_{N})=2^{-U_{N}}\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N},B_{N},\tau^{(N,2)}>\tilde{b}).$
Therefore
$\mathbb{E}_{\rm{same}}\left[(\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})-2^{-U_{N}}e^{-\frac{2}{2-\alpha}t})^{2}\right]$
$=$ $\mathbb{E}_{\rm{same}}\left[(\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})-2^{-U_{N}}e^{-\frac{2}{2-\alpha}t})^{2}\middle|B_{N}\right]+o(1)$
is equal to
$\mathbb{P}_{\rm{same}}(N^{-2}(\tau\wedge\tau^{\prime})>t)+\mathbb{E}_{\rm{same}}\left[4^{-U_{N}}\right]e^{-\frac{4}{2-\alpha}t}$ (61)
$-2\mathbb{E}_{\rm{same}}\left[4^{-U_{N}}\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t\middle|\tau^{(N,2)}>\tilde{b},B_{N},\mathcal{A}_{N})\right]e^{-\frac{2}{2-\alpha}t}+o(1).$
See that the sum of the first and second summands converges to $\frac{8(1-\alpha)}{2-\alpha}e^{-\frac{4}{2-\alpha}t}$ by Corollary 6.15 and that $\mathbb{E}_{\rm{same}}\left[4^{-U_{N}}\right]=\frac{4(1-\alpha_{N})}{2-\alpha}$. It suffices therefore to show that the third summand converges to the negation of this term.
The key observation is that $4^{-U_{N}}$ is independent to the pairwise coalescence time conditional on the pedigree given that the coalescence time occurs after the split. In particular,
$\mathbb{E}_{\rm{same}}\left[4^{-U_{N}}\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t\big{|}N^{-2}\tau^{(N,2)}>\tilde{b},B_{N})\;\Big{|}\;\mathcal{A}_{N}\right]$
decomposes into a product
$\mathbb{E}_{\rm{same}}\left[4^{-U_{N}}\right]\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\tau^{(N,2)}>\tilde{b},B_{N}).$
$\mathbb{E}_{\rm{same}}\left[4^{-U_{N}}\right]$ may be calculated directly as $\frac{4(1-\alpha_{N})}{4-\alpha_{N}}$ and $\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|N^{-2}\tau^{(N,2)}>N^{-2}\tilde{b},B_{N})$, which converges to $e^{-\frac{2}{2-\alpha}t}$ by Corollary 8.4 and the fact that conditioning on $N^{-2}\tau^{(N,2)}>\tilde{b}$ is equivalent to conditioning on there being a split before coalescence. Therefore the third summand in (61) converges to
$\frac{-8(1-\alpha)}{2-\alpha}e^{-\frac{4}{2-\alpha}t},$
which we established was sufficient for the claim. ∎
The special case of $\alpha=0$ is of interest and is an immediate consequence.
Corollary 6.17.
Suppose $\alpha_{N}\to 0$. Then for any fixed $t>0$, as $N\to\infty$,
$\mathbb{P}_{\rm{same}}\left(N^{-2}\tau^{(N,2)}>t\,\mid\mathcal{A}_{N}\right)\to e^{-t}$
in probability with respect to $\mathbb{P}_{\rm{same}}$.
6.4 Summary of the main proofs
Proof of Theorem 4.3.
The proof of Theorem 4.3 in the partial selfing, limited outcrossing, and negligible-outcrossing regimes are contained in Theorem 6.13, Theorem 6.4, and Corollary 6.6, respectively.
The proof of Theorem 6.13 is obtained by expanding the $L^{2}(\mathbb{P}_{{\rm diff}})$ distance between the conditional survival probability and the survival probability of an exponential random variable. This is performed using a second moment estimate of the conditional survival probability provided by Lemma 6.7. The cross-term and the square of the survival probability of the exponential simplify by Theorem 3.1, reducing the problem to convergence in distribution of the minimum of two conditionally independent copies of the pairwise coalescence time. This convergence is established in Theorem 6.12.
The proof in the limited-outcrossing and negligible-outcrossing regimes are the same, both following from Theorem 6.4. The subgraph of the pedigree generated by following all outcrossings and the random walks thereon are shown to converge in Skorokhod space by Lemma 6.2. This generates a coupling between random walks on the pedigree and random walks on this convergent subgraph. The convergence given by Lemma 6.2 is sufficient to imply that the conditional survival probability converges due to its dependence only on the discrete structure of the subgraph and the lengths of its edges. This sufficiency is demonstrated in Theorem 6.4. ∎
Proof of Theorem 4.4.
The proof of Theorem 4.4 in the $\alpha=0$, $\alpha\in(0,1)$, and $\alpha=1$ regimes are contained in Corollary 6.17, Theorem 6.16, and Lemma 6.5, respectively.
The proof of Theorem 6.16 follows the same $L^{2}(\mathbb{P}_{\rm{same}})$ expansion argument as in the partial-selfing regime of Theorem 4.3, including the same use of a second moment calculation as in Lemma 6.7. The first and third summands in the expansion can be calculated to be the same by showing that the minimum of two conditionally independent realizations of $\tau^{N,2}$ under the "same" sampling configuration is 00 with positive probability, and is otherwise exponential with rate $\frac{4}{2-\alpha}$ using Lemma 6.14. The cross term can be shown to converge to negative of the above sum using independence of the number of selfing events before the two sample lineages split and the pairwise coalescence time conditional on undergoing a split before coalescence, and then also using the fact that the time of the first splitting event converges to 00 with the $N^{2}$ time rescaling by Lemma 6.14.
In the case where $N(1-\alpha_{N})\to 1$, it follows that the conditional coalescence time converges to 00 in distribution by Theorem 3.1 and the conditional Markov inequality, per Lemma 6.5. ∎
7 Discussion
In this work, we have assumed that the pedigree of a population of constant size $N$ is the outcome of a random process of reproduction in which offspring are produced by self-fertilization with probability $\alpha_{N}$. By conditioning on the pedigree and considering pairwise times to common ancestry, we found three different coalescent models in the limit $N\to\infty$, when time is measured in units of $N$ generations (which is $N^{2}$ time-steps in the Moran model). The deciding factor is $N(1-\alpha_{N})$, which may be interpreted either as the rate of outcrossing events along an ancestral lineage or as the expected number of outcrossed offspring in the population each generation. The three models are ‘negligible outcrossing’ which applies when $N(1-\alpha_{N})\to 0$, ‘limited outcrossing’ which applies when $N(1-\alpha_{N})\to\lambda\in(0,\infty)$, and ‘partial selfing’ which applies when $N(1-\alpha_{N})\to\infty$. Negligible outcrossing and limited outcrossing both require $\alpha_{N}\to 1$, while partial selfing includes this as a special case.
Previous population-genetic analyses of selfing populations have assumed a fixed selfing rate, $\alpha_{N}=\alpha$ in our notation, and have implicitly averaged over pedigrees. When this averaging is done, just one coalescent model for all $\alpha\in[0,1]$ emerges in the limit, specifically the coalescent model with partial selfing described by Nordborg and Donnelly (1997) and Möhle (1998). When coalescence is conditioned on the pedigree and $\alpha$ is fixed, our very similar partial-selfing model is obtained but only for $\alpha\in[0,1)$. Assuming a fixed selfing rate of $\alpha=1$ leads to an entirely different model, namely negligible outcrossing.
A more detailed look at how $\alpha_{N}$ approaches $1$ reveals the critical case of limited outcrossing in between negligible outcrossing and partial selfing and characterized by the parameter $N(1-\alpha_{N})\to\lambda\in(0,\infty)$. In this regime, the parts of the population pedigree directly ancestral to the sampled individuals can be replaced by a random graph like the ones previously used to model recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). It is distinguished from these only by what the splitting events in the graph represent and how these are treated in modeling gene genealogies. Each splitting event in the graph corresponds to an outcrossing event in the genealogical ancestry of the sampled individuals. Ancestral lineages tracing back to such a node in the graph make a $50$:$50$ choice of which branch they follow. The rate at which ancestral lineages split is precisely our parameter $\lambda$.
As with the well known Wright-Fisher diffusion and the corresponding standard neutral coalescent model, these new models of coalescence conditional on the pedigree are meant as robust approximations for large populations. We expect, for example, that the Wright-Fisher model with selfing used to produce Figure 1 will have the same three limiting cases we found for the Moran model with selfing. Which of the three is most applicable for a given (necessarily finite) population will depend on $N$ and $\alpha_{N}$. The special case of negligible outcrossing involves the strong predictions of identical gene genealogies across all loci in the genome and zero heterozygosity, and so may be of limited utility. The critical case of limited outcrossing allows gene genealogies to differ across the genome but also predicts zero heterozygosity. Still it is clear from the jumps in probability in Figure 1 that if $N(1-\alpha_{N})$ is not large, insights from the limited-outcrossing model are helpful even when the value of $\alpha_{N}$ would not prevent heterozygosity. On the other hand, both Figure 1 and Figure 8 suggest that if $N(1-\alpha_{N})$ is greater than about $100$, partial-selfing is a better description than limited-outcrossing, even if the outcrossing rate $1-\alpha_{N}$ is quite small.
We chose to use the name “partial selfing" here because in this case the coalescent process conditional on the pedigree is very similar to the previously described partial-selfing model which did not condition on the pedigree (Nordborg and Donnelly, 1997; Möhle, 1998). Our model may be considered an update of the previous one. In both, times to common ancestry for “diff” samples follow the Kingman coalescent process with effective population size $N_{e}=(2-s)N/2$. In our formulation, partial selfing is obtained when $N(1-\alpha_{N})\to\infty$. Similarly we may recall from Figure 8 that as $\lambda\to\infty$ the predictions for “diff” samples under limited outcrossing converge on those for partial selfing. The reason that pedigrees do not constrain “diff” coalescence times under the partial-selfing model is that outcrossing events dominate the ancestry of the population. The same underlying phenomena are at work here as in the standard models of randomly mating populations, where the fact that individuals have two parents with high probability strongly influences the pedigree (Chang, 1999; Derrida et al., 1999; Coron and Le Jan, 2022) resulting in a fast mixing time of ancestral genetic processes given the pedigree (Barton and Etheridge, 2011). Because of this, the conditional and unconditional coalescent processes for these standard population models converge to the same Kingman coalescent process (Tyukin, 2015; Diamantidis et al., 2024).
The key update in our partial-selfing model is the inclusion of the number of generations of selfing up to the first outcrossing event in the recent ancestry of sampled individuals. In our model, a sampled individual has a single realization of the random variable $U$ in Theorem 4.4 or in (2). In the previous model, an average is taken over this distribution then this average, specifically the inbreeding coefficent $F=s/(2-s)$, is applied to every individual. Conditioning on the pedigree produces a coalescent model which aligns better with other work on partial selfing in population and evolutionary genetics. This is particularly true for multi-locus data, which are an explicit concern of conditional coalescent processes because all loci are transmitted through the same pedigree.
For the purposes of discussion, in what follows we take for granted that our conjecture about the coalescent process of a “diff” sample of size $n>2$ in Section 4.2 holds for partial selfing, namely that it is given by a Kingman coalescent process. We further conjecture that a sample of size $n$ containing $2m$ gene copies together as pairs in $m$ individuals and $n-2m$ gene copies each in different individuals, cf. Nordborg and Donnelly (1997), will experience a random number of effectively instantaneous coalescent events, $X_{i}$, now explicitly at locus $i$, then the remaining $n-X_{i}$ lineages will enter the Kingman coalescent process. Each of the $m$ individuals will have its own number of selfing generations before its first outcrossing event, given by independent draws of the random variable $U$. If these outcomes are $\{k_{1},k_{2},\ldots,k_{m}\}$, then $X_{i}\sim\sum_{r=1}^{m}\textrm{Bernoulli}\left({1-2^{-k_{r}}}\right)$, as in (3) and (4). An unlinked locus, $j\neq i$, has the same $\{k_{1},k_{2},\ldots,k_{m}\}$ but a conditionally independent value of $X_{j}$ and an independent realization of the Kingman coalescent process.
The coalescent with partial selfing conditional on the pedigree accurately predicts the variation in the level of inbreeding among individuals which Haldane (1949), Bennett and Binet (1956) and Kimura (1958) highlighted at the population level, and Weir and Cockerham (1973) developed into a theory of identity disequilibrium; recall (1) and (4). But like all coalescent models, the focus is now on the sample. If we write $Y_{i,r}$ for the indicator random variable of coalescence in the recent ancestry of sampled individual $r\in\{1,2,\ldots,m\}$ for the sample of the previous paragraph, then $E[Y_{i,r}]=s/(2-s)$ as in (3) and $\textrm{Var}[Y_{i,r}]$ is given by (1). If we were able to observe the outcomes $y_{i,r}$ for $r\in\{1,2,\ldots,m\}$, then the mean among loci $\bar{y}_{r}$ would be an estimate of $1-2^{-k_{r}}$ and the sample variance of $\bar{y}_{r}$ among individuals would be an estimate of the identity disequilibrium (1).
Note that our partial-selfing model permits another axis of sampling to be considered, namely the sampling of pedigrees. For example, one may imagine a number of pedigrees independently generated under the model and a single individual sampled from each. The inbreeding coefficient $F=s/(2-s)$ and the identity disequilibrium (1) predict the mean and variance of inbreeding among these individuals.
Identity by descent (i.e. the outcomes $y_{i,r}$ above) cannot be observed directly but only inferred based on patterns of identity in state (Pollak, 1987; Rousset, 2002; Uyenoyama, 2024). A number of methods have been developed and applied to estimate inbreeding coefficients. Some have made use of $U$ in (2) or similar distributions to estimate average inbreeding coefficients of populations (Enjalbert and David, 2000; David et al., 2007; McClure and Whitlock, 2012). Others have employed identity disequilibria (Ritland and Jain, 1981; Ritland, 1996; Sweigart et al., 1999; Yang, 2002; Jarne and David, 2008; Szulkin et al., 2010; Vieira et al., 2013) or runs of homozygosity (Franklin, 1977; Yi et al., 2022; Zeitler and Gilbert, 2024) to estimate individual inbreeding coefficients. Still others have taken up the specific problem of estimating the number of generations back to the most recent outcrossing event for each individual together with the selfing rate (Gao et al., 2007; Wang et al., 2012; Redelings et al., 2015).
Of these empirical works, the ones which appeal to the previous, unconditional coalescent model with partial selfing for justification do so with respect to the properties of “diff” samples, i.e. that they follow a Kingman coalescent with diploid effective size $N_{e}=(2-s)N/2$. Again, this is also a property of the conditional model we have described. But the coalescent model with partial selfing conditional on the pedigree further includes rigorous justification for the modeling of individual ancestries of selfing already done in these works. For example, in the inference method for infinite-alleles data developed by Redelings et al. (2015), individual ancestries are precisely those of the conditional model. Then, cf. Figure 1 in Redelings et al. (2015), the allelic states of the distinct ancestral lineages which emerge from these random outcomes at a locus, in our notation the $n-X_{i}$ lineages described above, are given by the Ewens sampling formula (Ewens, 1972) with the mutation parameter for a Kingman coalescent with diploid effective size $N_{e}=(2-s)N/2$.
The existence of the limited-outcrossing coalescent model, with its simple structure of the random graph, can aid in the interpretation of multi-locus genetic data even if its extreme prediction of zero heterozygosity of individuals is not met. Each branching event in the ancestral graph corresponds to an outcrossing event between ancestral individuals in the pedigree of the sample, in the case that these are rare, i.e. with rate proportional to $1/N$ in the model. Forward in time and to the extent that the two ancestral individuals differ genetically, each such event would create a potentially novel recombinant inbred line. The partial-selfing model approximates this when the selfing rate is high, except that some of these new lines may not become fully inbred.
Samples from predominately selfing populations are often dominated by individuals homozygous for one of only a few distinct mutli-locus genotypes, together with some recombinant inbred or partially inbred individuals (Bonnin et al., 2001; Bakker et al., 2006; Bomblies et al., 2010; Hartfield et al., 2017; Jullien et al., 2019). The Medicago truncatula data for $24$ unlinked loci in a sample of $200$ individuals presented and analyzed by Siol et al. (2008) illustrates this. They found four common multi-locus genotypes present in $76$, $34$, $25$ and $17$ individuals in the sample. A further $21$ recombinant inbred genotypes, resulting from relatively ancient crosses between the common types, were distributed among $39$ individuals in the sample. At lowest frequency in the sample were $9$ recombinant partially inbred (heterozygous) genotypes present in single individuals, presumably with relatively recent outcrossing events in their ancestries.
We have analyzed a simple population model in some detail to understand how the population pedigree shapes the coalescent process for all loci in a genome. The model includes just reproduction and selfing. There is no selection, no mutation, no recombination, no population structure, and no change in the population size over time. We expect that neutral mutation and changes in population size could be included the same way they are included in coalescent models which do not condition on the pedigree. So, for example, applications of coalescent theory to “diff” samples under partial selfing which infer the ancestry of populations, such as those in St. Onge et al. (2011), Brandvain et al. (2013), Beissinger et al. (2016), are as well justified under our conditional partial-selfing model as they are under the unconditional model of Nordborg and Donnelly (1997).
It would not be as straightforward to incorporate selection or population structure because these can have their own effects on the pedigree. Some simulations demonstrating this can be found in Kuo and Avise (2008), Wakeley et al. (2016) and Wilton et al. (2017). It is an open question how recombination and coalescence operate conditional on the population pedigree. We have dealt with free recombination here. Theorem 4.3 and Theorem 4.4 are about the distribution of conditionally independent coalescence times given the pedigree. For the unconditional or pedigree-averaged case, Kogan et al. (2023) recently considered the full range of recombination rates and how these might be assumed to scale with $N$ in an analysis of correlations of coalescence times in samples of size two at two loci. It may be that different results for recombination and coalescence will be obtained in the conditional case depending on how recombination scales with $N$.
Finally, selfing is just one form of inbreeding. The importance of modeling and disentangling the sources of inbreeding is well appreciated in the unconditional or pedigree-averaged case (Uyenoyama, 1986; McClure and Whitlock, 2012; Jarne and David, 2008; Zeitler and Gilbert, 2024). It would be of interest to understand the pedigree effects of inbreeding generally in the conditional case. Campbell (2015) and Severson et al. (2019) studied expected coalescence times for given rates of mating between cousins of varying degree, including sibs or “00-degree cousins”. Severson et al. (2021) and Cotter et al. (2021) extended these models and applied the separation-of-timescales result of Möhle (1998) to obtain the limiting distributions of coalescence times in the unconditional case. Based on the results we presented here, we would expect to see some effects of the population pedigree in conditional coalescent processes for these and other forms of inbreeding.
Acknowledgements
This work was supported by National Science Foundation grants DMS-2152103 and DMS-2348164.
Conflicts of interest
The authors declare no conflicts of interest.