### v2

# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

Abstract
We introduce a stochastic model of a population with overlapping generations and arbitrary levels of self-fertilization versus outcrossing. We study how the global graph of reproductive relationships, or population pedigree, influences the genealogical relationships of a sample of two gene copies at a genetic locus. Specifically, we consider a diploid Moran model with constant population size $N$ over time, in which a proportion of offspring are produced by selfing. We show that the conditional distribution of the pairwise coalescence time at a single locus given the random pedigree converges to a limit law as $N$ tends to infinity. The distribution of coalescence times obtained in this way predicts variation among unlinked loci in a sample of individuals. Traditional coalescent analyses implicitly average over pedigrees and generally make different predictions. We describe three different behaviors in the limit depending on the relative strengths, from large to small, of selfing versus outcrossing: partial selfing, limited outcrossing, and negligible outcrossing. In the case of partial selfing, coalescence times are related to the Kingman coalescent, similar to what is found in traditional analyses. In the case of limited outcrossing, the retained pedigree information forms a random graph, with coalescence times given by the meeting times of random walks on this graph. In the case of negligible outcrossing, which represents complete or nearly complete selfing, coalescence times are determined entirely by the fixed times to common ancestry of diploid individuals in the pedigree.
1 Introduction
The organismal genealogy of all individuals in a population for all time, also known as the population pedigree, has not traditionally been of concern in population genetics. But simulation studies, beginning with Ball et al. (1990) and including Kuo and Avise (2008), Wakeley et al. (2016) and Wilton et al. (2017), have suggested that the population pedigree can be important in structuring genetic variation. Of course, some of the earliest work in population genetics concerned how pedigrees affect patterns of variation, specifically identity by descent (Wright, 1921a, b, c, d, e, 1922). But the pedigrees in these early works were either ones laid down by regular systems of mating—see Lachance (2009) for a recent treatment—or pedigrees of arbitrary structure extending back a limited number of generations. These latter, relatively small pedigrees were assumed to be embedded in an infinite population with global properties, such as Hardy-Weinberg equilibrium and no identity by descent, and within which there was no explicit accounting of individuals or their relationships.
Finite population models including coalescent models are models of identity by descent. But population pedigrees did not figure in their development. The approach instead was to average over pedigrees by averaging over the outcomes of reproduction in each generation. See Diamantidis et al. (2024) for more about this history and a discussion of the roles of pedigrees in forward-time versus backward-time models. Averaging over pedigrees is usually done implicitly, as in the statement that two ancestral lines coalesce with probability $1/(2N)$ each generation under the diploid monoecious Wright-Fisher model. It is done explicitly, for example, in equation (4.1) in Kingman (1982) for the coalescence probability under Cannings’ haploid exchangeable model (Cannings, 1974). Predictions for diploid organisms based on this traditional approach are marginal to the population pedigree.
Coalescent theory describes the ancestry of gene copies or alleles in a sample of individuals from a population. Coalescent models are often applied by comparing predicted distributions of genetic variation at a single locus to data from multiple unlinked loci in a sample of individuals. This is done for example in analyses of the site-frequency spectrum (Tajima, 1989; Braverman et al., 1995; Fu, 1995; Achaz, 2009; Gutenkunst et al., 2009; Gao and Keinan, 2016) or when likelihoods are multiplied across unlinked loci (Watterson, 1985; Wakeley, 1999; Nielsen, 2000; Adams and Hudson, 2004; Wang and Hey, 2010; Lohse et al., 2011; Costa and Wilkinson-Herbots, 2017, 2021). But a sample of individuals has a single shared pedigree, fixed in time by past outcomes of reproduction. Further, in large populations with substantial outcrossing, the pedigree of any sample of individuals quickly spreads out to encompass the whole population (Chang, 1999; Derrida et al., 1999; Lachance, 2009; Barton and Etheridge, 2011). So it is important to know whether any specific features of the population pedigree might affect our predictions about genetic variation.
Coalescent models conditional on the pedigree give the correct sampling structure for multiple loci in a sample of individuals. In particular, unlinked loci represent conditionally independent realizations of the process of genetic transmission within the same fixed pedigree. Methods such as multiplying likelihoods across unlinked loci are obviously justified when conditioning on the pedigree. Modeling unlinked loci as independent draws from the distributions of traditionally formulated coalescent models implicitly assumes that each locus has its own population pedigree.
So far there have been only two analytical treatments of coalescence conditional on the population pedigree. Tyukin (2015) proved that the standard neutral coalescent model (Kingman, 1982; Hudson, 1983; Tajima, 1983) can still be used when coalescence is conditioned on the pedigree, for populations which would be described by that model under the traditional approach. But it is not true in general that extensions of the standard neutral coalescent model are similarly applicable to coalescent processes conditional on pedigrees generated under the corresponding population models. Diamantidis et al. (2024) proved this for populations in which very large families occur with some frequency, and went on to describe the resulting conditional coalescent models, in which predictions about genetic variation depend on the fixed times and sizes of large families.
In this work, we consider how coalescence conditional on the population pedigree is affected by autogamy or self-fertilization (hereafter just ‘selfing’). Modes of reproduction are of course fundamental aspects of evolution (Charlesworth and Wright, 2001; Glémin et al., 2019). Selfing is when two gametes produced by meiosis by the same individual unite to form a zygote. This type of inbreeding has been especially important in the evolution of plants (Jain, 1976; Uyenoyama et al., 1993; Charlesworth, 2006; Olsen et al., 2021). The common alternative to selfing in both plants and animals is outcrossing, in which zygotes are formed by the union of two gametes from different individuals. We consider just these two possibilities, and characterize populations by the probability individuals are produced by selfing. We describe three distinct kinds of conditional coalescent models which hold for different magnitudes of the selfing probability.
These models exist in the limit as the population size $N$ tends to infinity in the same way that the Wright-Fisher diffusion and the corresponding standard neutral coalescent model exist in this limit (Ewens, 2004). We use $\alpha_{N}$ to denote the selfing probability, with explicit dependence on $N$ so we can control the importance of selfing in the model. Our results for coalescence in large populations conditional on the pedigree depend strongly on the relative magnitude of $\alpha_{N}$, specifically how $\alpha_{N}$ changes as $N\to\infty$. Most previous analyses, in addition to not conditioning on the pedigree, have assumed a fixed selfing probability. This probability is often denoted $s$, though various notations have been adopted; e.g. $l$ in Haldane (1924), $h$ in Wright (1951), $\beta$ in Pollak (1987) and $\alpha$ in Karlin (1968a, b). As Uyenoyama (2024) recently emphasized, some authors use $s$ for the entire selfing probability while others use it to measure selfing in excess of a background probability of $1/N$ due to random mating. Our model includes these as special cases, namely if $\alpha_{N}=s$ or $\alpha_{N}=s+(1-s)/N$, respectively. Both of these fixed-$s$ models make the same predictions in the limit and both have been called ‘partial self-fertilization’ or ‘partial selfing’.
We use $\alpha\in[0,1]$ to denote the limit of $\alpha_{N}$ as $N\to\infty$. Conditioning coalescence on the pedigree reveals three markedly different behaviors near the boundary $\alpha=1$, depending on how quickly $\alpha_{N}\to 1$. The critical case, which we call ‘limited outcrossing’, is when $1-\alpha_{N}$ is of order $1/N$. Then the conditional coalescent process is characterized by a system of coalescing random walks on a random ancestral graph like the ones previously described for recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). If outcrossing is even less frequent, for example $1-\alpha_{N}$ of order $1/N^{2}$, the limiting conditional coalescent process is fully determined by the random pedigree of the population, which becomes just a Kingman coalescent tree connecting diploid individuals. We call this ‘negligible outcrossing’. If outcrossing is more frequent than in the critical case, for example $1-\alpha_{N}$ of order $1/\sqrt{N}$, the limiting conditional coalescent model resembles the pedigree-averaged or unconditional coalescent model of Nordborg and Donnelly (1997) and Möhle (1998). Here we adopt the previous term ‘partial selfing’. The latter model holds for all $\alpha\in[0,1]$ as long as the rate of approach to $\alpha=1$ is not too fast, whereas limited outcrossing and negligible outcrossing hold just for $\alpha_{N}\to 1$.
In the next two sections, we review previous works which generally use $s$ for the fixed selfing probability, take $s\in[0,1]$, and do not condition on the population pedigree. We will reserve our notation $\alpha_{N}$ for Section 1.3 and later. In Section 1.3, we explain how previous population-genetic models of selfing in fact average over pedigrees. In Sections 1.1 and 1.2 we use $s$ for continuity with what one finds in much of the evolutionary and population-genetic literature.
1.1 Prevalence and evolutionary-genetic consequences of selfing
Although the ability to self-fertilize is rare among animals (Jarne and Auld, 2006; Avise and Mank, 2009; Escobar et al., 2011) it is found in other taxa including eukaryotic microbes and marine invertebrates (Sasson and Ryan, 2017; Yadav et al., 2023) and is common in plants (Abbott and Gomes, 1989; Hartfield et al., 2017; Teterina et al., 2023). About $10$-$15$% of plant species are predominantly selfing (Wright et al., 2013), as are two of the best-studied genetic model species, Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). The empirical distribution of selfing probabilities among plant species is markedly bimodal, with relatively fewer species having $s\in(0.2,0.8)$ (Schemske and Lande, 1985; Vogler and Kalisz, 2001; Goodwillie et al., 2005). Whether selfing is favored over outcrossing or vice versa in a given species depends on a number of ecological and evolutionary factors (Fisher, 1941; Moran, 1962; Lande and Schemske, 1985; Charlesworth, 2006; Kamran-Disfani and Agrawal, 2014).
Selfing is a form of inbreeding which has three important consequences for the distribution of genetic variation. First, selfing increases single-locus homozygosity via increased identity by descent (Haldane, 1924; Rousset, 2002). In terms of Wright’s individual inbreeding coefficient $F$ (Wright, 1931, 1951), what Haldane (1924) showed was that the reduction in the proportion of heterozygous individuals in the population at equilibrium is given by $F=s/(2-s)$. A simple coalescent interpretation is that $F$ is the probability that the two alleles or gene copies at a locus in an individual coalesce in the recent ancestry of that individual, rather than being separated into different individuals by an outcrossing event (Nordborg and Donnelly, 1997).
Second, selfing reduces the effective rate of recombination (Bennett and Binet, 1956; Weir and Cockerham, 1973; Golding and Strobeck, 1980; Vitalis and Couvet, 2001). Nordborg (2000) provided an interpretation in terms of coalescence. Backward in time, a recombination event places ancestral genetic material onto two chromosomes. These are necessarily in the same parental individual which might itself have been produced by selfing. Then the same coefficient $F$ is the probability that the recombination event gets healed or repaired by a coalescent event within that parental individual’s recent ancestry. So only a fraction $1-F=2(1-s)/(2-s)$ of recombination events is observable.
Third, selfing generates non-random associations of diploid genotypes at different loci, even if the loci are unlinked (Haldane, 1949; Bennett and Binet, 1956; Kimura, 1958; Narain, 1966; Cockerham and Weir, 1968, 1977, 1983, 1984; Smith, 2023; Uyenoyama, 2024). Weir and Cockerham (1973) partitioned this effect into a sum of two terms, one due to linkage, which is equal to zero for unlinked loci, and one due to selfing, which is called the ‘identity disequilibrium’ and is equal to the theoretical variance of $F$ among individuals in the population:
$\sigma^{2}_{F}=\frac{4s(1-s)}{(2-s)^{2}(4-s)}.$ (1)
For reference, the quantity (1) is denoted $\sigma^{2}_{F_{1}}$ in Weir and Cockerham (1973, p. 260).
A coalescent interpretation of identity disequilibrium is as follows. Let $Y_{i}$ be an indicator variable, equal to $1$ if the two gene copies at locus $i$ in a particular individual coalesce in its recent ancestry, and equal to 00 otherwise. The recent ancestry of a randomly sampled individual is given by the number of generations of selfing in its ancestral line before the first outcrossing event. Weir and Cockerham (1973) note that this idea originates with Cockerham and Rawlings (1967). For an individual with $k$ generations of selfing in its immediate ancestry, the probability the two gene copies at a locus remain distinct (i.e. do not coalesce) in any of these generations is $2^{-k}$. If $U$ denotes the random number of selfing generations for an individual,
$\mathbb{P}(U=k)=s^{k}(1-s)\qquad k=0,1,2,\ldots$ (2)
and the coefficient $F$ can be seen as the average
$F\coloneq\mathbb{E}\left[Y_{i}\right]=\sum_{k=0}^{\infty}\left(1-2^{-k}\right)s^{k}(1-s)=\frac{s}{2-s}.$ (3)
The covariance of $Y_{i}$ and $Y_{j}$ for two unlinked loci, which are conditionally independent given $k$, is
$\mathrm{Cov}{\left[Y_{i},Y_{j}\right]}=\sum_{k=0}^{\infty}\left(1-2^{-k}\right)^{2}s^{k}(1-s)-\left(\frac{s}{2-s}\right)^{2},$ (4)
and this is identical to (1) because $\sigma^{2}_{F}\coloneq\mathrm{Var}{\left(Y_{i}\right)}$ is also given by (4). Interpreting identity disequilibria as covariances is sensible because they occur between loci.
In anticipation of later results conditional on the population pedigree, note that (2) can be interpreted in two different ways. A given pedigree has fixed values of $k$ for every individual in the population. Following Haldane (1949), Weir and Cockerham (1973) and others, (1), (3) and (4) can be interpreted as averages over individuals in an effectively infinite population. Alternatively, (2) can be seen as the prior distribution for a single individual in a given population. In this case, (1), (3) and (4) are seen as averages over all possible population pedigrees. One reason this distinction is important is that while (1) or (4) may be greater than zero, the covariance of $Y_{i}$ and $Y_{j}$ for two unlinked loci in a single individual conditional on the population pedigree must be equal to zero.
Owing to these three effects—increased homozygosity, reduced recombination, and identity disequilibrium—partially selfing populations follow different evolutionary paths than randomly mating populations (Charlesworth et al., 1990, 1993; Uyenoyama and Waller, 1991a, b, c; Lande et al., 1994; Nordborg et al., 1996; Roze and Lenormand, 2005; Roze, 2016; Abu Awad and Roze, 2020). The evolution of selfing also involves a variety of ecological phenomena and associated morphologies (Baker, 1955; Stebbins, 1957; Ornduff, 1969; Jain, 1976; Cutter, 2019). We note that some but not all evolutionary studies have explicitly modeled variation in the number of selfing generations in the recent ancestries of individuals (Cockerham and Rawlings, 1967; Campbell, 1986; Kelly, 1999a, b, 2007; Lande and Porcher, 2015), given here under neutrality by (2).
1.2 Standard neutral models with a fixed selfing probability but no pedigree
Our goal in this work is to understand how selfing affects the sampling structure of gene genealogies under neutrality when the process of coalescence is conditioned on the pedigree of the population. For simplicity we use a diploid Moran model of reproduction (Moran, 1958, 1962; Linder, 2009; Coron and Le Jan, 2022), which we describe in Section 2. Here we review some relevant previous theoretical treatments which do not include pedigrees and which assume Wright-Fisher reproduction (Fisher, 1930; Wright, 1931).
Both the Wright-Fisher model and the Moran model assume a single well mixed population of constant in size $N$ without selection. In the usual simplest application, i.e. to diploid, monoecious, randomly mating organisms, selfing occurs with probability $1/N$. When $N$ can be assumed to be large, both converge to the Wright-Fisher diffusion with its corresponding coalescent process (Ewens, 2004), which exist in the limit $N\to\infty$ with time measured in units proportional to $N$ generations. Arbitrary levels of selfing have been incorporated by assuming a fixed selfing probability of $s\in[0,1]$ in the limit $N\to\infty$. Based on the considerations of Section 1.1, it may be that selfing species are even less likely than outcrossing ones to meet the assumptions of these standard models of population genetics, but they provide a starting point for studying more complicated populations.
Pollak (1987) took a forward-time approach to studying, among other things, the equilibrium probabilities of identity by descent of two gene copies in the same individual or in different individuals, when $s\gg N^{-1}$. Let us call these probabilities $\Phi_{1}$ and $\Phi_{2}$, respectively. For large $N$, Pollak (1987) found that the expression for $\Phi_{2}$ was identical to that for a randomly mating population, if $N_{e}=N/(1+F)=(2-s)N/2$ is used in place of $N$. As $s$ increases from 00 to $1$, $N_{e}$ decreases from $N$ to $N/2$, effectively taking the value of $\Phi_{2}$ from that for a diploid (randomly mating) population to that for a haploid population. Pollak (1987) identified this as a separation-of-time-scales result, stemming from the assumption $s\gg N^{-1}$, and suggested that the two-times-scales diffusion approximation of Ethier and Nagylaki (1980) could be applied to prove convergence to the Wright-Fisher diffusion with this value of $N_{e}$. The separation of time scales can be seen in the relationship Pollak (1987) found between $\Phi_{1}$ and $\Phi_{2}$, namely that $\Phi_{1}=F+(1-F)\Phi_{2}$, which may best be understood in terms of subsequent descriptions of the following coalescent process.
[…]
It was in this context of partial selfing that Möhle (1998) proved an important general result (Möhle, 1998, Lemma 1, Theorem 1) for Markov processes with two time scales. Owing to the duality relation between the coalescent and diffusion models (Möhle, 1999) this result can be seen as the backward-time counterpart to the result of Ethier and Nagylaki (1980) cited by Pollak (1987). For the selfing model with fixed $s$ as $N\to\infty$, this gave rigorous justification to the model of Nordborg and Donnelly (1997): whenever a pair of lineages is in the same individual, they coalesce with probability $F=s/(2-s)$ instantaneously, otherwise ending up in different individuals with probability $1-F$, while samples from different individuals obey the standard neutral coalescent process with time in units of $(2-s)N$ generations (Möhle, 1998).
1.3 Accounting for the pedigree of the population is important
The coalescent models of partial selfing in Fu (1997), Nordborg and Donnelly (1997) and Möhle (1998) average over the process of reproduction and thus implicitly over the pedigree. Clearly, one consequence of conditioning on the pedigree should be that each sampled individual has its own realized number of generations of selfing before the first outcrossing event in its ancestry. This, as we will show, is a single draw from the distribution (2). Previous approaches instead used the average $F$ in (3). The averaging procedure is illustrated by the matrix $A^{m}$ in Möhle (1998, p. 496) which gives the $m$-generation transition probabilities of the fast process. For example,
$\left(A^{m}\right)_{32}=(1-s)\frac{1-(s/2)^{m}}{1-s/2}$
is the probability that the ancestors of two gene copies currently in the same individual remain distinct and are in two different individuals in past generation $m$. This can be expressed as a sum over pedigrees using the distribution (2), namely
$\left(A^{m}\right)_{32}=\sum_{k=0}^{m-1}\left(\frac{1}{2}\right)^{k}s^{k}(1-s).$ (5)
Lemma 1 and Theorem 1 of Möhle (1998) use the limit of the fast process, $P\coloneq\lim_{m\to\infty}A^{m}$, in which (5) converges to $2(1-s)/(2-s)=1-F$, or $\mathbb{E}\left[1-Y_{i}\right]$ for the indicator random variable $Y_{i}$ defined in Section 1.1.
Partial selfing is just one of three possibilities when the coalescent process is conditioned on the pedigree. Here we reintroduce our notation, in which $\alpha_{N}$ denotes the selfing probability in a population of size $N$, and $\alpha$ denotes its value in the limit $N\to\infty$. When selfing is the primary mode of reproduction, an entirely new sort of model emerges in the limit, one which is completely invisible in the unconditional pedigree-averaged process. This novel limit process occurs when the outcrossing probability $1-\alpha_{N}$ is of order $1/N$. It is parameterized by $\lambda\coloneq\lim_{N\to\infty}N(1-\alpha_{N})$, which we note is the same way mutation rates are assumed to scale in the Wright-Fisher diffusion and corresponding standard neutral coalescent model (Ewens, 2004) and how migration rates are scaled in the structured coalescent model (Takahata, 1988; Notohara, 1990; Herbots, 1997). In the limit we consider, $\lambda$ is the instantaneous rate of outcrossing events per ancestral lineage.
Figure 1 shows simulation results of pedigrees and pairwise coalescence times for the model of selfing used by Möhle (1998), which corresponds to $\alpha_{N}=\alpha$ constant in our model. Specifically, for each of three selfing probabilities, $50$ independent pedigrees were simulated and for each of these two different individuals were sampled at random. Colored lines, one for each pedigree-plus-sample, depict the cumulative distribution function (CDF) of the coalescence time for two gene copies, one from each of the two sampled individuals. On the scale used in Figure 1, the corresponding expectation for the pedigree-averaging model in all three cases would be the CDF of an exponential distribution with parameter $(2-\alpha)^{-1}$. Many of the CDFs in Figure 1(a) have something close to this shape, but those in Figure 1(b) and Figure 1(c) look nothing like it. Instead they are characterized by random jumps in probability at random times. For any one population, these times are fixed and determine the predicted distribution of coalescence times among unlinked loci.
   Refer to caption   Refer to caption by the method in Wakeley et al. (2012).
Figure 1: Cumulative distribution functions of the coalescence times for a sample of size $n=2$ on each of $50$ population pedigrees simulated under the Wright-Fisher model with partial selfing and $N=1000$ individuals. Panels (a), (b) and (c) show the results for three different values of the probability of selfing, respectively, for $\alpha\in\{0.9,0.99,0.999\}$. Coalescence probabilities in each generation given the pedigree and the sampled individuals were calculated exactly, up to numerical precision
The deviations from the predictions of the unconditional coalescent model displayed in Figure 1 are not due to the relatively small population size used in these simulations ($N=1000$). Instead they illustrate that what we will prove is true in the limit $N\to\infty$ with $N(1-\alpha_{N})\to\lambda$ holds even for such small populations. The value $\alpha=0.99$ in Figure 1(b) is similar to estimates of the selfing probabilities for the model species Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). Of course, $N=1000$ would be a very small population size for either of these species, though it may be realistic for some local populations. In any case, species or populations for which $N(1-\alpha)$ is not very large will have genetic ancestries very unlike what is expected under the unconditional or pedigree-averaging model. Instead, they are captured by the conditional coalescent model with ‘limited outcrossing’ described in Section 4.
1.4 Plan of the present work
Our analysis is based on a diploid Moran model of reproduction with constant population size $N$ and selfing probability $\alpha_{N}$, described in detail in Section 2. We focus on the coalescence time for a sample of size $n=2$ at a single locus, which already reveals the non-trivial effects of the pedigree on gene genealogies. In the case of partial selfing, we prove convergence in distribution of the coalescence time via an analysis of two independently assorting or unlinked loci on the population pedigree, cf. Birkner et al. (2013) and Diamantidis et al. (2024). In the case of limited outcrossing, we take a different approach to establish the conditional limit. Here we focus on the set of diploid individual ancestors of the sample and prove that this converges weakly to an ancestral graph.
The organization of the paper is as follows. In Section 2, after detailing our Moran model, we make precise the pedigree and the gene genealogy. Our results for the unconditional and the conditional distributions of pairwise coalescence times are presented in Section 3 and Section 4, respectively. We briefly discuss the case $n>2$ and offer a conjecture in Section 4.1. In Section 5, we establish some fundamental properties of the scaling limit under limited outcrossing and of the associated ancestral graph. We present the proofs of our results in the Appendix, Section 7.
2 A discrete-time diploid Moran model with selfing
We consider a diploid, monoecious, panmictic (well mixed and randomly mating) population of constant size $N$. Individuals are labeled $I=\{1,...,N\}$. The population has overlapping generations and changes at discrete time-steps. At each time-step, a randomly chosen individual dies and is replaced by an offspring produced in one of two ways. With probability $1-\alpha_{N}$ the offspring has two distinct parents; with probability $\alpha_{N}$ the offspring has exactly one parent. The parent(s) are chosen uniformly at random from the previous time-step. In the case of two parents, these are chosen without replacement. The reproduction events at different time-steps are independent and identically distributed.
Explicitly, for each non-negative integer $k\in\mathbb{Z}_{+}=\{0,1,2,3,...\}$, the reproduction event between consecutive time-steps $k$ and $k+1$ in the past is as follows:
  1. 1.
    (distinct parents) With probability $1-\alpha_{N}$, a single offspring is produced by outcrossing. A triplet containing the two parents and one offspring $(\pi_{k,1},\,\pi_{k,2},\,\varepsilon_{k})\in I^{3}$ is chosen uniformly at random with $\pi_{k,1}\neq\pi_{k,2}$ (i.e. the parents are distinct). The offspring $\varepsilon_{k}$ is an individual in time-step $k$, while the offspring’s parents are the individuals in time-step $k+1$ with labels $\pi_{k,1}$ and $\pi_{k,2}$.
    The offspring has two genes copies, one inherited from each parent. Genetically, the offspring is produced according to Mendel’s laws, which means each of the two gene copies in a parent is equally likely to be the one transmitted to the offspring. An example outcrossing event in a population of size $N=7$ is depicted below.
    ($\varepsilon_{k}=3$)   time-step $k$($\pi_{k,1}=2,\pi_{k,2}=4$)  time-step $k$$+$$1$
  2. 2.
    (selfing) With probability $\alpha_{N}$, a parent-offspring pair with labels $(\pi_{k},\,\varepsilon_{k})\in I^{2}$ is chosen uniformly at random without replacement. The offspring $\varepsilon_{k}$ in time-step $k$ is the offspring of a single individual (the parent) with label $\pi_{k}$ at time-step $k+1$.
    As above, the offspring is produced according to Mendel’s laws which means each of the two gene copies in the parent is independently and equally likely to be inherited by each of those of the offspring. An example selfing event in a population of size $N=7$ is depicted below.
    ($\varepsilon_{k}=3$)    time-step $k$($\pi_{k,1}=\pi_{k,2}=4$)   time-step $k$$+$$1$
Reproduction events in different time-steps are independent. Note that at time-step $k$, the offspring $\varepsilon_{k}$ is a new individual introduced and it replaces the one with the same label in the previous step $k+1$ (death-birth for the label $\varepsilon_{k}$). The remaining $N-1$ individuals in time-step $k$ are the same individuals as those in time-step $k+1$. Hence this model has overlapping generations.
Remark 2.1.
When $\alpha_{N}=0$, our model is exactly the one introduced in Coron and Le Jan (2022). When $\alpha_{N}$ is equal to a fixed $p$ in $[0,1)$, our model is exactly the one in Linder (2009).
2.1 The pedigree as important partial information of the population
The above population dynamics generates a random graph, which we will call $\mathcal{G}_{N}$, with vertex set $I\times\mathbb{Z}_{+}$ (the $N$ individuals at all time-steps) and directed edges joining each individual with its parent(s). This random graph $\mathcal{G}_{N}$ is the population pedigree. For the moment, let us consider this pedigree without specifying outcomes of genetic transmission.
Between consecutive time-steps with two distinct parents, the pedigree has one edge from offspring $(\varepsilon_{k},k)$ to parent $(\pi_{k,1},k+1)$, one from offspring $(\varepsilon_{k},k)$ to parent $(\pi_{k,2},k+1)$, and $N-1$ single edges for the same individual from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\varepsilon_{k}\}$. The portion of the pedigree corresponding to the example outcrossing event above is
[…]
Between consecutive time-steps with selfing, the pedigree has a double edge from $(\varepsilon_{k},k)$ to $(\pi_{k},k+1)$, and again $N-1$ single edges from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\varepsilon_{k}\}$. The portion of the pedigree corresponding to the example selfing event above is
[…]
Hence, the offspring always has 2 edges coming out of it which trace upward (towards the past). All the other $N-1$ individuals have a single edge.
Repetition of this process with these two possibilities for each time step backward into the indefinite past results in a single realized pedigree of the population. Patterns of genetic ancestry at each locus are outcomes of Mendelian transmission conditional on this one pedigree. Unlinked loci are transmitted independently through the pedigree. In Figure 2, we illustrate a realization of both the population dynamics with genetic transmission and the corresponding pedigree for the first 4 generations in the past for a population of size $6$.
Figure 2: A realization of our diploid Moran process with $N=6$ individuals including the genetic transmission events at one locus (left image) and the corresponding pedigree (right image). Also on the left, two gene copies $(X_{0},Y_{0})=(2,9)$ are sampled in the present time-step 00, and their lineages are highlighted by the solid dots. Namely, $\{(X_{k},Y_{k})\}_{k=0}^{4}=\{(2,9),(2,9),(2,5),(6,5),(4,4)\}$. These two lineages coalesce in time-step $4$ because $X_{4}=Y_{4}$ but $X_{k}\neq Y_{k}$ for $k=0,1,2,3$.
2.2 Pairwise coalescence time: unconditional versus conditional on the pedigree
In our population model each of the $N$ individuals has two copies of each genetic locus. Suppose we number these $J=\{1,...,2N\}$ such that the individual with label $i\in I$ has gene copies $12i-$ 1 and $2i$. Broadly, our aim is to describe the statistical properties of random samples of gene copies from $J$ by following their ancestral genetic lineages backward in time.
Suppose we sample two gene copies $X_{0}$ and $Y_{0}$ without replacement from the $2N$ present at time $k=0$. Let $X_{k}$ and $Y_{k}$ be their ancestral lineages $k$ time-steps in the past. Then $(X_{k})_{k\in\mathbb{Z}_{+}}$ and $(Y_{k})_{k\in\mathbb{Z}_{+}}$ are two correlated Markov chains with state space $J$ that have the same transition probabilities. Our main object of study, for sample size $n=2$, is
$\tau^{(N)}:=\inf\{k\in\mathbb{Z}_{+}:\,X_{k}=Y_{k}\},$ (8)
the (pairwise) coalescence time of the sample of two genes. The left image in Figure 2 shows a realization of the population process in which $\tau^{(6)}=4$.
The value of $\tau^{(N)}$ is completely determined by the population process. However, knowing the pedigree is not enough to determine the exact value of $\tau^{(N)}$, due to Mendelian randomness. For instance, even if we know the pedigree is exactly the one on the right in Figure 2 and we know that $X_{0}=2$, we do not know what $X_{3}$ is: $X_{3}$ can be in one of the genes $\{5,6,7,8\}$ with equal probability.
3 Unconditional distribution of pairwise coalescence time
[…]
${\text{diff}=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}&\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\bullet}&\end{pmatrix}\,\text{same}=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}&{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\bullet}\end{pmatrix},\ \text{coal}=\begin{pmatrix}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\bullet}&\end{pmatrix}}$
The one-step transition matrix $\mathbf{\Pi}_{N}$ of this Markov process, under our diploid Moran model, is
$\mathbf{\Pi}_{N}\,:=\,$ diff same coal diff $1-\frac{2}{N(N-1)}$ $\frac{1}{N(N-1)}$ $\frac{1}{N(N-1)}$ same $(1-\alpha_{N})\frac{1}{N}$ $\frac{N-1}{N}+\alpha_{N}\frac{1}{2N}$ $\alpha_{N}\frac{1}{2N}$ coal 00 00 $1$
For $q\in\{\text{diff},\,\text{same},\,\text{coal}\}$, we let $\mathbb{P}_{q}$ be the probability when the initial state (i.e. the sampling configuration) is $q$. We will be interested in the probability the coalescence time exceeds some particular value, or the survival function, which is one minus the cumulative distribution function of the coalescence time. Thus our statements about convergence, in Theorem 3.1 and later Theorems 4.1 and 4.5, are about convergence in distribution.
Theorem 3.1 (Unconditional limiting distribution).
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ as $N\to\infty$. Then $N^{-2}\tau^{(N)}$ converges in distribution: for any fixed $t\geq 0$
$\mathbb{P}_{q}{\left(N^{-2}\tau^{(N)}>t\right)}\to\begin{cases}\frac{2-2\alpha}{2-\alpha}e^{-\frac{2}{2-\alpha}t}\qquad&\text{when }\;q=\rm same\\[4.5pt] e^{-\frac{2}{2-\alpha}t}\qquad&\text{when }\;q=\rm diff\end{cases}$
The convergence under $\mathbb{P}_{{\rm diff}}$ of the rescaled coalescence time $N^{-2}\tau^{(N)}$ to an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$ in Theorem 3.1 shows that our Moran model gives the same result as the Wright-Fisher model with partial selfing (Nordborg and Donnelly, 1997; Möhle, 1998) when one averages over the pedigree for a fixed probability of selfing $\alpha\in[0,1]$. Kogan et al. (2023, Lemma 2) recently obtained this result in the broader context of selfing plus two-locus recombination under Wright-Fisher reproduction (and averaging over the pedigree).
The proof of Theorem 3.1 is given in the Appendix, Section 7.1. Two kinds of events are key to our analysis: overlap events where two sample lineages in distinct individuals transition in one time-step to belonging to the same individual, and splitting events where two sample lineages in the same individual transition to belonging to two distinct individuals. We note that these correspond, respectively, to the ‘parent-sharing’ events and ‘separation’ events in the two-locus unconditional model of Uyenoyama (2024). Splitting events occur at outcrossing events, whereas overlap events may be due either to selfing or outcrossing.
Remark 3.2.
For any $k\in\mathbb{Z}_{+}$,
$\mathbb{P}_{\rm diff}(\tau^{(N)}>k)=$ $\,\left(1,0,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}\text{ and }$
$\mathbb{P}_{\rm same}(\tau^{(N)}>k)=$ $\,\left(0,1,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}$
exactly in discrete time. Like its Wright-Fisher counterpart (Nordborg and Donnelly, 1997; Möhle, 1998), Theorem 3.1 is a separation-of-timescales result. Direct application of Möhle (1998, Lemma 1) or its generalization by Möhle and Notohara (2016) is complicated by the extra factor of $N$ in our diploid Moran model, i.e. to go from time in discrete steps to time in units of $N$ generations. Kroumi and Lessard (2015, Proposition 1) have extended Möhle’s lemma to the Moran model. In Section 7.1, we instead use our notions of overlap and splitting events to prove convergence.
4 Main results: conditional coalescence times given the pedigree
Our main results are about the conditional distribution of $\tau^{(N)}$ given the pedigree and the sampled pair of individuals. We consider a sequence of our diploid Moran model indexed by $N$, under three different regimes for the selfing probability $\alpha_{N}$ in the limit. The first corresponds to the model of partial selfing previously studied by others, but holds only when $\alpha_{N}\to 1$ more slowly than $1/N$. The second is what we call limited outcrossing, and happens when $\alpha_{N}\to 1$ at rate $\lambda/N$ for some $\lambda\in(0,\infty)$. The third is when outcrossing is negligible, because $\alpha_{N}\to 1$ faster than $1/N$. As we will show, limited outcrossing interpolates between partial selfing and negligible outcrossing. As $\lambda\to\infty$, it becomes indistinguishable from partial selfing with $\alpha=1$. As $\lambda\to 0$ it becomes indistinguishable from negligible outcrossing.
Following the notation in Diamantidis et al. (2024), we let $\mathcal{A}_{N}$ be the $\sigma$-algebra generated by both the pedigree $\mathcal{G}_{N}$ and the labels $\hat{X}_{0}$ and $\hat{Y}_{0}$ of the sampled individuals. That is, we let
$\mathcal{A}_{N}:=\sigma{\left(\mathcal{G}_{N},\,\hat{X}_{0},\,\hat{Y}_{0}\right)}$ (9)
which by definition contains all the information relevant to the distribution of the coalescence time. Our main results (Theorems 4.1 and 4.5) say that the conditional distributions of $N^{-2}\tau^{(N)}$ given $\mathcal{A}_{N}$ converge to different distributions, depending on the sampling configuration (diff or same) and on the three regimes of the selfing probability. Furthermore, the limiting distributions can retain information of the random pedigree even as $N\to\infty$.
We now state our main results for coalescence times in Theorems 4.1 and 4.5.
Theorem 4.1.
For any $t\in\mathbb{R}_{+}=[0,\infty)$, as $N\to\infty$,
$\mathbb{P}_{{\rm diff}}{\left(N^{-2}\tau^{(N)}>t\,|\,\mathcal{A}_{N}\right)}\to\begin{dcases}e^{-\frac{2}{2-\alpha}t}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\infty\;\text{ and }\,\alpha_{N}\rightarrow\alpha\in[0,1]\\[2.25pt] \mathbb{P}(T_{\lambda}>t\,|\,G_{\lambda})&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\\[2.25pt] \mathbbm{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow 0\end{dcases},$
where $G_{\lambda}$ is the ancestral graph, defined in Definition 4.3, and $T_{\lambda}$ is the coalescence time for a pair of independent simple random walks on $G_{\lambda}$, defined in Definition 4.4.
The proof of Theorem 4.1 is given in the Appendix, Section 7.2.
Remark 4.2.
This result contrasts with the asymptotic behavior of the marginal or unconditional probability $\mathbb{P}_{{\rm diff}}{\left(N^{-2}\tau^{(N)}>t\right)}$ in Theorem 3.1. For example, assuming that $\alpha_{N}=1$ and assuming that $\alpha_{N}\to 1$ at rate $N^{-2}$ lead to the same result in Theorem 3.1 but correspond to negligible outcrossing and partial selfing, respectively, in Theorem 4.1.
Theorem 4.1 offers a description of the conditional distribution of the coalescence time $\tau^{(N)}$ for a sample of two gene copies from different individuals in a population of size $N$ given the pedigree. Simulations of the CDFs for five pedigrees under each of the three cases in Theorem 4.1 are shown in Figure 3. Theorem 4.1 says that the law of $N^{-2}\tau^{(N)}$ for two lineages in different individuals is robust to the structure of the pedigree so long as the rate at which two lineages in the same individual undergo a splitting event, $(1-\alpha_{N})N^{-1}$, is much greater than the rate at which two sample lineages in distinct individuals coalesce, $N^{-2}+O(N^{-3})$, specifically that the ratio of the former over the latter, approximately $N(1-\alpha_{N})$, tends to infinity. This robustness for the case of partial selfing is simply that the conditional limit agrees exactly with the marginal limit in Theorem 3.1.
If, on the other hand, the rate of coalescence is comparable to or dominates the rate of splitting of an individual, as in the cases of limited outcrossing and negligible outcrossing, Theorem 4.1 demonstrates the non-robustness of the “diff” pairwise coalescence time to the pedigree. That is, the survival function of the conditional pairwise coalescence time does not converge to its mean. Indeed, in this case the rates at which lineages split backwards in time is then comparable to the rate at which they coalesce. The subgraph of the pedigree which encodes the possible trajectories of the sample lineages takes the shape of an ancestral graph $G_{\lambda}$ as the population size $N$ goes to infinity, and the ancestries of the two sample lineages correspond to coalescing random walks thereon with coalescence time $T_{\lambda}$. Correspondingly, $N^{-2}\tau^{(N)}$ conditional on the pedigree asymptotically takes the distribution of $T_{\lambda}$ conditional on the ancestral graph.
Figure 3: The limiting conditional CDF $t\mapsto\lim_{N\to\infty}\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N)}\leq t|\mathcal{A}_{N})$ for our Moran model under three assumptions about $\alpha_{N}$: partial selfing, limited outcrossing, and negligible outcrossing. (Left) Partial selfing, where $\lim_{N}N(1-\alpha_{N})=\infty$. The limiting CDF is deterministic and is exactly the CDF of an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$, with $\alpha=1$ for comparison with those at center and right. (Center) Five realizations of the limiting CDF for $\lambda=5$ under limited outcrossing, i.e. with $\lim_{N}N(1-\alpha_{N})=\lambda=5$. The limiting CDF is random and is described precisely in Theorem 4.1. (Right) Five realizations of the limiting CDF for the case of negligible outcrossing, where $\lim_{N}N(1-\alpha_{N})=0$. The limiting CDF is random, and it is the CDF of a random constant that is exponentially distributed with rate 2, i.e. the CDF is a Heaviside function with exponentially distributed jump time.
To complete the exposition of Theorem 4.1, we now formally define the ancestral graph $G_{\lambda}$ and the pairwise coalescence time $T_{\lambda}$ of random walks theoreon.
Definition 4.3 (The ancestral graph $G_{\lambda}$).
Fix $\lambda\in\mathbb{R}_{+}$. By the ancestral graph $G_{\lambda}=\left(G_{\lambda}(t)\right)_{t\in\mathbb{R}_{+}}$ we denote a particle system beginning with two different particles, where each particle splits into two with independently with rate $\lambda$ and each pair of particles coagulate independently at rate $2$.
The ancestral graph is equal in distribution to an ancestral recombination graph (Griffiths, 1991; Griffiths and Marjoram, 1997) or an ancestral selection graph (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). Now consider coalescing random walks $x_{\lambda}$ and $y_{\lambda}$ on $G_{\lambda}$. Note that since $G_{\lambda}$ is the limiting object and $\alpha_{N}\to 1$, these walks coalesce with probability one the first time they meet. Let $G_{\lambda}(0)=\{x(0),y(0)\}$ consist of the two different particles at $t=0$. Furthermore, let $\left(x_{\lambda}(t),y_{\lambda}(t)\right)_{t\in\mathbb{R}_{+}}$ be a $G_{\lambda}$-valued continuous-time process starting at $(x(0),y(0))$ such that
  • (i)
    $\{x_{\lambda}(t),y_{\lambda}(t)\}\subset G_{\lambda}(t)$ for all $t\in\mathbb{R}_{+}$,
  • (ii)
    at any splitting event, each of $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ will follow each of the two paths available with equal (i.e. 1/2) probability.
Then, $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ are coalescing random walks on the ancestral graph $G_{\lambda}$ starting at two different particles.
Definition 4.4 (The coalescence time $T_{\lambda}$).
The first meeting time of $x_{\lambda}$ and $y_{\lambda}$ is
$T_{\lambda}:=\inf\{t\in\mathbb{R}_{+}:\,x_{\lambda}(t)=y_{\lambda}(t)\}.$ (10)
How one may reconstruct the CDF of the $T_{\lambda}$, conditional on a realization of $G_{\lambda},$ is discussed in Figure 4. Two realizations of $T_{\lambda}$ on the same ancestral graph are demonstrated in Figure 6.
Figure 4: (Left) A realization of the random ancestral graph $G_{\lambda}$ starting with two particles/nodes. For this particular realization, the overlap times of the sample lineages are $t_{1}<t_{2}<t_{3}$. (Right) The CDF corresponding to this $G_{\lambda}$. The conditional distribution of the limiting coalescence time $T_{\lambda}$, given $G_{\lambda}$, is obtained by tracing ancestral genetic lineages backwards in time along the graph $G_{\lambda}$. Given this $G_{\lambda}$, $T_{\lambda}$ must take values in $\{t_{1},t_{2},t_{3}\}$. One can read off the conditional CDF of $T_{\lambda}$ as follows, by tracing the ancestries of a hypothetically infinite number of unlinked loci. The node of $G_{\lambda}$ at time $t_{1}$ contains $3/4$ of the ancestries from the right sample lineage and the whole of those from the left sample lineage, so $\mathbb{P}(T_{\lambda}=t_{1}|G_{\lambda})=3/4$. Between $t_{1}$ and $t_{2}$, half of the loci following the left ancestral lineage go left at the split and the other $1/2$ continue up. The lineage on the right between $t_{1}$ and $t_{2}$ contains $1/4$ of the ancestries from the right sample lineage. So when these two lineages overlap at $t_{2}$ we have $\mathbb{P}(T_{\lambda}=t_{2}|G_{\lambda})=1/8$. Finally the remainder of the right sample lineage meets the remainder of the left sample lineage at $t_{3}$ for $\mathbb{P}(T_{\lambda}=t_{3})=1/8$.
Analogously to Theorem 4.1, we obtain the limiting conditional distribution of the coalescence time when our samples are taken from the same individual.
Theorem 4.5.
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ as $N\to\infty$. For all $t\in\mathbb{R}_{+}$ we have convergence in distribution
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N)}>t|\mathcal{A}_{N})\to\begin{dcases}e^{-t}&\quad\text{if }\alpha=0\\ 2^{-U}e^{-\frac{2}{2-\alpha}t}&\quad\text{if }\alpha\in(0,1)\\[2.25pt] 0&\quad\text{if }\alpha=1\end{dcases},$
where $U$ is a random variable satisfying $\mathbb{P}(U=k)=\alpha^{k}(1-\alpha)$ for $k\in\mathbb{Z}_{+}$.
The proof of Theorem 4.5 is given in the Appendix, Section 7.2.
As in (2), the quantity $U$ represents the number of selfing events in the ancestry of the sampled individual, before the first splitting or outcrossing event. To make the connection with Theorem 4.1, the first two cases in Theorem 4.5 pertain to partial selfing (or case one) in Theorem 4.1, where is the possible for a “same” sample to enter the “diff” process. The third case in Theorem 4.5 pertains to both limited outcrossing and negligible outcrossing (or cases two and three) in Theorem 4.1, where the two gene copies in a “same” sample coalesce instantaneously with probability $1$.
4.1 Extension to larger samples
We anticipate that our main results can readily be extended to samples of arbitrary size $n>2$. We offer a detailed conjecture here but leave further exploration and proof to future work.
Suppose we have a sample of $n$ genes from the population at time-step 00, where $n>2$ is an arbitrary fixed number. As reviewed in Berestycki (2009), it is customary in formal treatments of coalescent theory to summarize the ancestral relationships of the sample by an ancestral process which we can call $(\Pi_{k}^{N,n})_{k\in\mathbb{Z}_{+}}$ and which takes values in the space $\mathcal{E}_{n}$ of partitions of $\{1,\ldots,n\}$, where indices belong to the same block of $\Pi^{N,n}_{k}$ if and only if the samples with these indices have a common ancestor $k$ steps in the past. Under $\mathbb{P}_{{\rm diff}}$, which now pertains to $n$ samples from $n$ distinct individuals, the initial state $\Pi_{0}^{N,n}$ of the ancestral process has $n$ singleton blocks. Furthermore, let $\mathcal{A}_{N}$ now be the $\sigma$-algebra generated by both the pedigree and the labels of the $n$ individuals from whom we sample the $n$ sample lineages. We make the following conjecture about diff samples.
Conjecture 4.6.
Under the conditional law $\mathbb{P}_{{\rm diff}}(\cdot|\mathcal{A}_{N})$, the sequence of $\mathcal{E}_{n}$-valued processes $\Pi^{N,n}=(\Pi^{N,n}_{\lfloor tN^{2}\rfloor})_{t\in\mathbb{R}_{+}}$ converges in finite dimensional distributions, as $N\to\infty$, to the random process
$\Pi=\begin{cases}\mathcal{K}^{(\alpha)}_{n}&\quad\text{if }N(1-\alpha_{N})\rightarrow\infty\quad\text{and }\alpha_{N}\rightarrow\alpha\\ P_{\lambda}\quad\text{under }\mathbb{P}(\cdot|G_{\lambda}^{n})&\quad\text{if }N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\end{cases},$ (11)
where $\mathcal{K}^{(\alpha)}_{n}$ is the the Kingman $n$-coalescent (Kingman, 1982) with time rescaled by $\frac{2}{2-\alpha}$, $G_{\lambda}^{n}$ is an ancestral graph starting with $n$ nodes, and $P_{\lambda}=(P_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ is a $\mathcal{E}_{n}$-valued process on $G_{\lambda}^{n}$ which is the extension of (10) to larger samples, so that $i\sim j$ (coalesce) if and only if $x_{i,\lambda}(t)=x_{j,\lambda}(t)$.
Furthermore we expect that as $\lambda\rightarrow\infty$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to the Kingman $n$-coalescent $\mathcal{K}^{(\alpha)}_{n}$ as implied by (19) below in Section 5, and that as $\lambda\rightarrow 0$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to a single realization of $\mathcal{K}^{(1)}_{n}$, which is the limit in the negligible-outcrossing regime.
For the first case in (11) we further conjecture, for a sample of size $n$ containing $2m$ gene copies together as pairs in $m$ individuals and $n-2m$ gene copies each in different individuals, that each of the $m$ individuals will have its own number of selfing generations before its first outcrossing event and that these will be given by $m$ independent draws of the random variable $U$ in (2). This is a straightforward extension of the coalescent with partial selfing in Nordborg and Donnelly (1997).
5 Additional properties of limited outcrossing
Here we provide further analyses and illustrations of the novel scaling limits in Theorems 4.1 and 4.5, in particular the coalescing random walk on the ancestral graph. We note that the simulations for the Wright-Fisher model in Figure 1 show similar behavior as those for our Moran model in Figure 3.
Observe first that the limited-outcrossing regime in Theorem 4.1 includes the case $\lambda=0$, which means that the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ when $\lambda=0$ is equal to that expected under negligible outcrossing, namely a single draw from an exponential distribution with rate 2. The other extreme, $\lambda\to\infty$, is covered by (19) below. Namely, as $\lambda\to\infty$, the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ converges to ${\rm Exp}(2)$, which corresponds to the partial-selfing regime with $\alpha=1$. Hence, for lineages in different individuals, the limited-outcrossing regime covers all possible cases for which $\alpha_{N}\to 1$, as we tune the magnitude of $\lambda$. This highlights the significance of the limited-outcrossing regime.
To understand how information relevant to conditional gene genealogies can persist even in the limit $N\to\infty$, we introduce a discrete-time ancestral graph, $G^{N}=\left(G^{N}(k)\right)_{k\in\mathbb{Z}_{+}}$, which is a subset of the pedigree containing all potential individuals through which our sample of $n=2$ gene copies may have been transmitted. Each realization of $G^{N}$ is constructed by the following two rules. We refer to potential ancestral individuals of the sample as nodes. When a node is the result of an outcrossing event, it fragments into two nodes. When one node is the offspring of another node in a selfing event, the two coagulate into a single node. Individual genetic lineages are transmitted through this graph according to Mendel’s laws as in Section 2. A realization of $G^{N}$ is shown in Figure 5.
Figure 5: A realization of $G^{N}$ containing potential ancestors of the two sampled individuals (ovals) at the bottom. In past time-step $3$, the ancestral individual on the left undergoes an outcrossing event. In past time-step $4$, one of the nodes from this event is the offspring of one of the potential ancestral individuals of the sampled individual on the right, and these two nodes coagulate. In the fifth time-step, the right most particle experiences an outcrossing in which exactly one of the parents is a node in the graph. This may appear in the discrete-time ancestral graph but will not occur in the limiting ancestral graph.
This graphical structure will be poorly behaved as $N\to\infty$ if the number of nodes diverges as the population size goes to infinity. For the random variable $G^{N}$ to have a suitable limit as the population size goes to infinity we require that the rate at which nodes fragment, approximately $(1-\alpha_{N})N^{-1}$, and the rate at which pairs of nodes coagulate, of order $N^{-2}$, are comparable. This comparability is taken to mean that the ratio of these two rates, $N(1-\alpha_{N})$, converges as $N$ goes to infinity to $\lambda\in\mathbb{R}_{+}$. This reasoning gives rise to the following lemma.
Lemma 5.1.
Suppose $N(1-\alpha_{N})\to\lambda\in\mathbb{R}_{+}$. The discrete-time ancestral graph under the coalescence timescale $\left(G^{N}(\lfloor tN^{2}\rfloor)\right)_{t\in\mathbb{R}_{+}}$ converges in distribution to the ancestral graph $G_{\lambda}$ as $N\to\infty$.
The proof of Lemma 5.1 is given in the Appendix, Section 7.2.2. This requires additional details which we have omitted here for simplicity. The significance of Lemma 5.1 is that the simple subgraph $G^{N}$ captures all the relevant information in the pedigree for the limiting ancestral graph $G_{\lambda}$. One difference between these graphs is that ancestral genetic lineages do not coalesce immediately at each overlap event in $G^{N}$, whereas they do necessarily at $T_{\lambda}$ in the limiting graph $G_{\lambda}$.
Next, we consider the variances of the conditional survival probability $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ and of the conditional expectation $\mathbb{E}\left[T_{\lambda}|G_{\lambda}\right]$. We obtain explicit formulas and their asymptotics as $\lambda\to 0$ and $\lambda\to\infty$. The starting point of our analysis is a pair of conditionally independent gene genealogies given the graph for the sampled individuals, cf.  Diamantidis et al. (2024) and Birkner et al. (2013). Figure 6 depicts realizations of coalescence times for two such loci.
Lemma 5.2.
For any $\lambda\in(0,\infty)$ and $t\in(0,\infty)$,
$\mathrm{Var}{\left(\mathbb{P}(T_{\lambda}>t|G_{\lambda})\right)}=\mathbb{P}(T_{\lambda}\wedge T_{\lambda}^{\prime}>t)-e^{-4t}\quad\text{ and }\quad\mathrm{Var}{\left(\mathbb{E}\left[T_{\lambda}|G_{\lambda}\right]\right)}=\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)},$ (12)
where $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are two conditionally independent copies of the pairwise hitting time given the same $G_{\lambda}$, as defined in (10) with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$ and $x_{\lambda}(0)\neq y_{\lambda}(0)$.
Proof.
First, note that $\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t|G_{\lambda})$ has mean $\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t)=e^{-2t}$ which is the same for all $\lambda$. The last equality holds because every pair of particles in the ancestral graph $G_{\lambda}$ coagulates at rate $2$. Therefore, the first equality in (12) follows from the following equality:
$\mathbb{E}\left[\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t|G_{\lambda})^{2}\right]=$ $\,\mathbb{E}\left[\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t|G_{\lambda})\;\mathbb{P}_{x_{\lambda}^{\prime}(0)\neq y_{\lambda}^{\prime}(0)}(T^{\prime}_{\lambda}>t|G_{\lambda})\right]$
$=$ $\,\mathbb{E}\left[\mathbb{P}_{x_{\lambda}^{\prime}(0)=x_{\lambda}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)}(T_{\lambda}>t,\;T^{\prime}_{\lambda}>t|G_{\lambda})\right]$
$=$ $\,\mathbb{P}_{x_{\lambda}^{\prime}(0)=x_{\lambda}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)}(T_{\lambda}\wedge T^{\prime}_{\lambda}>t).$
The second equality in (12) follows from the conditional independence of $T_{\lambda}$ and $T_{\lambda}^{\prime}$. ∎
Figure 6: Two conditionally independent realizations of the pairwise process, $(x_{\lambda},y_{\lambda})$ on the same ancestral graph $G_{\lambda}$. The hitting times of these random walks are marked by a purple dot. These times correspond to $T_{\lambda}$ and $T_{\lambda}^{\prime}$, the coalescence times of two unlinked loci given the same pedigree. In each realization, $x_{\lambda}$ is in blue and $y_{\lambda}$ in red. At each splitting event in the graph, where a particle splits into two, each walk chooses which particle path to follow fairly, i.e. with probability $1/2$ each. With respect to $\{t_{1},t_{2},t_{3}\}$ shown in Figure 4, here $T_{\lambda}=t_{1}$ on the left and $T_{\lambda}^{\prime}=t_{2}$ on the right.
The random variables $T_{\lambda}$ and $T_{\lambda}^{\prime}$ in Lemma 5.2 are the pairwise meeting times for two conditionally independent processes $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ with the same starting point, i.e. with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$. They correspond to the ancestral processes for $n=2$ at two unlinked loci starting from the same two sampled individuals, as in Figure 6. To model the full two-pairs process, we consider all possible possible ancestral states of the samples. For simplicity, we combine states where possible by symmetry and restrict ourselves to a process with the following five states.
$s_{0}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{1}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{2}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{\Delta,1}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
$s_{\Delta,2}$ $=\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\end{pmatrix}\begin{pmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}\,\textrm{or}\,\begin{pmatrix}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bullet}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\end{pmatrix}$
Here, two dots of the same color (blue or red) correspond to two particles of the same pair; say, the two blue dots correspond to $(x_{\lambda},y_{\lambda})$ and the two red dots correspond to $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$. Two dots in parentheses correspond to taking the same value in $G_{\lambda}$, i.e. being in the same individual. Hence, for $i\in\{0,1,2\}$, the state $s_{i}$ corresponds to when $i$ individuals or particles in $G_{\lambda}$ contain two conditionally independent lineages.
[…]
Lemma 5.3.
Let $S=(S_{t})_{t\in\mathbb{R}_{+}}$ be the process, with state space $\{s_{0},\,s_{1},\,s_{2},\,s_{\Delta,1},\,s_{\Delta,2}\}$, which tracks the two pairs $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ at time $t$ (backward) on the coalescence time-scale, then its rate matrix $Q_{\lambda}$ is
$Q_{\lambda}=\begin{pmatrix}-12&8&0&4&0\\ \frac{\lambda}{2}&-6-\frac{\lambda}{2}&2&4&0\\ 0&\lambda&-2-\lambda&0&2\\ 0&0&0&-2&2\\ 0&0&0&0&0\end{pmatrix}.$ (13)
[…]
Then, the covariances of $T_{\lambda}$ and $T_{\lambda}^{\prime}$ for the different sample configurations are exactly the same as the classical formulas in two-locus models with recombination, given for example by equations (7.30), (7.29) and (7.28) in Wakeley (2009). Thus we have
$\mathrm{Cov}_{s_{0}}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ $=\frac{4}{\lambda^{2}+26\lambda+72},$ (14)
$\mathrm{Cov}_{s_{1}}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ $=\frac{6}{\lambda^{2}+26\lambda+72},$ (15)
$\mathrm{Cov}_{s_{2}}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ $=\frac{1}{2}\frac{\lambda+36}{\lambda^{2}+26\lambda+72}.$ (16)
Proposition 5.4.
For any $\lambda\in\mathbb{R}_{+}$, the variance of $\mathbb{E}{\left[T_{\lambda}|G_{\lambda}\right]}$ is equal to
$\mathrm{Var}{\left(\mathbb{E}{\left[T_{\lambda}|G_{\lambda}\right]}\right)}=\mathrm{Cov}{\left(T_{\lambda},T^{\prime}_{\lambda}\right)}=\frac{1}{2}\frac{\lambda+36}{\lambda^{2}+26\lambda+72}.$ (17)
In particular,
$\mathrm{Var}{\left(\mathbb{E}{\left[T_{\lambda}|G_{\lambda}\right]}\right)}=\begin{dcases}\frac{1}{4}-\frac{1}{12}\lambda+O(\lambda^{2})&\text{ as }\lambda\to 0\\[4.0pt] \frac{1}{2\lambda}+\frac{5}{\lambda^{2}}+O(\lambda^{-3})&\text{ as }\lambda\to\infty\end{dcases}.$
Proof.
The first equality in (17) follows directly from (12), the second from (16). ∎
Next, we consider asymptotics of the variance of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$.
Proposition 5.5.
For any $t\in(0,\infty)$,
$\mathrm{Var}\left(\mathbb{P}(T_{\lambda}>t|G_{\lambda})\right)=\begin{dcases}e^{-2t}-e^{-4t}+\lambda\frac{e^{-2t}}{8}\left(1-4t-e^{-4t}\right)+O(\lambda^{2})&\quad\text{ as }\lambda\to 0\\ \frac{2e^{-4t}}{\lambda}\left(1+\frac{2(3+16t)}{\lambda}\right)+O(\lambda^{-3})&\quad\text{ as }\lambda\to\infty\end{dcases}.$ (18)
This follows from the computation of the second moment of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ in Lemma 7.23 in the Appendix. Then we have, for any $t\in(0,\infty)$,
$\mathbb{P}(T_{\lambda}>t|G_{\lambda})\to\begin{dcases}\textbf{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{ as }\lambda\to 0\\ e^{-2t}&\quad\text{ as }\lambda\to\infty\end{dcases}.$ (19)
The proofs of both Lemma 7.23 and Proposition 5.5 are given in the Appendix, Section 7.3.
We observe that the covariance $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is connected to the difference between the conditional survival function $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ and the unconditional survival function $\mathbb{P}(T_{\lambda}>t)=e^{-2t}$.
Proposition 5.6.
For any $\lambda\in(0,\infty)$,
[…]
Proof.
By Lemma 5.2, the left-hand side is equal to
[…]
where by $T_{\lambda}\wedge T_{\lambda}^{\prime}$ we mean the smaller of $T_{\lambda}$ and $T_{\lambda}^{\prime}$. ∎
Remark 5.7.
In agreement with previous work—see (3) and (52) in Simonsen and Churchill (1997)—from (16) and the fact $\mathrm{Var}{\left(T_{\lambda}\right)}=1/4$, we see that $\mathrm{Corr}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is exactly the same as $\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$. Equivalently, $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}=\frac{1}{4}\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$, and it is also equal to the left hand side of (20). The covariance formula obtained here is consistent with Wilton et al. (2015, Eq. 9) by taking $2T_{\lambda}=T_{1}$ and $2T_{\lambda}^{\prime}=T_{2}$; the rescaling is due to the fact that the marginals on $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are exponentially distributed with rate $2$, while those of $T_{1}$ and $T_{2}$ are with rate $1$.
Depending on $\lambda\in(0,\infty)$, limited outcrossing displays a range of behaviors between the two qualitatively different extremes of negligible outcrossing and partial selfing. Figure 7 illustrates this using the law of total variance for the coalescence time of two gene copies sampled from two different individuals. Let $T$ and $G$ represent this “diff” coalescence time and its associated ancestral graph, for simplicity dropping the subscript $\lambda$ we attached to these previously. Then we can write
$\textrm{Var}{\left(T\right)}=\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}+\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}.$ (21)
Note that $\textrm{Var}{\left(T\right)}=1/4$, because $T$ by itself is exponentially distributed with rate parameter $2/(2-1)=2$. The second term on the right-hand side of (21), $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, is given by (17). The first term on the right-hand side of (21) is just $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}=1/4-\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, but both terms are displayed in Figure 7 for the sake of illustration.
Figure 7: The two components, $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$ and $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, of $\textrm{Var}{\left(T\right)}$ as functions of $\lambda$ under limited outcrossing. By the law of total variance, $\textrm{Var}{\left(T\right)}=\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}+\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$. Also, $\textrm{Var}{\left(T\right)}=1/4$ in this case because it is the average over ancestral graphs ($G$), that is to say over pedigrees, and because $\alpha_{N}\to 1$.
When $\lambda$ is very small, on the far left in Figure 7, variation in the mean coalescence time among graphs, $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, accounts for the bulk of the total variation in $T$. Correspondingly, there is little variation in coalescence times given the graph. Figure 7 displays the expectation of the latter, $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$, but $\textrm{Var}{\left(T|G\right)}$ itself will be small on any graph in this case. With probability approaching one as $\lambda\to 0$, the individuals containing the two ancestral lineages will experience only selfing events in their ancestries, up to and including the time they descend from a common ancestral individual. Pairwise coalescence times at every locus in the genome starting from these two individuals will converge on the time of this ancestor and $\textrm{Var}{\left(T|G\right)}$ should be minimal. Variation in the waiting time to this first common ancestral individual among graphs will be the primary source of variation in $T$. In the limit $\lambda\to 0$, this waiting time is exponentially distributed with rate parameter $2$, so $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ approaches $\textrm{Var}{\left(T\right)}=1/4$. It is given by a Kingman coalescent process made effectively haploid by selfing, just as under negligible outcrossing.
Across the middle of Figure 7, ancestral graphs will include increasing numbers of splitting or outcrossing events as $\lambda$ increases. Such graphs contain multiple pathways to coalescence because genetic lineages may trace back to either parent at each outcrossing event. They include multiple possible coalescence times, as already evident in the CDFs of $T|G$ in Figure 4 for a hypothetical graph with three splitting events and three possible coalescence times, in Figure 3 for each of five simulated ancestral graphs with $\lambda=2$, and in Figure 1 for each of fifty pedigrees in three versions of a finite-$N$ Wright-Fisher model corresponding to $\lambda\in\{1,10,100\}$. As a result, $\textrm{Var}{\left(T|G\right)}$ is positive and $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$ becomes the main source of variation in $T$ as $\lambda$ increases, eventually swamping variation of the mean coalescence times among graphs.
When $\lambda$ is very large, on the far right in Figure 7, single ancestral graphs do little to constrain $T|G$. These graphs are large. In their construction backward in time, the number of ancestral lineages grows quickly to reach a quasi-stable distribution, specifically a zero-truncated Poisson distribution with mean $\lambda$, as Mano (2009) showed for the ancestral recombination graph. The waiting time to the event that completes the graph, when there is just one lineage left, will be long, $\sim 2e^{\lambda}/\lambda^{2}$ in expectation using equation (1.1) in Griffiths and Marjoram (1997). Given the graph, each of the pair of genetic lineages starting from the two sampled individuals will trace back through a large number of splitting events before the two meet in a single individual and coalesce. As $\lambda\to\infty$, all variation in $T$ will be due to this random process of coalescence given the graph. Equation (19) shows that this conditional coalescent process converges on that of the partial-selfing model with $\alpha_{N}\to 1$. Proposition 5.6 shows further that $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ in (17) is a key measure of the distance of the CDF under limited outcrossing to that under the corresponding model of partial selfing.
To summarize what is shown in Figure 7, limited outcrossing becomes like negligible outcrossing as $\lambda\to 0$ and like partial selfing as $\lambda\to\infty$. With respect to the pre-limiting Moran model, the results and intuitions about Figure 7 concern the behavior at the boundary of complete or nearly complete selfing ($\alpha_{N}\to 1$). All at this boundary, we find: first negligible outcrossing, then limited outcrossing, then an extreme case of partial selfing with $\alpha=1$. The entire rest of the range of selfing probabilities $\alpha\in[0,1]$ is covered by the partial-selfing model.
Finally, although a treatment of recombination is beyond the scope of this work, it is relatively straightforward to describe the joint ancestral process of a sample of size two at two linked loci under limited outcrossing. That is, we can extend the analysis giving the rate matrix (13) and the covariances (14)-(16) to the case of two linked loci.
[…]
An equilibrium is established by the balance of two events in a purely selfing individual lineage: gene copies on different chromosomes in an offspring will be on the same chromosome in the parent with probability $1/2$, and gene copies on the same chromosome in an offspring will be on different chromosomes in the parent with probability $r$. The equilibrium probability that they are on different chromosomes is
$p_{r}\coloneq\frac{2r}{1+2r}$ (22)
and it is only in this case that the two will separate, one going to each parent, when a splitting event occurs. Thus, the effective rate of splitting is $\lambda p_{r}$. Note that $r=1/2$ for unlinked loci and this gives $p_{r}=1/2$, as already noted regarding (13).
In sum, the two locus ancestral graph process with recombination can be modeled with the same states as above and with rate matrix (13) but with $\lambda p_{r}$ in place of $\lambda/2$. Similarly, the covariances of coalescence times at two linked loci are given by (14)-(16) after making this same substitution. We emphasize, cf. Lemma 5.2 and Figure 7, that these covariances are averages over the ancestral graph or the pedigree. Kogan et al. (2023) recently derived such covariances for arbitrary $N$, $r$ and $\alpha_{N}$ in a model of Wright-Fisher reproduction. Their detailed exact expressions collapse to ours when the same limit is taken with $N(1-\alpha_{N})\to\lambda$ and accounting for the difference in time scale between the Wright-Fisher and Moran models (results not shown).
6 Discussion
In this work, we have assumed that the pedigree of a population of constant size $N$ is the outcome of a random process of reproduction in which offspring are produced by self-fertilization with probability $\alpha_{N}$. By conditioning on the pedigree and considering pairwise times to common ancestry, we found three different coalescent models in the limit $N\to\infty$, when time is measured in units of $N$ generations (which is $N^{2}$ time-steps) in the Moran model. The deciding factor is $N(1-\alpha_{N})$, which may be interpreted either as the rate of outcrossing events along an ancestral lineage or as the expected number of outcrossed offspring in the population each generation. The three models are ‘negligible outcrossing’ which applies when $N(1-\alpha_{N})\to 0$, ‘limited outcrossing’ which applies when $N(1-\alpha_{N})\to\lambda\in(0,\infty)$, and ‘partial selfing’ which applies when $N(1-\alpha_{N})\to\infty$. Negligible outcrossing and limited outcrossing both require $\alpha_{N}\to 1$, while partial selfing includes this as a special case.
Previous population-genetic analyses of selfing populations have assumed a fixed selfing probability, $\alpha_{N}=\alpha$ in our notation, and have implicitly averaged over pedigrees. When this averaging is done, just one coalescent model for all $\alpha\in[0,1]$ emerges in the limit, specifically the coalescent model with partial selfing described by Nordborg and Donnelly (1997) and Möhle (1998). When coalescence is conditioned on the pedigree and $\alpha$ is fixed, our very similar partial-selfing model is obtained but only for $\alpha\in[0,1)$ when $\alpha$ is fixed. Assuming a fixed selfing probability of $\alpha=1$ leads to an entirely different model, namely negligible outcrossing.
A more detailed look at how $\alpha_{N}$ approaches $1$ reveals the critical case of limited outcrossing in between negligible outcrossing and partial selfing, and characterized by the parameter $N(1-\alpha_{N})\to\lambda\in(0,\infty)$. In this regime, the parts of the population pedigree directly ancestral to the sampled individuals can be replaced by a random graph like the ones previously used to model recombination (Griffiths, 1991; Griffiths and Marjoram, 1997) and selection (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). It is distinguished from these only by what the splitting events in the graph represent and how these are treated in modeling gene genealogies. Each splitting event in the graph corresponds to an outcrossing event in the genealogical ancestry of the sampled individuals. Ancestral lineages tracing back to such a node in the graph make a $50$:$50$ choice of which branch they follow. The rate at which ancestral lineages split is precisely our parameter $\lambda$.
As with the well known Wright-Fisher diffusion and the corresponding standard neutral coalescent model, these new models of coalescence conditional on the pedigree are meant as robust approximations for large populations. We expect, for example, that the Wright-Fisher model with selfing used to produce Figure 1 will have the same three limiting cases we found for the Moran model with selfing. Which of the three is most applicable for a given (necessarily finite) population will depend on $N$ and $\alpha_{N}$. The special case of negligible outcrossing involves the strong predictions of identical gene genealogies across all loci in the genome and zero heterozygosity, and so may be of limited utility. The critical case of limited outcrossing allows gene genealogies to differ across the genome but also predicts zero heterozygosity. Still it is clear from the jumps in probability in Figure 1 that if $N(1-\alpha_{N})$ is not large, insights from the limited-outcrossing model are helpful even when the value of $\alpha_{N}$ would not prevent heterozygosity. On the other hand, both Figure 1 and Figure 7 suggest that if $N(1-\alpha_{N})$ is greater than about $100$, partial-selfing is a better description than limited-outcrossing, even if the outcrossing probability $1-\alpha_{N}$ is quite small.
We chose to use the name “partial selfing” here because in this case the coalescent process conditional on the pedigree is very similar to the previously described partial-selfing model which did not condition on the pedigree (Nordborg and Donnelly, 1997; Möhle, 1998). Our model may be considered an update of the previous one. In both, times to common ancestry for “diff” samples follow the Kingman coalescent process with effective population size $N_{e}=(2-s)N/2$. In our formulation, partial selfing is obtained when $N(1-\alpha_{N})\to\infty$. Similarly we may recall from Figure 7 that as $\lambda\to\infty$ the predictions for “diff” samples under limited outcrossing converge on those for partial selfing. The reason that pedigrees do not constrain “diff” coalescence times under the partial-selfing model is that outcrossing events dominate the ancestry of the population. The same underlying phenomena are at work here as in the standard models of randomly mating populations, where the fact that individuals have two parents with high probability strongly influences the pedigree (Chang, 1999; Derrida et al., 1999; Coron and Le Jan, 2022) resulting in a fast mixing time of ancestral genetic processes given the pedigree (Barton and Etheridge, 2011). Because of this, the conditional and unconditional coalescent processes for these standard population models converge to the same Kingman coalescent process (Tyukin, 2015; Diamantidis et al., 2024).
The key update in our partial-selfing model is the inclusion of the number of generations of selfing up to the first outcrossing event in the recent ancestry of sampled individuals. In our model, a sampled individual has a single realization of the random variable $U$ in Theorem 4.5 or in (2). In the previous model, an average is taken over this distribution then this average, specifically the inbreeding coefficent $F=s/(2-s)$, is applied to every individual. Conditioning on the pedigree produces a coalescent model which aligns better with other work on partial selfing in population and evolutionary genetics. This is particularly true for multi-locus data, which are an explicit concern of conditional coalescent processes because all loci are transmitted through the same pedigree.
Let us take for granted the conjecture we made in Section 4.1 concerning a sample of size $n$ with $2m$ gene copies together as pairs in $m$ individuals. At a locus $i$ in such a sample, there will be a random number of effectively instantaneous coalescent events $X_{i}$, then the remaining $n-X_{i}$ lineages will enter the Kingman coalescent process. If $\{k_{1},k_{2},\ldots,k_{m}\}$ are the realized numbers of selfing generations in the ancestries of the $m$ individuals, then $X_{i}\sim\sum_{r=1}^{m}\textrm{Bernoulli}\left({1-2^{-k_{r}}}\right)$. An unlinked locus, $j\neq i$, has the same $\{k_{1},k_{2},\ldots,k_{m}\}$ but a conditionally independent value of $X_{j}$ and an independent realization of the Kingman coalescent process.
This coalescent model with partial selfing conditional on the pedigree accurately predicts the variation in the level of inbreeding among individuals which Weir and Cockerham (1973) developed into a theory of identity disequilibrium; recall (1) and (4). Like all coalescent models, the focus is now on the sample rather than the population. If we write $Y_{i,r}$ for the indicator random variable of coalescence in the recent ancestry of sampled individual $r\in\{1,2,\ldots,m\}$, then $E[Y_{i,r}]=s/(2-s)$ as in (3) and $\textrm{Var}[Y_{i,r}]$ is given by (1) or (4). If we were able to observe the outcomes $y_{i,r}$ for $r\in\{1,2,\ldots,m\}$, then the mean among loci $\bar{y}_{r}$ would be an estimate of $1-2^{-k_{r}}$ and the sample variance of $\bar{y}_{r}$ among individuals would be an estimate of the identity disequilibrium (1).
Identity by descent cannot be observed directly but only inferred based on patterns of identity in state (Pollak, 1987; Rousset, 2002; Uyenoyama, 2024). A number of methods have been developed and applied to estimate inbreeding coefficients or selfing probabilities. Some have made use of $U$ in (2) or similar distributions to estimate selfing probabilities of populations (Enjalbert and David, 2000; David et al., 2007; McClure and Whitlock, 2012). Others have employed identity disequilibria (Ritland and Jain, 1981; Ritland, 1996; Sweigart et al., 1999; Yang, 2002; Jarne and David, 2008; Szulkin et al., 2010; Vieira et al., 2013) or runs of homozygosity (Franklin, 1977; Yi et al., 2022; Zeitler and Gilbert, 2024) to estimate individual inbreeding coefficients due to selfing. Others have taken up the problem of estimating the number of generations back to the most recent outcrossing event for each individual together with the selfing probability (Gao et al., 2007; Wang et al., 2012; Redelings et al., 2015). Our results for partial selfing establish a coalescent framework for these methods, and provide for new coalescent-based estimators like the one described in Nordborg and Donnelly (1997, Eq. 20) but incorporating (2) and $\{k_{1},k_{2},\ldots,k_{m}\}$ above in place of just $F=s/(2-s)$.
Of these empirical works, the ones which appeal to the previous, unconditional coalescent model with partial selfing for justification do so with respect to the properties of “diff” samples, i.e. that they follow a Kingman coalescent with diploid effective size $N_{e}=(2-s)N/2$. Again, this is also a property of the conditional model we have described. But the coalescent model with partial selfing conditional on the pedigree further includes rigorous justification for the modeling of individual ancestries of selfing already done in these works. For example, in the inference method for infinite-alleles data developed by Redelings et al. (2015), individual ancestries are precisely those of the conditional model. Then, cf. Figure 1 in Redelings et al. (2015), the allelic states of the distinct ancestral lineages which emerge from these random outcomes at a locus, in our notation the $n-X_{i}$ lineages described above, are given by the Ewens sampling formula (Ewens, 1972) with the mutation parameter for a Kingman coalescent with diploid effective size $N_{e}=(2-s)N/2$.
The existence of the limited-outcrossing coalescent model, with its simple structure of the random graph, can aid in the interpretation of multi-locus genetic data even if its extreme prediction of zero heterozygosity of individuals is not met. Each branching event in the ancestral graph corresponds to an outcrossing event between ancestral individuals in the pedigree of the sample, in the case that these are rare, i.e. with probability proportional to $1/N$ in the model. Forward in time and to the extent that the two ancestral individuals differ genetically, each such event would create a potentially novel recombinant inbred line. The partial-selfing model approximates this when the selfing probability is high, except that some of these new lines may not become fully inbred.
Samples from predominately selfing populations are often dominated by individuals homozygous for one of only a few distinct mutli-locus genotypes, together with some recombinant inbred or partially inbred individuals (Bonnin et al., 2001; Bakker et al., 2006; Bomblies et al., 2010; Hartfield et al., 2017; Jullien et al., 2019). The Medicago truncatula data for $24$ unlinked loci in a sample of $200$ individuals presented and analyzed by Siol et al. (2008) illustrates this. They found four common multi-locus genotypes present in $76$, $34$, $25$ and $17$ individuals in the sample. A further $21$ recombinant inbred genotypes, resulting from relatively ancient crosses between the common types, were distributed among $39$ individuals in the sample. At lowest frequency in the sample were $9$ recombinant partially inbred (heterozygous) genotypes present in single individuals, presumably with relatively recent outcrossing events in their ancestries.
We have analyzed a simple population model in some detail to understand how the population pedigree shapes the coalescent process for all loci in a genome. The model includes just reproduction and selfing. There is no selection, no mutation, no recombination, no population structure, and no change in the population size over time. We expect that neutral mutation and changes in population size could be included the same way they are included in coalescent models which do not condition on the pedigree. So, for example, applications of coalescent theory to “diff” samples under partial selfing which infer the ancestry of populations, such as those in St. Onge et al. (2011), Brandvain et al. (2013), Beissinger et al. (2016), are as well justified under our conditional partial-selfing model as they are under the unconditional model of Nordborg and Donnelly (1997).
It would not be as straightforward to incorporate selection or population structure because these can have their own effects on the pedigree. Some simulations demonstrating this can be found in Kuo and Avise (2008), Wakeley et al. (2016) and Wilton et al. (2017). It is an open question how recombination and coalescence operate conditional on the population pedigree. We have dealt mostly with free recombination here. In particular, Theorem 4.1 and Theorem 4.5 are about the distribution of conditionally independent coalescence times given the pedigree. In the case of limited outcrossing, we were able to include an analysis of two loci with recombination which gives covariances of coalescence times consistent with the discrete, exact results for the unconditional or pedigree-averaged case in Kogan et al. (2023).
Finally, selfing is just one form of inbreeding. The importance of modeling and disentangling the sources of inbreeding is well appreciated in the unconditional or pedigree-averaged case (Uyenoyama, 1986; McClure and Whitlock, 2012; Jarne and David, 2008; Zeitler and Gilbert, 2024). It would be of interest to understand the pedigree effects of inbreeding generally in the conditional case. Campbell (2015) and Severson et al. (2019) studied expected coalescence times for given rates of mating between cousins of varying degree, including sibs or “00-degree cousins”. Severson et al. (2021) and Cotter et al. (2021) extended these models and applied the separation-of-timescales result of Möhle (1998) to obtain the limiting distributions of coalescence times in the unconditional case. Based on the results we presented here, we would expect to see some effects of the population pedigree in conditional coalescent processes for these and other forms of inbreeding.
Acknowledgements
This work was supported by National Science Foundation grants DMS-2152103 and DMS-2348164.
Conflicts of interest
The authors declare no conflicts of interest.